A Bibliography of Accuracy and Stability of Numerical Algorithms, Second Edition

Nicholas John Higham University of Manchester Department of Mathematics

Manchester M13 9PL England

Tel: +44 (0)161 275 5800 FAX: +44 (0)161 275 5819

E-mail: higham@ma.man.ac.uk (Internet)

17 November 2023 Version 2.33

Abstract

This is a bibliography of publications in the second edition of the book Accuracy and Stability of Numerical Algorithms.

Title word cross-reference

(A, B) [KP96a]. (AR − LB, DR − LE) = (C, F) [Kåg94]. (Y A − DZ, Y C − BZ) = (E, F) [Chu87]. 1 [IEC94]. 2 [BZ98, Matc]. 27 [Gol67]. 2 × 2 [Win71]. 3 × 3 [Lad76]. 8 [Gol67]. {a_{ij}} [Asp59]. a_{ij} = 0 [Asp59]. AMD5_{K}86 [MLK98]. AX + XB = C [BS72, GNV79]. AX + Y B = C [Ziȩ85]. AX − XB = C [Hig93c, SN91]. AX − XB = Y [BR97]. AX − XB^{T} = C [Bye84]. AX − Y B = C [BK79]. Ax = b [Har80]. AXB + CY D = E [BK80]. AXB − CXD = E [Chu87]. AXB^{T} + CXD^{T} = E [GWL^{+}92, GLAM92]. AXD − BXC = E [Ept80]. ℓ^{p} [Boy74]. ℓ_{2} [GV74, Wat91]. H [ABO88, Neu86, NP84, Pol87]. I − 2ww^{H} [Wil65b]. j > i + p [Asp59]. l_{2} [CP76, Sha85]. l_{p} [GS83]. LDL^{H} [Bar91b]. LDL^{T} [FP74, Sun92d, Hig99b]. LU [ABO88, Bar91b, Cry73, DHS95, ES92, FNP82, Mat84a, Ste97b]. M [Neu86]. O(n) [Dhi98]. P [Rum02, Hig92a, Yip86]. π [Bai88a]. P ≤ A ≤ Q [Har80]. p ≤ b ≤ q [Har80]. px = xq [Syl84]. QR [Bar94, Dem89a, Dem90, Gen75, Gul95, Hig90g, Hig91a, Kar74, Kni95, Ste77b, Sun91, Zha93a]. sep^{−1} [KP92]. T A − BT = C [Hea77]. x ∗ (1/x) ≠ 1 [Ede94b]. x^{(i+1)} = P x^{(i)} + q [AR92]. XA + BX + C = 0 [Mil88b]. XA − BX = R [Dat88]. ∥A^{−1}∥_{∞} [Var76].

-Decomposition [Kar74]. -Decompositions [Mat84a]. -Digit [Gol67]. -Factorization [Cry73]. -Matrices [ABO88, Neu86, NP84, Pol87]. -Norm [BZ98, Hig92a]. -norms [Rum02]. -Scaling [GV74, Sha85]. -Solutions [CP76].

1-Norm [CH01a, CH01b, HT00]. 10967-1 [IEC94]. 115 [CL83a]. 14th [Tuc02]. 15C [Kah83, Hew82]. 16 [For60]. 16th [Gol77]. 1900-1920 [Bel31]. 1994-12 [IEC94]. 19th [Gol77].

2-by-2 [Mol74]. 2-Norm [MAV^{+}97]. 2.0.1 [Sun92b].

3 [DCDH90b, DCDH90a].

48G [Hew93].

524 [Bre78b].

64 [CHN99].

722 [CC93]. 754 [Hou81]. 754-1985 [IEE85].

854-1987 [IEE87]. 860 [HGD90].

90 [Kea96a]. 90-Based [Bai95]. 90/95 [MR99].

A. [Bod49, Bri92, CD86, Inc92]. Aasen’s [Han97]. Absolute [BSW61]. Accelerate [BLS91]. Accelerated [LPS92]. Acceptability [BRB76]. Acceptance [Coo90]. Access [BRK^{+}87]. Accumulated [Bar81a, Lin76]. Accumulators [Ros65]. Accuracy [BRB76, BRR85, Boh77, CHKL01, CH99a, DK85, DGE^{+}99, Dum98, GS90, GAKK93, Gol67, Gre97a, Hig89a, Hig93a, Hig99a, Hig02, Hof89a, JW86, Kah94, KC73, Lin74, Sch96, CL79]. Accurate [AGL98, DK88, DK90, Dem99, Dem01, FB91, GB91, Hig90f, Hig00, JW85, Lin75, Lin70, Mal71, Mat95a, Pri92, Tan90a]. Accurately [Cli90, SW90, GP90]. ACE [Tur45, CK81, CD86, Wil54a, Wil54b, Wil80]. ACM [ACM87]. ACRITH [BRR85, BRK^{+}87, JW86, KL85]. Activity [CK81]. Adaptive [FGP91, FJ98, PP92, She97, SB92]. Adding [Vit68]. Addition [FG82, FT86, Møl65b, RK75, RS88, Swe65]. Additions [Syl50]. Adjoint [Han97]. Adjugate [Ste98]. Admissible [OPW65]. Admits [Mar94]. Ado [Kah87a]. Advanced [BRR85, Hew82]. Aggregating [Pan78]. Aid [Ver49]. Alan [CK84, Hod83]. Albers [Alb91]. Algebra [Bas02, BL77, BDD^{+}02, Bun87, ILA93, Dem92b, Dem97, DGK84, DDSvdV98, DCHH88b, DCDH90b, DCHH88a, DCDH90a, Ede93b, Ede94a, Fad59, For67, Fox64, GPS90, GAKK93, Gus97, HS87, Hig94b, Hig99a, Kah66, KS88, KS92, LHKK79, Mir61, Mol78, Sto91, Str93, Tod77, Tre90, TB97, Wil54a, WR71, Wil74b]. Algebraic [AL84, Ata40, BCS97, Bye85, Bye87, For53, FM67, FHW48, GL91, GL95a, KH90, KFC^{+}01, Kup71, Lau79, Rob80, vdS70, Sto77, Wil63b, Wil65a, Wil68, Wil70]. Algebraicznych [Wil67]. ALGOL [Nic70]. Algorithm [And99, AL85, Ari00, Bai88a, BF88, BLS91, Bai93, BHH93, Bar91a, BS72, BP92, Bur97, CH98a, CH01a, CH01b, Chu00, Cod93, CT65, DH89, DK85, DCHH88b, DCDH90b, DHSS94, Dum98, GWL^{+}92, GM73, HP94a, HT00, JP91, KF63a, KF63b, Kap99, Kar74, Kea96a, Lad76, LPW83, LZ92, McK62, MS78b, Mol2b, Pan78, Smi62, Smi65, Smi91, Smi98, Ste71, Tha61, Tor89, Tor91, War76, Wat91, Win68, Yal00, Kie73, Bre78b, Cod88a, CC93, CL83a, For60, Hig88c]. Algorithmen [Spi76]. Algorithmic [CdB80]. Algorithms [AL86, AL84, BDG97, BR77, BG76, BL80, BP94b, BE73a, BKO94, Bre70a, Bre70b, Bun87, BDV89, CGL83, CF92, Csa76, DGKS76, DH92, DL94, DDR95, DGK84, Fis74, GPS90, GL91, GM74, Gre94, GK69, GA96, Gus97, HHKR93, HLB01, Hig86b, Hig87b, Hig90h, Hig99a, Hig02, HJ88, Hof89b, Hul82, KP92, KP96a, KP96b, Kah73, KW00, KLP92, Knu97, Knu98, Kre76, KW92, LPS92, LS80, LO82, Lu95, Mal72, Mat96, Mil88a, MW80, Pan78, Pan01, Par00, Par71, Pow75, Pow98, Pri91, Pry85, Row90, Rus98, Sch88, Sed88, Smo95, Ste01, Stu80, Stu81, Tan91, Tor97, Tsa83, Vit68, Woź80, Yal95, Mul97]. Almost [Yal00]. AMD [Rus98]. AMD-K7 [Rus98]. Analyses [CPS96, CPS97, CP01, Kah96a, Mei83, Wil74a]. Analysis [Abd71, AGP97, AM94, And71, AL85, AL86, Bab69, Bar81b, BB85b, Bar88, Bar91a, Bar91c, Bel70, Bha97, Bjö87, Bjö91, BBG92, dBP77, Bre70b, BC86, BT92, Bun71a, But87, Car59, CF92, CH94, CdB80, Dem87b, Eir93, Fox78, Gas83, Gen69, Gen75, Gol77, Gul95, Hen64, Hig87b, Hig90a, Hig90h, HK93a, HK00, Hil74, HJ85, HJ91, Hou64, IK66, JP91, JO74, Kåg94, Kah72, Kah91, Kie87, Kni93, Kör88, Lan66, Lan56, LS80, LPW83, Lin74, Man81, Mat96, Mil75b, Mil76, MW80, MS78b, MS78a, MM83b, Moo66, Moo79, Mül83, NP92, Neu85, NP84, New73, New74, Nic77, O’C93, Oli77, Olv83, Ort63, Ort72, Pai73, Par71, Par90, Pry81, Pry85, Ram71, Row90, Sau78, Sca50, SB94, Sma90, Smi75, Sor85, Ste71, SB80, Sto77, Stu80]. Analysis [Stu85b, Stu85c, Swe65, Tan91, TZ01, Tre92a, Tsa83, VA86, Wen66, vW66, Wil60, Wil61, Wil62, Wil65b, Wil68, Wil71a, Wil85, Wil86, Woź80, YC97, Zha93a, Ziv82, Woź78, CGL83]. Analyst [Wil71b]. André [Bre96]. Andrew [CK84]. Anecdotes [Tro84a]. Angles [Kah97]. Annotated [NR76]. Annual [Ede91b]. Anomalies [KL85]. Anomalous [Ben38]. ANSI [IEE85]. ANSI/IEEE [IEE85]. Answers [Fox71]. Application [Boy94, Ept80, HT00, PV74, Ste80, Tur36, Ben24, Ben24, CB90]. Applications [AB79, ABD92, BS68, BBG92, CH92a, Dat88, Fer95, GDD94, HSS99, Hig86a, Hig88c, Hig89b, Hod92, Hou81, MN79, MN99, MD64, MP77, Moo79, NR76, PS91, Ral81, vR01, Smi65, TG81, Tyr91, Dah83, Kie73]. Applied [Bar89, CH94, Dem97, Lan56, Par64]. Applying [PR69]. Appraisal [Lon67]. Approach [ADdR89, CdB80, DEL00, HP96, Jan97, KM83, KM86, LS80, Man81, Olv78, Tsa83]. Approche [Bru89]. Approximants [Hig01]. Approximate [Kov76, Kup71, OP64, Ziv95]. Approximating [CHKL01]. Approximation [ADW84, EY36, KW00, Lás94, Rei91, Sch83, Tre90]. Approximations [BP67, BBB89, FB91, Fik67, Hun97, SR80]. Arabia [Uni92]. Arbitrary [Pri91]. Architecture [FJ98, HP96, HJ88]. Architectures [JP94]. Area [Kah97]. Argument [Fer95]. Arguments [Var76]. Arising [FGS96, GMPS92, Mat84a, PW71, Rut47, Rut52, Var72]. Arithmetic [AB79, BSBF89, Bar81a, BB85b, BB85a, BRR85, BHY80, Bre78b, Bre78a, Bro81, Car59, COT89, CLP01, Cod73, CCG^{+}84, CC93, CHH83, CW87, CHN99, Dem87b, Dem91, DDR95, Fro01, Gar76, GM74, Gol91, Gre94, IEE81, Hen80, HLB01, Hig91b, HK95, Hou81, JW85, JW86, Kah80, Kah81, Kah87b, Kah90, Kah96b, Kie81, KM81, KM86, IEC94, Mal72, MK85, Met77, Mil88a, Olv78, Olv83, Olv86, Ove01, Par00, Pla88, Pri91, Rum99a, Sch81, She97, Smi91, Smi98, Stu81, Stu82, Swa80, Tan89, Tan90b, Tan92, Tie70, Tis01, Tur91, VCV01a, VCV01b, Wil97, Wil63a, Yoh79b, Yoh80, IEE85, IEE87]. Arithmetics [Bar81b, Cra79, Pri92, SS84, Yoh79a]. Arithmétique [Pic72, PV74, Mul89]. Arrays [BI87, HSS99]. Art [PTVF92, Wil85]. Articles [Syl50]. Aspects [Bar88, Eir93, Mui82, Pai90, Ruh83]. Assertions [BGM^{+}94]. Associated [Jor68, Kni93, NP84, Ste73a, Tra66]. Asymptotic [BW93, Gau53]. ATLAS [WPD01]. Attainable [Bre73, Gre97a, Lás94]. Attained [SW98]. AUGMENT [BHY80]. Augmented [ADdR89, Bjö92, BP94c]. Ausgleichung [Läu61]. Automata [KO63]. Automated [WPD01]. Automatic [BBG92, CR72, Gil51, Gus97, Iri91, Lan01, LK76, Ral81, Sto77, Har46, Tur45, Wil48a, Wil48b]. Available [Dek71]. Average [TZ01, TS90, WWWS86]. Average-Case [TS90]. Award [ACM87].

Babuskasche [Nic70]. Backward [AGP97, AM94, ADD89, BGM^{+}94, dBP77, CI94, CH99b, CH99c, Eir93, Fox78, GL95a, Gu98, Gul95, HH92a, Hig93c, KF63a, KW97, Mal01, MS02, NP84, Sun95a, Sun96, Sun97, SS97, Sun98, WKS95]. Ball [Tao84b]. Banach [Ost73]. Band [Boh75]. Banded [BW93, JP93]. Base [Mat70]. Based [Bai95, CH98a, Lin92, Pug92, Wil62, Wil65b]. Bases [ABCR93]. Basic [AvL81, DCHH88b, DCDH90b, DCHH88a, DCDH90a, LHKK79, Tod77, Bas02, BDD^{+}02, HHKR93, VCV01a]. Basis [Bur97, SB95]. Basis-Kernel [SB95]. Be [Bus68, For53, Kah90]. Became [Tau88]. Beeblebrox’s [Gra92]. Before [Coo94]. Behavior [Gre89, GS92, HK93b]. Behaviour [Gau53, HW76, vdV90]. Beitrag [Hil94]. Benchmark [Kar85]. Benchmarks [Kah94]. Berechnung [Sch33]. Best [FS55, GV74, Lás94, MAV^{+}97, RS88, Sau78, Sto62]. Better [GA96]. Between [KP96b, Syl51]. Beyond [CO84]. Bibliography [NR76]. Bidiagonal [DK90, Fal01]. Bifurcations [Gov00]. Billion [BBB89]. Binary [Buc59, CC93, CHN99, Gay90, IEE81, Kah96b, Pax91, SS84, IEE85]. Binary-Decimal [Gay90]. Bisection [Bye88, DDR95]. Bisection-Like [DDR95]. Bisectors [Par71]. Bit [Kah87a, Ziv91, Tro84b]. Bits [Gol67]. Björck [Hig87b]. BLAS [BDD^{+}02, DH92, DHSS94, Hig90e, LDB^{+}00]. BLAST [Bas02]. Bledy [Wil67]. Block [CH01a, CH01b, CGM85, Dem83, DH92, DHS95, ES92, FV62, GP90, Hig99b, HT00, JP91, KP92, Mat84a, Mat84b, Meu92, Pol87, Sch88, SP88, Sha85, TG81, Var72]. Block- [ES92]. Block-Tridiagonal [Var72]. Blocking [Gus97]. Blocks [BD93b, BBC^{+}94]. Blockwise [Pol87]. Boeing [DGL92]. Book [For70, Wol99]. bordered [GP90]. Borweins’ [Bai88a]. Bound [Lás94, MAV^{+}97, Tao84a, Var75]. Boundary [BR77, CH94]. Bounding [Hig90b, IK00, Var76]. Boundless [Ske89]. Bounds [Bar91b, Bar94, Bjö91, Boh75, CG67, CR79, CH99b, DH93, DDM01, GL95b, Gol62, Gu98, Hen62a, Hig90c, HT01, Jen82, KW97, Kit95, Lem75, Mal01, MS02, OP64, OW82, Olv86, Peñ96, SW98, Sha91, Ste73a, Ste77b, Sun91, Sun92c, Sun92d, Sun95b, Sun96, Sun97, SS97, Sun98, WKS95, Ziv95]. Brain [Gra92]. Branch [Kah87a]. Brent’s [BHY80]. Brief [AM87]. British [WCK89, Lav80]. Buckley’s [IK00]. Buffon [WT94]. Bug [Ede97]. Bugs [Gla95]. Building [BBC^{+}94]. Built [Lav80]. Bunch [JP93]. BVP [Mat84a]. Byte [Buc81].

C [KKL^{+}93]. C-XSC [KKL^{+}93]. CADAC [CHH83]. Calculating [BK77, Ver49, Har48]. Calculation [AM65, CT65, Hot43, LS78]. Calculations [Act96, Pow98, Str86]. Calculator [Blo48, Sch83, Smi75, Har46]. Calculus [MN99]. Can [Bus68, For53, Fos94, Pan84]. Canceling [Pan78]. Cancer [Ber91]. Canonical [GW76, TA32]. carrés [Ben24]. Cascades [SS84]. Cascading [Ros65]. Case [BR77, TZ01, TS90, Ziȩ85]. Cases [LM01].

Cauchy [Tyr91, Cho68]. Cauchy-Toeplitz [Tyr91]. Causes [Mar94]. Celebrating [EG91]. CELEFUNT [Cod93]. celui [Ben24]. Century [Ewi94, Gol77]. Certain [And90, CH94, JW85, Peñ96, Ste73a, Sto62, Var72]. Certification [PY01]. CESTAC [BC86]. CG [vdV90]. CG-S [vdV90]. Chains [IM94, O’C93, Ste97a]. Change [DW93]. Chaos [PJS92, SSL93]. (Procédé [Ben24]. Characteristic [Giv54, Tau49]. Characteristics [Cod73]. Characterization [GV74]. Chebychev [CR92]. Chebychev-Vandermonde [CR92]. Chebyshev [ADW84, BKO94, Cle55, GO94, Oli77, RO91, Woź77b]. Checked [MLK98, Rus98]. Checking [PR93]. Chemistry [Rut47, Rut52]. Chemotherapy [Ber91]. China [LYK89]. Chip [Gla95, Mar94]. Chips [SW89]. Choice [Buc59, Dan76, Hul82]. Cholesky [Ano22, Bre96, Car91, CPS96, CH98a, Dem89b, DOV94, FGM95, GSS96, Hig90a, Kie87, Mei83, Pow86, Sau72, Ste93b, Ste97d, Ste97b, Sun91, Sun92d]. Cholesky) [Ben24]. Circle [FV62]. Circuit [Mar94]. Circulant [Dav79]. Circular [Lin92, Yal00]. Class [Dub83, IK00, KKL^{+}93, MP74, Syl50, Woź80]. Classical [Wil74a]. Classroom [PJS92]. Clenshaw [Oli77]. Code [DR95]. Codes [Hig88c]. Coefficient [BR77, Doy57]. Coefficients [Cro41, Gen69, OP64, OPW65, Oet65]. Collected [Bri92, Inc92, Tau63, Syl04]. Collection [BDDD97, DGL92, GK69, Wri93]. Collections [BPR^{+}97]. Collocation [CH94]. Combination [Gau95]. Combinatorial [Hal67]. Combinatorics [WSW72]. Combining [Fis74]. Commandant [Ano22]. Commemorative [Fox78]. Comment [Bjö78]. Comments [IEE81, LK76, Wil71b]. Commutation [MN79]. Compact [Pug92]. Companion [EM95b, KL88, Kit95, TT94]. Comparison [BG76, Neu86, SS62]. Compatibility [Kov76, OP64, RG67]. Compiler [Far88b]. Complement [Cot74]. Complete [Ede92a, EM95a, Gou91, Hig00, HV97, Tho42, Vav89]. Complex [AvL81, BMPW66, Cod93, CT65, Cro41, Fra58, Hig88c, Hig92b, Hig98a, HFT94, IK00, Kah87a, Olv83, Olv86, Smi98, Ste85, Smi62]. Complexity [AvL81, BCS97, DDM01, Lic87, Mil75a]. Compliance [Rus98]. Component [Bjö91]. Component-Wise [Bjö91]. Components [CI95]. Componentwise [CP01, GK93, HH92b, HK93a, Hig94b, Rum99b, Rum02, Smo95, Sun92c, Zha93a, Dem92a]. Comportement [Bru89]. Composite [Per02]. Computable [CG67, Tur36]. Computation [Bai88a, BKP93, BRR85, BRK^{+}87, Boh77, CH90, Dhi98, Dub83, DG94, Far88b, Fer95, For70, Fox71, Giv54, GW76, GDD94, Hig00, Hof89a, Kah90, KFC^{+}01, KM83, Leh96, Mai84, MS02, Pan00, PY01, Ral91, SP88, Ske89, SA56, Ste74a, Thi88, Var90, Wil60, Sun92a]. Computational [Bau74, BRB76, CF92, CKP95, Fad59, For67, GK69, Ker90, Lic87, Van92, Mil75a, vR01, Tod61]. Computations [AH83, BP94b, CCF96, Dwy51, Far88a, FMM77, GPS90, Gen73, GV96, GS92, Hig93b, Kah96a, Kea96b, Lin75, Lin81, Mat95a, Mol2a, Mol90b, New62, Pri92, Ste73b]. Compute [Bar91a, BBB89, MV78]. Computed [GL95a, Gre97a]. Computer [AM65, BN71, Fik68, FM67, For69a, FMM77, Fra92, Gar76, GL81a, Gol91, Gol72, HP96, Hof87, Kah94, Kah01, KM81, KM86, Lon67, Par00, PH98, SB77, Sch81, Sto77, Sto91, Swa80, War60, WCK89, Knu62]. Computerorientierte [KS88]. Computers [BDD^{+}97, BMSV92, DDSvdV98, Don01, HJ88, Lav80, Ran75, Vit68, War76, Wil74b, Woo58]. Computing [Act96, AL86, Ata40, BDM93, BD93a, Bjö90a, BCP^{+}97, BC86, CL79, CGL83, CV86, DK87, DGE^{+}99, DDM01, Ede93b, FGM95, Fra58, Gen69, Gil51, HHKR93, Hig86a, Hig86b, Hig87a, HS87, Hig88a, Hig90c, HP94a, Inc85, KP96a, KW00, KG80, KKL^{+}93, Ove01, PTVF92, SK93, SS74, Tao84a, Tao84b, Thi88, Tur45, Wil48a, Wil48b, SA73]. Concerning [Har80, Tau49, Tod61]. Concrete [GKP94]. Condition [And91, BDM93, Bau63a, Bec00, Bis90, BLP90, Bye84, Bye85, CF92, CH94, CMSW79, CCV82, CR83, Dem83, Dem87a, Dhi98, Dix83, Ede88, Ede92b, FGP91, Gau83, Geu82, GK93, GL81b, GEJ95, Hag84, Hig86b, Hig87c, Hig88c, HH92a, Hig95a, KW89, KP96a, KL94, KLR98b, KLR98a, Lem75, Van87, Loi69, MAV^{+}97, O’L80, PP92, Ric66b, Roh89a, SB92, vdS69, vdS70, Ste80, Tod54, Tod68, VT98, WWWS86, Zie88, ZWS86]. Conditioned [CF88, FGS96, FS55, GW76, Men56, Rum99b]. Conditioning [AR92, Fle85, MS74, Rhe76]. Conditions [Sha82]. Conferences [WCK89]. Confluent [BE73a, Dem90, Gau62, Gau78, Hig90h, Lu94, Lu95, Lu96]. Conjecture [Ede92a, EM95a, PC83]. Conjectures [Fro01, Var90]. Conjugate [Ari00, CN96, CGM85, Gre89, GS92, Gre94, Not93, Woź80]. Conjugate-Gradient [Gre89, Woź80]. Connection [BHH93]. Conquer [BDG97]. Constant [BR77, BR71b]. Constrained [Bar88, BNP88, BV92a, CH99a, CH99b, CH99c, FGS96, GW92, Gul94, Gul95, Jan97, Van85, Swa74, Wri92]. Constraints [Eld80]. Constructing [Chu99, Chu00, Ikr98, KS00, Pan78, ZZ95]. Construction [BP67, Wil80]. Continuant [Rut47, Rut52]. Continued [Bla64, JT74]. Continuous [JW85]. Contrôle [PV74]. Contribution [Bru89, Geu82, Par90]. Contributions [Gar76, Bel31]. Control [DK88, Enr89, Hod92, Hul82, SR80]. Controllability [KL88]. Controlled [CHH83, Har46]. Controlled-Precision [CHH83]. Controlling [GMSW91].

Convergence

[AN63, AR92, Dax90, LRWW98, McK98, Not93, Tao84a, Tor91, Tor97, vdV90]. Convergent [Bai88a, MP77]. Conversation [Woe97]. Conversion [Mat70, Pax91]. Conversions [Gay90, Mat68, VCV01b]. Converting [Ziv95]. Coordinex [PC83]. Corner [Mol90a]. Correct [LM01, Sch84]. Correction [Bar88, BV92a, Lan01, Pic72]. corrections [Kie73]. Correctly [Gay90, LMT98, Ziv91]. Correctness [DDR95, Hul79]. Correlation [DH00]. Cost [Pri92]. Counter [CR83]. Counter-Examples [CR83]. Course [Har67, Ort72]. Courses [ILA93]. Cramer [Mol74]. Cray [Bai88b, Kah90]. Cray-2 [Bai88b]. CRD [Bar81b]. Crest [Jos91]. Crimes [Sto91]. Criteria [ADR92]. Criterion [Ada67, Ari00, Kar74]. Critical [Wic92]. Cross [GA96]. Crout [For60, McK62, Tha61]. Cryptograms [MM83b]. CS [PW94]. Cubic [Kah86]. Curve [Smi65]. Curve-Fitting [Smi65]. Cuts [Kah87a]. Cyclic [AM94, BG94, Zaw91].

D. [Kah73]. d’Algorithmes [PV74]. d’Analyse [Cau41]. Dangers [BCP^{+}97]. Danilewski [Fra58]. Dans [PV74]. Data [CR79, Kea96a, RG67, Roh90]. David [IEE81]. Decimal [Bai88a, Buc59, CHH83, Gay90, Pax91]. Decimal-Binary [Gay90]. Decimals [Leh43]. Decomposition [BD93a, BDD^{+}97, Bre70b, BKP76, Bun82, Cho68, DGE^{+}99, GDD94, GW92, Hig86a, Hig90a, Hig94a, HP94a, HV97, Kar74, Kni95, Liu87, PW94, Pow86, Rei91, Ste93a, TG81, Tol97, Zha93a, Bar86]. Decompositions [Bar91b, CI94, Dem99, FNP82, Gen75, Mat84a, Mat84b, Ste97c, Sun92c]. Deconvolution [Lin92, Yal00]. Defense [Uni92]. Deferred [Bar88, BV92a]. Defined [OPW65]. défini [Ben24]. Definite [BR71a, Bec00, CV86, Doy57, GL81a, GV79, Ham82, Hig98a, MPW66, Mat92, Van95, Hig90a]. Definiteness [Ker90]. Definiteness/Semidefiniteness [Ker90]. Definition [dJ77, Tre92a]. Deivations [CL79]. Denormalized [Coo81]. Dense [DGK84, Ede91b, Ede93b, Ede94a, GPS90, Gus97]. d’équations [Ben24]. Derivatives [Kub91, Pow75, Smi65]. Deriving [HS81a]. Descent [Bol84]. Design [LDB^{+}00, PH98]. Designing [DDK92]. Designs [Gla95, Par00]. Detection [BV92b]. Determinant [Fra58, LZ92, PY01]. Determinants [Caf63, Cro41, Mak91, Rut47, Rut52, Syl51, Tur29, Bel31]. Determination [Man81, Sch84]. Determine [Cod88a]. DEUCE [Wil80]. Development [Ste77c, Tur45, Ypm95]. DFT [BH95]. Dhahran [Uni92]. Diagonal [BD93b, Bun71a, CL83a, CL83b, Chu99, Hig97b, Ikr98, Peñ98a, Sha85, Sha91, SW62, Var76, ZZ95]. Diagonalize [Dem83]. Diagonally [FV62, FNP82, Mal00]. Dictionary [Nel98]. Difference [Fer95, Hen63, Per02, Var72]. Differences [KF63a, KF63b, SvdVM00]. Different [New81]. Differential [BP67, But87, CKP95, Ept80, Gil51, HW91, Hen62b, KH90, MN99, Sha94, Vit69, Dah83]. Differentiation [Iri91, LM66, Ral81]. Digit [Boy94, Gol67, Rai76]. Digital [Gil51, Knu99, KM86, Ran75, Wil74b, Har48, Woo58]. Digits [Bai88a, BBB89, New81, Rai69, Tur82]. Dimensions [LRWW98]. Direct [Ber91, BP71, BGN70, CG67, DT91, DRMN79, DER86, Hig93b, JO74, LTT00, Pow98, Swa72, Swa74, Tor89, Wil61, Wri96, DR95]. Directional [Tor89]. Directions [FGM95]. Discovery [Coo94]. Discrete [ADW84, BP67, BH95, GLKH90, GL95b, Hen62b, Ple76, Sch96]. Discrete-Time [GLKH90, GL95b]. Distance [Bye88, Dem87a, Rum02, Ziv82, Dem92a]. Distillation [And99]. Distributed [BDD^{+}97, KP92]. Distribution [Bar81a, DG87, Ede91a, Ede92b, Rai69, Tur82, Ura68]. Distributions [BB85a, BG79]. Diversions [Gar61]. Divide [BDG97]. Divided [KF63a, KF63b]. Division [Ede97, KM97, Lic87, MLK98, Rus98, Smi62, Ste71, Ste85, Tsa83, Tur45]. Division-Free [Tsa83]. Do [Fox71, GA96, Kah81, Pow75, Kah01, KS92]. Dominance [Peñ98a, Var76]. Dominant [FV62, FNP82, Mal00]. Donald [Alb91]. Done [Kah90]. Doolittle [Far87]. Double [HLB01, LM01, Møl65a, Møl65b, Mol90b]. Double-Precision [Møl65a, Møl65b]. Double-rounding [Mol90b]. Doubled [Kah87b, Lin81, SS84]. Doubled-Mantissa [SS84]. Doubled-Precision [Kah87b, Lin81]. Draft [IEE81]. Drift [RK75]. Driven [Tan89, Tan90b, Tan92]. Dual [Wri97]. Dubious [MV78]. d’un [Ben24]. d’une [Pic72]. durch [Spi76]. Dynamic [AP94, Cod73]. Dynamical [Gov00]. Dynamically [Cod88a].

Early

[HGD90, Lav80, Ste93a, WCK89, CK81]. Earth [QTD91]. Econometric [MV99]. Econometrics [MN99]. Edition [Tho42]. Editor [Edi91]. EDSAC [Wil97]. Effect [Ske81, Ypm83a]. Effective [CH94]. Effectively [CF88]. Effects [LS78, SvdVM00]. Efficiency [Fos97]. Efficient [BMSV92, GDD94, Hig86b, LS78, SV89, Ste80, Tan90a, VA86, KFC^{+}01]. Eigendecompositions [DK87]. Eigenproblem [AL84, BDM93]. Eigenproblems [BDG97]. Eigenspaces [KP96a]. Eigensystem [GBDM77, Mat95a, SBD^{+}76]. Eigensystems [GW76, Ste01]. Eigenvalue [BDDD97, CI94, CB90, CV86, DDR95, Ede91a, Ede93a, KW92, LZ92, Ort63, Par64, Par65, Par98, Rut58, Ste73a, Sun95a, Tis01, TM01, Wil62, Wil65a, Wil70]. Eigenvalues [And90, CL83a, CL83b, Cha93, Chu00, Dix83, Ede88, EKS94, EM95b, Fra58, HT01, Hym57, Ikr98, KP96a, Van87, RT92b, TM62, ZZ95]. Eigenvectors [Hym57, Van87]. Einführung [KS88, Sto72]. einiger [Neu74]. EISPACK [GBDM77, SBD^{+}76]. Electrodynamics [vR01]. Electronic [DG87, Lon67]. Element [Ari00]. Elementary [CW80, Cod93, CdB80, GB91, HFT94, Kah87a, LM01, Leh96, Men56, Mul97, SK93, Stu80, Tan91, Ziv91]. Elements [CL83a, CL83b, CF92, Chu99, Hen64, Ste74b, Thi88]. Elicited [Sev98]. Elimination [AGP97, BB85b, BZ98, Boh75, CG85, Cli73, Coh74, CH99c, Cry68, CR72, DK77, Dax82, DP88, Ede92a, Fos94, Fos97, GP92, GIK02, Gou91, HH89, Hig90b, Hig90f, IK00, LD98, Mac89, Mal00, MRW01, NP92, OW82, PW75, Ple74, Rei71, Sau78, Ske79, Ske80, Sor85, Ste74b, Ste95, Ste97c, Str69, Stu82, Stu85b, Stu85c, Stu85a, Tre85, TS90, Vav89, Wri93, YC97, van77, Bus71, GP90, Läu61]. Elliptic [BR77]. Empirical [WPD01]. endlicher [Neu74]. Engine [Tur45, Wil48a, Wil48b]. Engineering [Act96, BRK^{+}87, IBM88, MD64, Mil84]. Engineering/Scientific [BRK^{+}87]. Engineers [Ham73]. Enhances [Hig90g, Hig91a]. Enigma [CK84, Hod83]. Enough [For70, Gol67]. Entries [Ikr98, ZZ95]. Entrywise [O’C93]. Entscheidungsproblem [Tur36]. Environment [For78]. Equality [Bar88, BNP88, BV92a, CH99a, CH99c, Eld80, Van85]. Equality-Constrained [BNP88, Van85]. Equation [AGL98, BK79, BK80, BS68, BS72, BR97, Bye84, Bye85, Bye87, Dat88, Dou59, For69a, For69b, GLKH90, GWL^{+}92, GLAM92, Ham82, HK88, HK00, Hod92, HR92, KW89, KP92, Kåg94, KP96b, Kah86, Mil88b, Mol2b, Rob80, Ste92, Syl84, Wac88, Ziȩ85].

Equations [ABCR93, AL84, Ata40, BSD60, BG76, BP70, Bjö87, Bol84, BBB89, BMPW66, BP71, But87, BGN70, CI95, CG67, Chu87, CKP95, Cro41, Dax90, DK85, DS83, Don01, DRMN79, DR82, Ept80, For53, Fos91, FHW48, GL91, GL95a, GL95b, Gil51, GS76, Gri85, HW91, Hen62b, Hig90g, Hig91a, HS95, Kel95, KH90, KLP92, Kov76, Kup71, Lan70, Lau79, LYK89, MPW66, Men56, OP64, Ost73, Pai73, Pen98b, PW71, Rhe76, Roh89b, Sha94, Sny55, Usp48, Var72, Ver49, Vit69, WWWS86, Wes68, Woź77a, Zie88, Ben24, Dah83]. Equilibration [Bun71b, McK62, Ske81, vdS69, vdS70]. Equilibria [Gov00]. Equivalence [Dem83]. Equivalent [Syl51]. Ergodic [Pet81]. Erreurs [PV74]. Error [Abd71, AB79, AGP97, AM94, AL85, AL86, ADD89, AM65, Bar81a, Bar81b, BB85b, BB85a, Bar88, Bar91a, BR71b, Bjö91, BBG92, BGM^{+}94, dBP77, Bre70b, BC86, Car59, CG67, CB90, CH94, CG85, CH99b, Dem87b, DH93, DDM01, Dou59, Eir93, Enr89, Fox78, Gen69, Gen75, GL95a, GF75, GFB85, Gul95, Hen63, Hen80, Hig87b, Hig90b, Hig90c, HH92a, HK93a, Hig93c, Iri91, JO74, Jor68, Kah72, Kah96a, Kie87, Kni93, Kub91, Lan66, LS80, LPW83, Lin76, Lyn64, Mei83, Mül83, Neu85, NP84, New73, New74, O’C93, OP64, Oli77, Oli79, Olv78, OW82, Olv83, Olv86, Ort63, Pai73, Pry81, Pry84, Pry85, Qui94, QT90, Ram71, Sau78, SR80, Ste71, Ste73a, Sto77, Stu80, Stu82, Stu85b, Stu85c, Stu85a, Sun92d, Tan91, TZ01, Tie70, Tsa83]. Error [Ura68, VA86, Wil60, Wil61, Wil62, Wil65b, Wil68, Wil71a, Wil74a, Wil85, Wil86, Woź80, Zie88, Ziv82, Ziv95, Woź78, Stu85b]. Error-Control [Enr89]. Errors [Act96, Bau74, Bod49, Boh75, CI94, Dem89b, Des63, For59, Gau95, Gol62, Hen66, HS66, Kah65, KW00, Lan01, LS78, Lin75, Not93, Peñ96, Ros65, SZ80, Ske89, SvdVM00, Sun98, Tur48, Wil63b, Wol64, Ypm83a, vdV90]. Estimate [BR71b, CMSW79, GEJ95].

Estimates [Gau75, GKL95, Hag84, KL94, Kub91, Stu82, Stu85b, Stu85a, Zie88]. Estimating [CG85, Dix83, Dum98, Gre97a, Hig88c, Hig92a, KP96b, KW92, Van87, O’L80]. Estimation [And91, AM65, Bar74, Bar91c, Bis90, BLP90, CH01a, CH01b, FGP91, GL81b, Hig87c, Hig88c, HT00, Iri91, KP96a, KW97, KLR98b, KLR98a, Lin75, Mei69, PP92, SB92]. Estimator [Bye84, CCV82, Hig90d]. Estimators [CR83, KW89, KP92, Ste80, ZWS86]. etc [Kah83]. Etude [PV74]. Euclidean [Blu78, Ost73]. European [Jos91]. Evading [RK75]. Evaluate [PS73]. Evaluating [Cro41, Fik67, Hig01, JT74, Oli77]. Evaluation [Bla64, EGL^{+}87, Fik67, Fik68, Fra58, Knu62, KFC^{+}01, Van79, New73, New74, NT58, Oli79, Olv86, Stu81, Ziv91]. Ever [Qui83]. Every [Gol91]. Exact [Fer95, LO82, Ziv95]. Examples [BN71, CR83, KKN^{+}92]. Exception [DL94, Hul82, HFT94]. Exceptions [Hau96]. Exchange [Qui83]. Execution [Bai93]. Exercices [Cau41]. Exercises [Ove01]. Existence [Pan00]. Expansion [Lin76]. Expected [Fle85]. Experience [BCP^{+}97, HGD90, Hig90d]. Experiments [Jor68, Ric66a]. Explained [MR99, Sch89]. Explicit [Lan70]. Exploiting [Hig90e]. Expm1 [Tan92]. Exponential [MV78, Tan89, Tan90a]. Expressions [AB79, BP67, Stu81]. Extended [DCHH88b, DCHH88a, KKL^{+}93, LDB^{+}00]. Extending [Dek71]. Extension [BRK^{+}87, GBDM77, Lic87]. Extra [Bai88b]. Extremal [Dix83]. Eyes [Ewi94].

Factor [Fos97, GIK02, IK00]. Factoring [JP94]. Factorization [ABO88, ABD92, Bar94, Car91, CH92a, CPS96, CPS97, CP98, CP01, CH98a, CH98b, Cry73, DGKS76, Dem89a, Dem90, DHS95, DEL00, DOV94, DGR^{+}91, ER74, FGM95, GSS96, Gul95, Hig90g, Hig91a, Hig99b, Hig00, JP93, Kie87, Mei83, Sau72, Ste77b, Ste97d, Sun95b]. Factorizations [BP94c, Fal01, FP74, Pai90, Pan00, Pol87, SB92, Ste93b, Sun91, Sun92d]. Factorizing [Fle76, Hig98a]. Factors [DH00, HH89, Ste97b]. Fail [Fos94]. Failure [Uni92]. Fallacies [Ker90]. False [Ede92a]. Far [Mat24]. Fast [ABCR93, AP94, BL80, BKO94, CR93, Chu00, CLP01, Csa76, Dem89a, DH92, EGK90, GS66, GO94, Hig88b, Hig90e, Kni95, Lu94, Lu95, MP74, Pan78, PP92, Ram71, Rat82, Rei91, Rum99a, Sch96, She97, Ziv91, Van92]. Faster [DL94, Inc85, Kap99]. Fault [BGM^{+}94]. Fewer [Pry81]. FFT [CT93, Coo94, FJ98, TZ01, Coo90]. FFTW [FJ98]. Fiabilité [Bru89]. Fields [Hen62a, Lic87]. Fifty [Gra92]. Fifty-Ninth [Gra92]. Figures [Qui83]. Find [Blu78]. Finding [Ada67, CL83a, Wil55, Ypm83b]. Fingers [Buc59]. Finite [Ari00, CCF96, GS92, Gre94, HK93b, HK95, MK85, Mil88a, SL54, Tau49, TM62, Tod54, Var72]. Finite-Difference [Var72]. First [Ede91b, Rai69, Rai76, ACM87]. Fists [Buc59]. Fit [AN63]. Fitting [Smi65]. Five [GA96]. Fixed [Ura68]. Fixed-Point [Ura68]. Flaw [Mar94, SB94]. Floating [And99, BSBF89, Bar81a, Bar81b, BB85b, BB85a, Bar89, Boh77, BGM^{+}94, Bre73, Bro81, Car59, Car91, CO84, Cli90, Cod73, CCG^{+}84, Cod88b, CC93, CHN99, Dek71, Dem89b, Dem92c, DDR95, Dem01, Esp95, Far88b, Fat82, FG82, FT86, GB91, GM74, Gol91, GFB85, IEE81, Hau96, Hen80, HLB01, Hig91b, Hig93a, Hou81, JSW83, KP79, Kah80, Kah81, Kah87b, Kah96b, Kar85, KC73, IEC94, LO82, Lin74, Lin75, Lin81, Lin70, Mal71, Mal72, MI81, Mat70, Mol67, Møl65b, MLK98, Ove01, Pla88, Pri91, Pri92, RK75, RS88, Rus98, Sch81, Sch84, Sev98, SB94, She97, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, Tis01, VCV01a, VCV01b, Vit68, Wal90, Wic89, Wic92, Wil60, Wil63a, NAG86, IEE85, IEE87]. Floating-Point [And99, Bar89, Bre73, Bro81, Cod73, CCG^{+}84, Cod88b, CC93, CHN99, Dek71, Esp95, Far88b, Fat82, FT86, Gol91, GFB85, Hau96, Hou81, JSW83, KP79, Kah81, Kah87b, Kah96b, Kar85, Lin74, Lin75, Lin81, Lin70, Mal71, Mal72, MI81, Mat70, MLK98, Pla88, RK75, RS88, Sch81, Sch84, Sev98, She97, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, VCV01a, VCV01b, Wal90, Wil60, Wil63a, NAG86, IEE85, IEE87, Boh77]. Flottante [Pic72]. Flushes [Gla95]. Form [Aas71, AR92, BD93b, Doy57, GM73, GW76, KL88, MH57, Tao84b, Tho42, Wil65b]. Formal [Bar89, dJ77, Wic89]. Formalization [Mat70]. Formula [Har72, Yip86]. Fortran [Kea96a, MR99, Bai93, BRK^{+}87, CL83a, GWL^{+}92, Hig88c, PTVF92, Smi91, Tro84a, Sun92a, Sun92b, Bai95, Blu78, Bre78b, Bre78a, CGT92, DCHH88b, DCHH88a, DS94, DR82, GMSW86, Kub91, LHKK79, MD64, Mol2a, DR95]. FORTRAN-77 [GWL^{+}92]. Fortran-90 [DS94]. Forum [Bas02]. Forward [KF63b, Stu85b, Stu85c]. Foundations [Sma90]. Fourier [Kör88, Van92, Boy94, BH95, CT65, GS66, Gen69, New73, Oli77, Ram71, Sch96]. FPV [NAG86]. Fractals [PJS92]. Fractions [Bla64, JT74]. Framework [Ari00]. Frameworks [Van92]. Free [BDG97, MI81, Tsa83, WSW72]. Frequency [New81]. Fun [GS66]. Function [Bye87, Cod82, Fik67, NM65, Rob80, Sch83, Tan89, Tan90b, Tan92]. Functional [Row90]. Functions [Boh77, CW80, Cod93, FB91, Fik68, HFT94, Kah87a, KL94, LM01, Mul97, Smi98, Syl51, Tan90a, Tan91, Ziv91, Hew82, CC93]. Fundamental [Knu97, BP94b]. Further [Fis74, Kah65, Tur84]. Future [BSBF89].

Gained [Coo90]. Gateway [ETxx]. Gauss [For51, AM87, BR71a, DH89, Ham87b, Mal00, PW75, Sau78, Ste95]. Gaussian [Fos94, Str69, Vav89, BB85b, BZ98, Boh75, Bus71, CG85, Coh74, Cry68, CR72, DK77, Dax82, DP88, Ede92a, Fos97, GIK02, Gou91, HH89, Hig90b, Hig90f, IK00, Kos92, LD98, Mac89, MRW01, NP92, OW82, Rei71, Ske79, Ske80, Sor85, Ste74b, Ste95, Ste97c, Stu82, Stu85b, Stu85c, Stu85a, Tor65, Tre85, TS90, Wri93, YC97, van77]. GEMMW [DHSS94]. Genauigkeitsfragen [Bau66]. General [But87, GMSW91, Ham87b, Hym57, KL94]. Generality [PW94]. Generalizations [FV62]. Generalized [ABD92, AL84, BD93a, Bar94, CM79, FNP82, GIK02, KW89, Kåg94, KP96b, Lu95, MN89, NR76, Pai90, PS91, Pen98b, SS74, Ste92, Tis01, Zie86]. Generalizing [CCV82]. Generated [BR71b]. Generation [AOU87, DH00, DM89, Ste80]. Genius [Dun90]. Geometric [Hof89a, KFC^{+}01, Mil88a, NP92, She97]. Gerling [For51]. Gerschgorin [FV62]. Get [Fox71]. Given [OP64, RG67]. Givens [BI87, Gen73, Gen75, Ham74, Rat82]. Gleanings [Mat24]. Gleichungssysteme [Bau66]. gleitendem [SB53]. GMRES [DGRS95]. Goertzel [Gen69]. Gradient [Ari00, CN96, CGM85, Gre89, GS92, Gre94, Woź80]. Gradients [Not93]. Graduate [ILA93]. Gram [Abd71, Bjö67b, BP92, Bjö94, DGKS76, Hof89b, JP91, Ric66a, Ruh83]. Graph [Mil76]. Graphs [Bau74, Yal95]. Great [Dun90]. Greedy [Bur97]. Grooks [Hei92]. Growth [BZ98, DP88, Fos97, FS80, GIK02, Gou91, HH89, IK00, Jor68, Tho42]. Guaranteed [GAKK93]. Guide [ABB^{+}99, BBD^{+}01, BCC^{+}97, DBMS79, DGL92, GBDM77, GMSW86, HH00, IBM88, SBD^{+}76, NAG86, Hew93, Matc, Sun92a, Sun92b].

H [Far87, Par90, Eva76]. Hadamard [Wal93, Bir75, EM95a, WSW72]. Halmos [EG91, Alb91]. Hand [HH92b, OP64, Sun96]. Handbook [Hig98b, Wes68, Hew82]. Handling [DL94, Hau96, Hul82, HFT94]. Hankel [Bec00]. Hard [PR93]. Hardness [Roh95]. Hardware [Par00, PH98]. Hardware/Software [PH98]. Hardy [Fox87]. Harwell [DGL92]. Haugeland [Kah73]. Heat [Dou59]. Hermitian [BDDD97, Mat92, Sun95a]. Hessenberg [CS86, GNV79, Ike79]. Hessenberg-Like [CS86]. Heterogeneous [BCP^{+}97]. High [Bai88b, BF88, BMvN46, DDK92, DGE^{+}99, DDSvdV98, EGK90, Fat82, GS90, GvN51, JW86, JP94, KM97, vNG47, BRR85]. High-Accuracy [JW86, BRR85]. High-Level [Fat82]. High-Order [EGK90]. High-Performance [DDSvdV98, JP94]. High-Precision [KM97]. High-Speed [BF88]. Higham [GIK02]. Highly [APS93, Per02]. Hilbert [Cho83, Pow86, SL54, Tau49, Tod54, Tod61]. Historical [You89, Ypm95]. History [AM87, Bel31, Boy89, Gol77, HPS83, Iri91, PW94, Ste93a, Woo58]. Hodges [CK84]. Homotopy [LZ92]. Homotopy-Determinant [LZ92]. Horner’s [Mül83]. Hough [Kah73]. Hough’s [Sch89]. Householder [BV87, BG67, BG65, CH98b, PR69, Pug92, SV89, Tsa75]. Householder’s [Dan76, Ort63]. HP [Kah83, Hew82, Hew93]. HP-15C [Kah83, Hew82]. Hundred [Ove01]. Hurts [Kah90]. Hyman’s [War76].

IA [CHN99]. IA-64 [CHN99]. IBM [KL85]. Identity [Bre70b]. IEC [IEC94]. IEEE [CC93, CHN99, Dem91, Fat82, GB91, Hou81, Kah80, Kah87b, Kah96b, Ove01, Pax91, Rus98, Tan89, Tan90b, Tan92, IEE85, IEE87]. II [Bjö68, GvN51, HW91, MS78a, NP92, Rut52, Ste01, Stu85c, Tha61, VCV01b]. III [Gau78]. Ill [Dem87a, DK88, FGS96, GW76, Men56, Rhe76, Rum99b]. Ill-Conditioned [FGS96, GW76, Rum99b]. Ill-Conditioning [Rhe76]. Ill-Posed [Dem87a, DK88]. ILLIAC [Sny55]. Imaginary [Hig98a]. Implementation [Bar88, Cap71, CH01a, CH01b, Cod82, DCHH88b, DCDH90b, Kah73, LDB^{+}00, Mul97, Tan89, Tan90b, Tan92]. Implementations [Mil88a]. Implementing [DGK84, HFT94, Tsa75, Yoh79a]. Implications [Fat82, Mol90b]. Implicit [Ept80, Rei00]. Implies [JW77, Ske80]. Importance [Bau63a]. Impossible [Mol90a]. Improbability [Kah96a]. Improved [DH93, PS91]. Improvement [DW84, JP91, Sny55]. Improving [Dum98, Lin74, Wal90]. In-and-Out [Mat68]. Inaccuracy [DW84]. Inaccurate [CR79, Kov76, Oet65]. Including [Ove01]. Incomplete [Pol87]. inconnues [Ben24]. Inconsistent [Ziȩ85]. Incremental [Bis90, BLP90]. Indefinite [AGL98, BG76, BP71, CH98a, Che98, DRMN79, DGR^{+}91, Fle76, GMPS92, JP93, JP94, SvdVM00, DR95]. Independent [CCG^{+}84, IEC94, VCV01a, VCV01b, vW66, IEE87]. Index [COT89]. Inequalities [FH55, HLP52, Mat92]. Inertia [BK77, Che98, GMSW91]. Inertia-Controlling [GMSW91]. Inexact [DES82]. inférieur [Ben24]. Influence [Ede93b, Ede94a]. Information [IEC94]. Initial [Enr89]. Inner [Neu85, Win68]. Instability [Mat95b, TT87]. Instrumentation [BBG92]. Integer [IEC94]. Integral [ABCR93]. Integration [DR84, Gil51, QTD91]. Integrations [Qui94, QT90]. Intel [HGD90, Kah83, Mar94]. Intelligence [Inc92]. Interesting [For53]. Interface [BHY80, PH98]. Interior [FGS96, Wri92, Wri97]. Interior-Point [Wri97]. Interpolating [ADW84].

Interpolation [EGK90, KF63a, KF63b, Rei90, Smi65, Wer84]. INTerval [Rum99c, AH83, DK85, Fro01, Kah80, Kea96a, Kea96b, Man81, Moo66, Moo79, Nic77, Roh89b, Rum99a, Yoh79b, Yoh80]. Interval-Analysis [Nic77]. INTERVAL ARITHMETIC [Kea96a]. Interview [Eva76, Sev98]. Interviews [Alb91]. INTLAB [Rum99c]. Introduction [AH83, Bel70, Fox64, Hil74, Mir61, PJS92, Ste73b, SB80, Str93, TA32, SA73, Kea96b]. Introductory [COT89, IEE81]. Invariants [Tur29]. Inverse [BDG97, Coh75, HS81a, HP94b, Meu92, PS91, SL54, SS74]. Inverses [Asp59, Caf63, CM79, CS86, Dia78, Gau62, Gau75, Gau78, Hen62a, Ike79, NR76, NP84, PW70, Ste77a, Wed73, Zie86]. Inversion [BF88, BHH93, BR71a, CR93, CLP01, CH92b, Csa76, Doy57, GO94, Mit48, NT58, QQSvdG01, Wes68, Wil61]. Inverting [BE73b, GvN51, MH57, vNG47]. Involving [CR93, Gau83, Hig88b]. iPSC [HGD90]. iPSC/860 [HGD90]. Isaac [Whi97]. Isn’t [For70]. ISO [IEC94]. ISO/IEC [IEC94]. Isotropic [BG79]. Iterates [Hen62a]. Iteration [Dia78, DG94, HK93b, Kel65, Kni93, McK62, PS91, PBRZ99, Zaw91]. Iterations [Hig97c, Woź78]. Iterative [AR92, ADR92, Axe94, BNP88, BBC^{+}94, BR71b, BG67, Bjö67a, Bjö68, Bjö78, Bjö90a, Dax90, Des63, Gre97b, Gul94, HW76, Hig90g, Hig91a, HK93a, Hig97a, Hof89b, JW77, Kel95, Kel99, Kie81, MPW66, MP77, Mil88b, Mol67, Ske80, SS84, SS74, Tis01, Wac88, Wil55, You71, You89, GW66, GP90, Sch33].

J [Alb91, Eva76]. J. [Par90]. Jacobi [Mat95a]. James [Syl04, Ano87, Fox87]. Jim [Fox78]. John [Tau63]. Jordan [AM87, BR71a, DH89, GW76, Läu61, Loi69, Mal00, PW75]. Jordan-Elimination [Läu61]. Joseph [Syl04]. Journey [Dun90].

K7 [Rus98]. Kahan [Sev98, Nic70, Woe97].

Kahan-Babuskasche [Nic70]. Kaufman [JP93]. Kernel [SB95]. Kind [ABCR93]. kleinsten [Läu61]. Know [Gol91]. Known [Lin74]. Komma [SB53]. Kreiss [WT94]. Kronecker [HS81b, HPS83]. Krylov [HR92, Bec00, HS95, SvdVM00]. Kutta [But87].

L.S.D. [Tur84]. LABoratory [Rum99c, CK81, HGD90, Wil48a, Wil80, Woo58]. Lagrange [Wer84]. Laguerre’s [Par64]. L’analyse [Bru89]. LANCELOT [CGT92]. Lanczos [Coo94, FGP91, Gre89, GS92, Gre94, KW92]. Language [Fat82, HC88, IEC94, KKN^{+}92]. Languages [Kah01]. LAPACK [ABB^{+}99, CH01a, Hig97a, KP96b]. LAPACK-Style [KP96b]. LAPACK95 [BBD^{+}01]. Large [Ata40, CGT92, Ede91b, Ede93b, ES92, Gem86, GL81a, Hea84, HH89, MS02, MH57, Sau72, Woź77b, You71, Har48, Woź78]. Large-Scale [CGT92, ES92, Sau72, Har48]. Largest [KW92]. Last [Whi97, Ziv91]. Latent [Wil55]. Law [Ben38]. LCAS [Kah91]. Leading [Peñ96, Peñ98a, Tur82]. Leads [Gus97]. Least [Abd71, ADdR89, BI87, Bar88, BNP88, BV92a, BG67, Bjö67a, Bjö67b, Bjö68, Bjö78, Bjö87, Bjö90b, Bjö91, Bjö96, BG65, Cli73, CH98b, CH99a, CH99b, CH99c, DW93, Eld80, Far88a, Gau95, Gen73, GS90, Gol65, GDD94, Gu98, GW92, Gul94, Gul95, Hea84, HV97, Jan97, JO74, Jor68, Kar74, KW97, KLR98a, LH95, Van85, Lon67, MS02, Man81, PW70, Ple74, PR69, Rei91, Rei00, Sau78, vdS75, Ste77a, Sun96, Sun97, WKS95, Wei90, GW66]. Least-Change [DW93]. Least-Squares [ADdR89, Bjö78, Jan97, Jor68, Van85, Sun96]. Lecture [Kah73, Kah96b]. Lectures [CCF96, Wed34, ACM87]. Led [Uni92]. Legendre’schen [Hil94]. Leja [Rei90]. Length [CCG^{+}84, Tie70]. Leur [Bru89]. Level [COT89, DH92, DHSS94, Fat82, Hig90e, DCDH90b, DCDH90a]. Level-3 [DH92]. Level-Index [COT89]. Libraries [DDK92]. Library [CGG^{+}91, FHS78, GB91, IBM88, Inc89, KKL^{+}93]. Life [Far87]. Like [ABCR93, Bar91c, CR93, CS86, DDR95, Gau83, Kah97, Lu96, Rei91, BH92, Hig88b, Hig90h, Ypm83a, Ypm83b]. linéaires [Ben24]. Linear [Abd71, AL85, AL86, ADD89, AR92, AGL98, Ata40, Bas02, BLS91, BP67, BMvN46, BI87, BBC^{+}94, BSD60, BG76, BG67, Bjö67a, Bjö67b, Bjö68, Bjö87, Bjö91, BP94c, BDD^{+}02, Bol84, dBP77, BMPW66, BP71, BK77, Bun87, BDV89, BG65, But87, CM79, CF88, CI95, CG67, Che98, Cli73, CP76, ILA93, CR79, Cro41, Dax90, DK85, Dem92b, Dem97, DGK84, DDSvdV98, DCHH88b, DCDH90b, DCHH88a, DCDH90a, Don01, DRMN79, DR82, Dwy51, Ede91b, Ede93b, Ede94a, Eld80, ES92, Fad59, Far88a, For53, FM67, For67, FHW48, Fox64, GPS90, Gas83, GS90, GAKK93, Gol65, GV79, Gre97b, Gu98, GEJ95, GW92, Gul94, Gul95, Gus97, HW76, Ham87b, Han97, Hea84, HS87, Hig90g, Hig91a, HH92a, HH92b, Hig94b, Hig97a, Hig99a, Hof87, Jor68, Kah66, KW00]. Linear [Kar74, KW97, Kat76, Kel65, Kel95, KLR98b, KLR98a, Kie81, Kov76, Kup71, Lan70, Lan01, LHKK79, LYK89, Lu94, Lu95, MP74, MS02, Man81, Mat92, MN89, Men56, MP77, Mir61, Mol78, NP84, OP64, OPW65, Oet65, Peñ96, Ple74, PR69, Rei00, RG67, Rob80, Roh89a, Roh89b, Roh95, Sau72, SvdVM00, vdS70, vdS75, Smo95, Sny55, Ste77a, Str93, Sun96, Sun97, Tra66, Tre90, TB97, Ver49, WKS95, WWWS86, Wes68, Wil54a, WR71, Wil74a, Wil74b, Woź77b, Yip86, You71, Zie88, Ziȩ85, DR95, GP90, Woź78, Mol2b, vR01]. Linear-Algebra [Gus97]. Linear-Time [KW00]. Lineare [KS88]. linearer [Bau66]. Linearly [Syl51]. Lines [RT92a]. Liniowa [KS92]. LINPACK [Bye84, CCV82, DBMS79, Ste77c].

LINPACK-Style [Bye84]. Local [DW84]. Locality [Tol97]. Logarithm [CHKL01, Hig01, SA75, Tan90a, Tan90b]. Logarithmic [BB85b, BB85a]. Logic [Gla95]. Logiciels [Bru89]. Long [Qui94, QT90]. Long-Term [Qui94, QT90]. Look [BRB76]. Lookup [Tan91]. Loss [BP92, FG82, FT86]. Lösung [Bau66]. Louis [Bre96]. Low [Cap71, LRWW98]. Lower [EY36, Lás94, Var75]. LR [Rut58]. LR-Transformation [Rut58]. LSS [Mal01]. LU [CP98, DEL00, Mat84b, Pan00, Ste93b, Tol97]. LU-Decompositions [Mat84b]. Lyapunov [BS68, BN87, GLKH90, GL95b, Ham82, HK88, HS95, Hod92, Pen98b, Wac88].

M [Bod49, Bri92, CD86, Inc92]. M. [Far87]. MA27 [DR82]. MA47 [DR95]. MACHAR [Cod88a]. Machine [Ata40, Bow53, Cod88a, CT65, DGK84, Gil51, BHH93]. Machinery [Har48]. Machines [DT91, Sch88, Tor89]. Macintosh [Com94]. Made [Act96]. Magnetic [BSD60]. Magnetic-Tape [BSD60]. Mail [DG87]. majorants [Dah83]. Making [LD98, SW89]. Man [Sev98]. Manifestations [Cot74]. Manifold [AL86]. Manipulation [Sto77]. Mantissa [SS84]. Manual [BH95, CGG^{+}91, CW80, Com88, Inc89, Har46]. Many [EKS94, JSW83, Leh43]. Maple [CGG^{+}91, Map]. Marcel [Gil91]. Marching [BR77]. Mark [Blo48]. Market [BPR^{+}97]. Markov [Ham87b, IM94, O’C93, Ste97a]. Massively [GA96]. Math [Cau41, For70, Inc89, Matc]. Mathematica [Mata, SK93, Wol99]. Mathematical [Bab69, BP94a, DG87, Fik68, FMM77, FHS78, GB91, Gar61, Gla95, Hig98b, Ral91, Sca50, Var90, Ziv91, Syl04]. Mathematics [Boy89, Bri92, Dun90, Ede97, EG91, Ewi94, GKP94, Har67, Jos91, Kah83, Nel98, Tod77, YG72, Tur45]. Mathematik [Sto72]. MATLAB [Hig95c, HH00, SR97, Matb].

Matrices [ABO88, AOU87, And87, And90, Asp59, BDD^{+}97, BV92b, BZ98, Bau63b, Bau63a, Bau69, BR71a, BW93, Bec00, BP94a, Bir75, BG79, BV87, BLP90, BE73b, Bun71b, Bun74, Bun82, Bus68, Bye88, CR93, CS86, Cha93, Che98, Cho68, Chu99, Chu00, Cry73, Cry76, CR72, Dan76, DH00, Dav79, Dem89a, Dem90, DK90, Dem99, Dix83, DGK84, DER86, DGR^{+}91, Ede88, Ede93a, EGL^{+}87, Fal01, FH55, FV62, Fle76, FS55, Fra58, FNP82, Gan59a, Gan59b, Gau53, Gau62, Gau75, Gau78, Gau83, Gem86, GIK02, GO94, GS83, GvN51, GK69, GL81b, Han97, HS81a, Hen62a, Hig87c, Hig92b, Hig98a, Hou64, Hym57, Ike79, IK00, JP93, JP94, Kit95, KS00, Kos92, Lad76, Leh96, Lem75, Loi69, Mal00, Mat92, Mat96, Mat84a, Mat84b, MS74, Men56, Meu92, MH57]. Matrices [Neu86, vNG47, NP84, PBRZ99, Pan01, Pol87, Rei91, RT92b, Rei87, Roh89a, Rum98, Rum99b, Rut68, SW98, Sha82, Sha85, Sha91, vdS69, Ste80, Ste97d, Ste97c, SW62, SB95, Syl84, Tao84a, TM62, TT94, Tre92b, Tur29, TA32, Tyr91, Van95, Var79, VT98, WSW72, Wal93, Wat91, Wed34, Wil55, Wil65b, Win71, Zie86, Bar86]. Matrix [Aas71, BDDD97, Bai88b, BF88, BK79, BK80, BHH93, BS68, BS72, Bar91c, Bel70, BR71b, BM94, Bha97, BL80, BP94b, BMSV92, Bod49, BPR^{+}97, Bre70a, Bre70b, BKP76, Bye87, Bye88, Caf63, CL83a, CL83b, CHKL01, CH01a, CH01b, Cho83, Chu87, CMSW79, CLP01, Coh75, CW87, CH92b, Csa76, Dat88, Dem83, DK87, DM89, Dhi98, DEL00, DHSS94, DS94, Doy57, DRMN79, DGL89, DGL92, EY36, Ede91a, EKS94, EM95a, EM95b, ER74, Fis74, FS80, Fri91, GBDM77, GWL^{+}92, GLAM92, Giv54, GV74, GV96, GKL95, HS81b, Hig86b, Hig87a, Hig88a, Hig88c, Hig89b, Hig90a, Hig90e, Hig90d, Hig92a, Hig92b, Hig93b, Hig94a, HK95, Hig95c, Hig97c, Hig99b, HT00, HK00, HT01, Hig01, HS95, HJ85, HJ91, Hot43, Ikr98, Inc85, Jen82, KP96a, KP96b, Kap99, Kau83, KLP92]. Matrix [KL94, Ker90, Kni95, LPS92, Lan70, Liu87, Van79, MN79, MN99, Mat95b, MAV^{+}97, MP77, MW80, Mil88b, Mit48, Mol2a, MV78, NT58, New62, O’L80, Pai73, Pan78, Pan84, PS91, PY01, Par64, Par65, Pry85, QQSvdG01, SL54, SBD^{+}76, Ste73b, Ste77b, SS90, Ste98, Ste01, Sto62, Str86, Sun92c, TG81, Tau49, Tau88, Tod54, Tod61, Tur48, Var75, Wac88, WT94, Wes68, Wil61, ZZ95, Zie88, Ziȩ85, ZWS86, Dem92a, Hou58, CV86, Dah83, BPR^{+}97, Spi76, Sch33]. Matrix-Matrix [DS94]. Matrix-Multiplikation [Spi76]. Matrix-Sign [KLP92]. Max [Bun71b]. Max-Norm [Bun71b]. Maximum [Boh77, Tao84b, Tor65]. Mead [LRWW98, McK98]. Mean [BR71b]. Meaningless [Fox71]. Measure [Pry84, Ziv82]. Measurements [Hun97]. Measures [Hig91b, Rhe76]. Measuring [Bye88]. Mechanical [Inc92]. Mechanically [MLK98, Rus98]. Memoir [Far87]. Memory [BDD^{+}97, KP92, Kre76]. Method [Abd71, Ari00, Bar91c, BR71a, Bjö87, Bjö92, Boy74, Bun71a, Bye88, Cap71, Cli73, CGM85, CH99a, Cro41, EGK90, Fos91, FM91, Gen69, Gol62, GNV79, GS76, Han97, Hig92b, HP94b, Hig97b, Inc85, LRWW98, Lan66, Lau79, Van85, Lyn64, MP74, McK98, MH57, NM65, Oli77, Ort63, Pai73, Par64, Pug92, RT92a, SS74, Tao84a, Tis01, VA86, War76, Woź77b, Ypm95]. méthode [Ben24]. Methods [Act70, AR92, Axe94, BNP88, BV92b, Bar89, BBC^{+}94, Ber91, Bjö90b, Bjö96, Bol84, BT92, BE73b, BP71, BK77, But87, BGN70, CN96, CH94, CH99c, CH92b, DBA74, DR84, DES82, DS83, DW84, DT91, DW93, Dub83, DER86, Ept80, Fad59, FGP91, Fik67, Fle87, FGS96, For67, FMM77, Fra58, GMSW91, Gol65, Gov00, Gre97a, Gre97b, HW76, Ham73, Hea84, Hen62b, Hen63, Hig90g, Hig91a, HK93a, HS95, Hot43, HR92, IK66, KW89, KMN89, Kau83, Kel95, Kel99, LYK89, LTT00, Lin74, Mat95a, MD64, Met77, MP77, Moo79, Pow70, Qui94, vR01, Rum01, Swa72, Tao84b, Tra66, TT87, Wil55, Wil61, Wri92, Wri96, Wri97, You89, Ypm83a, Ypm83b]. Metric [FH55]. MGS [Ple74]. Might [Kah90]. Million [QTD91]. Minimal [Bau63a]. Minimization [NM65, Pow75]. Minimum [CV86]. Minor [Syl51]. minorants [Dah83]. Miscalculate [Mar94]. Miscalculating [Kah97]. Misdemeanours [Sto91]. Missile [Uni92]. Mixed [GK93, LDB^{+}00]. Model [AB79, Bro81, CB90, Ham87b, Hen80, Rob80, Sch84, Tie70, DCHH88b, DCDH90b]. Models [Hen66, HS66, MN89]. Modern [BL77, Wil71a]. Modification [FP74, Pug92, Yip86]. Modifications [Fos91, Ham74]. Modified [BP92, CH98a, Che98, FGM95, Oli77, RSD77]. Modifying [GW92, Ste74b]. Modular [BBB89]. Module [Kea96a]. Modulo [Per02]. moindres [Ben24]. Moments [Ede91a]. Monde [LM66]. Monitoring [Bus71, ER74, Bar86]. Monotone [FB91]. Monotonic [BSW61]. Monthly [Ewi94]. Moore [SS74]. Morrison [Yip86]. Most [Boy94]. MP [Bre78b, Car91]. Much [Kah87a]. Muir [Bel31]. Multi [Tor89]. Multi-Directional [Tor89]. Multicomputer [Ede94a]. Multidigit [KO63]. Multidirectional [Tor91]. Multiple [BKP93, BHY80, Bre78b, Bre78a, HH92b, Smi91, Sun96, Ypm83b, Smi98]. Multiple-Precision [Bre78b, Bre78a, Smi91, Smi98]. Multiplication [Bai88b, BL80, BMSV92, Bre70a, Bre70b, CW87, DS94, GFB85, Hig90e, Kap99, KO63, Kni95, LPS92, Mat95b, Pan78, Pan84, Per02, Rus98, Ura68, Win71]. Multiplications [GA96, Hig92b, Lad76, PS73]. Multiplicative [Pry85]. Multiplicativity [GS83]. Multiplikation [Spi76]. Multiply [DHSS94]. Multiplying [Hig92b, Lad76, Vit68]. Multiprecision [Bai95, KFC^{+}01, Bai93]. Multishift [DG94]. Multistep [Qui94]. Multivariate [Mui82]. Mysteries [Tre85].

nach [Läu61]. National

[HGD90, CK81, Wil48a, Wil80, Woo58]. Natural [New81]. Nature [LK76]. Near [Rum99b, Mat24]. Nearest [Dem87a, Hig88a, Dem92a]. Nearly [Gri85, Sha91]. Nearness [Hig89b]. Necessary [PS73, Sha82]. Need [Kah81]. Needle [Kah97, WT94]. Needle-Like [Kah97]. Negative [Ham82]. Nelder [LRWW98, McK98]. Netlib [BDG^{+}94]. neue [Spi76]. Neumann [Gol72, Ple76, Tau63]. Neville [AGP97, GP92]. Newton [PBRZ99, Rei90, Tis01, BF88, DES82, DW84, EGK90, FGM95, KF63a, KF63b, Lan66, PS91, Wer84, Whi97, Ypm83a, Ypm83b, Ypm95]. Newton-like [Ypm83a, Ypm83b]. Nineteen [MV78]. Ninth [Gra92]. nombre [Ben24]. Non [BDDD97, Ham82, Hen62a, Jos91]. Non-European [Jos91]. Non-Hermitian [BDDD97]. Non-Negative [Ham82]. Non-Normal [Hen62a]. Noncommutative [Gil91, Lad76]. Nonlinear [Bar74, CGT92, DS83, GMSW86, GS76, Gri85, Kel95, Rhe76, Woź77a]. Nonnegative [BP94a, Fal01, FS80]. Nonnormality [Loi69]. Nonscalar [PS73]. Nonsingular [Hea77]. Nonsingularity [PR93, Roh90]. Nonstandard [Cra79, Yoh79a]. Nonstationary [McK98]. Nonsymmetric [BDM93, BDG97, BDD^{+}97, LZ92, Hou58]. Norm [BZ98, Bar91c, Blu78, Bun71b, CH01a, CH01b, Gau75, Hig90d, HT00, MAV^{+}97, SW98, Tao84a, Tao84b, Hig88c, Hig92a].

Normal [Fos91, Hen62a, Lás94]. normales [Ben24]. Normed [SW62]. Norms [BSW61, Boy74, GS83, GKL95, SS62, Sto62, Zie88, Rum02]. Notation [Knu92, SZ80]. Note [BV92a, Ben24, Bun82, CS86, CG85, Cle55, Coh74, Dem01, Des63, For59, GW66, Ham74, Hor79, Ikr98, JSW83, Kie87, Van79, Mal00, Mol90b, Møl65a, New81, QQSvdG01, Rei71, Smo95, Ste85, Ste92, Sun95a, Tsa75, Wic92, Yip86, ZZ95, Bar86, Edi91, van77]. Notes [FHW48, Hig99a, Kah73, Kah96b, Knu92]. Nothing [Kah87a]. NP [PR93, Roh95]. NP-Hard [PR93]. NP-Hardness [Roh95]. NPSOL [GMSW86]. Null [CH99a]. Number [Bar89, Bec00, Bre73, CH94, CMSW79, CR83, Dem83, Dhi98, Ede92b, GL81b, GEJ95, Ham87a, Hig86b, Hig87c, KC73, MK85, MAV^{+}97, PS73, SA75, WWWS86, ZWS86]. Numbers [BDM93, Ben38, Bir75, Cli90, Coo81, Dem87a, Dix83, Ede88, GK93, Hig95a, KO63, Lem75, LO82, Loi69, MI81, Mol95, New81, O’L80, Per02, Roh89a, Ske89, vdS69, SW90, Tod68, Tur36, VT98, Zie88]. Numeric [Hau96, Mat70, Mol90b]. Numerical [Act70, Bab69, Bjö96, BCP^{+}97, Bla64, BRR85, Bol84, BT92, Bun87, But87, Bye85, Cod73, CdB80, CH90, DBA74, DR84, Dem84, DDK92, Dem92b, Dem93, DL94, Dem97, DS83, DDSvdV98, Dou59, DGRS95, Dub83, Ede93b, Ede94a, Eir93, Ept80, For78, Fox64, Gas83, Giv54, GAKK93, GvN51, Gol77, Gol65, Gov00, Ham82, Ham87b, HHKR93, Ham73, Hea84, Hen64, HS87, Hig94b, Hig99a, HK00, Hig02, Hil74, Hou64, Hul79, Hul82, HAC^{+}85, IK66, JW77, JT74, dJ77, Kah66, Kah96a, KMN89, KW00, LM66, MD64, MV99, Mil75a, Mil84, Mol78, vNG47, NW99, Ort72, Ove01, PY01, Par90, Pen98b, Pow70, PTVF92, vR01, Row90, Ruh83, Sca50, Sha94, Ske79, Ske80, Ske89, SK93, Sma90, SS74, SB80, Stu80, Tod77, Tre90, Tre92a, TB97, Tsa83, Vit69, Wen66].

Numerical [Wes68, vW66, Wil71b, Wil74b, Woź77b, Woź77a, YG72, Zaw91, SA73, Bus71, Pri92, Thi88, Sun92a]. Numerically [DH00, Fos91, GM73]. Numerics [Bjö94, Com88, Com94]. Numériques [PV74, Bru89]. Numerische [KS88, Sto72]. Numeryczna [KS92].

Oak [HGD90]. Obliczeń [KS92]. Observations [Gau95]. ODE [SR97, Tuc02]. ODEs [Eir93, HT93]. Odor [Dem91]. Off [Abd71, AB79, Cap71, Dem93, Des63, Dia78, Dou59, For59, Gol62, GF75, Lyn64, Pry81, Qui94, Qui83, Tur48, VA86, Ziv82, BC86, CB90, Bod49]. Old [Sev98]. Once [Wri96]. One [BBB89, EY36, Gau95, Hig88c, Ove01, PJS92, SB92, GP90]. One-norm [Hig88c]. Ones [Ziv95]. Only [JW85]. onto [AL86]. Open [Dem92b]. Operation [Ham87a, Har46]. Operations [AvL81, CHN99, VCV01a]. Operator [BR97, BN87, Goo79, HS81b]. Operators [Kat76]. Optimal [Bau63a, Bur97, KW97, Mal01, MS02, MS74, Pan78, Rum02, Sha85, Sha91, Str69, Stu82, Stu85a, Sun96, Sun97, SS97, WKS95, Wat91]. Optimale [SB53]. Optimality [Har72]. Optimally [Bau63b, Bau69, Bus68, Sha82]. Optimisation [Ber91]. Optimization [CGT92, DS83, Fle87, FGS96, GMW81, GMPS92, Hig93b, Kel99, NW99, Pow70, Pow98, Swa74, WPD01, Wri92]. Optimum [Vit68]. Options [Kah80]. Orbit [QT90, QTD91]. Orbital [Qui94]. Orbits [CKP95]. Order [BMvN46, EM95a, EGK90, GvN51, Gol62, Mit48, vNG47]. Ordering [RS88]. Orderings [EGL^{+}87]. Ordinary [But87, CKP95, Ept80, HW91, Hen62b, Sha94]. Ordinateurs [Mul89, PV74]. Organization [Bow53, PH98]. Origin [Buc81, CT93, Tro84b]. Origins [Ran75]. Orthogonal [AOU87, BZ98, BP94c, CR93, Gau83, HSS99, Hig88b, HV97, Rat82, Smi65, Ste80, SB95, Wil62]. Orthogonality [BP92, ES93]. Orthogonalization [Bjö67b, Bjö94, Hof89b, Ric66a, Ruh83]. Orthogonalizing [Mat96]. Other [CD86, Hun97]. Over-Relaxation [Lyn64]. Overflow [FT86, MI81]. Overflow/Underflow [MI81]. Overflow/Underflow-Free [MI81]. Overview [You89]. Owner [BH95].

P [Vav89]. P-Complete [Vav89]. P754 [IEE81]. Package [BHY80, Bre78b, Bre78a, Cod93, CGT92, GWL^{+}92, GMSW86, KL85, Smi91, Yoh79b, NAG86]. Padé [Hig01]. PADRE2 [Kub91]. Paging [Mol2a]. Pair [KP96a]. Pairs [KP96b]. Pairwise [Bar91a, Sor85]. Paper [CT93]. Papers [CD86, Ran75, Syl04]. Parallel [APS93, BDG97, BDD^{+}97, CH01b, CLP01, Csa76, Dem92c, DDR95, DT91, Ede93b, EGK90, GPS90, GL91, GA96, HP94a, HP94b, Hig95b, HJ88, LO82, Mat95b, MRW01, QQSvdG01, Rum99a, SB77, Sch88, Tor89]. Parallelism [Dem93]. Parallelizable [BF88]. Parameter [Bar74]. Parameterization [For78]. Parameters [Cod88a, Sch84]. Paranoia [Kar85]. Part [BM94, Gau95, IEC94, Mat92, PJS92, Gau95, Stu85b, Stu85c]. Partial [Bun74, Dax82, FGM95, Fos94, Hig97b, Liu87, Mac89, Ske81, Tol97, Vav89, Wri93, van77]. Particular [HW76, Ziȩ85]. Partitioned [HP94b]. Parts [Alb91, Hig98a]. Pascal [Syl50, KKN^{+}92, Coh75, Gol72, HHKR93]. PASCAL-XSC [KKN^{+}92, HHKR93]. Pascal’s [Gra92]. Patriot [Uni92]. Pattern [Tor97]. Paul [EG91, Alb91]. Peacock [Jos91]. Peculiar [Rai69]. Pencils [Dem83, DK87]. Penguin [Nel98]. Penrose [SS74]. Pentium [Ede97, Mar94, SB94]. People [Lav80]. Pereyra [Hig87b]. Perfidious [Wil84]. Performance [BBG92, DDK92, DDSvdV98, Don01, JP94, RSD77]. Permutation [HS81b]. Perturbation [Bar91b, Bar94, Bjö91, CPS96, CPS97, CP01, DOV94, Eld80, FM91, Gu98, HH92b, Hig93c, Hig94b, Kåg94, KW97, Kat76, Mal01, MS02, O’C93, Ste73a, Ste77b, Ste77a, SS90, Ste90, Ste93b, Ste97b, Stu81, Sun91, Sun92c, Sun92d, Sun95b, Sun96, Sun97, SS97, WKS95, Wed73, Wei90, Zha93a]. Perturbations [Rum98, Sun95a]. Perturbed [Gre89]. Phenomenon [TT87]. Phys [Cau41]. Physical [CK81, Wil48a, Wil80, Woo58]. Physics [Rut47, Rut52]. Pi [BBB89]. Pilot [CK81, Wil54a, Wil54b, Wil80]. Pipeline [DGK84]. Pitfalls [For70, SA56]. Pivot [Coh74, Cry68, Ste74b, Tor65]. Pivoting [Bjö92, Bun71a, Bun74, CG85, DK77, Dax82, Ede92a, EM95a, For60, Fos94, Fos97, Gou91, HH89, Hig97b, Hig00, LD98, Liu87, Mac89, MRW01, Peñ96, Peñ98a, PW75, PN00, Ske81, vdS70, Sor85, Tha61, Tol97, Vav89, Wri93, YC97, van77]. Places [Leh43]. Plane [AP94, Ham74, Wil63a]. Plane-Rotations [Wil63a]. Planetary [QT90]. Please [Tur91]. Pocket [Smi75]. Point [And99, AL86, BSBF89, Bar81a, Bar81b, BB85b, BB85a, Bar89, BGM^{+}94, Bre73, Bro81, Car59, Car91, CO84, Cli90, Cod73, CCG^{+}84, Cod88b, CC93, CHN99, Dek71, Dem89b, Dem92c, DDR95, Dem01, Esp95, Far88b, Fat82, FG82, FT86, GB91, GM74, Gol91, GFB85, IEE81, Hau96, HLB01, Hig91b, Hig93a, Hou81, JSW83, KP79, Kah80, Kah81, Kah87b, Kah96b, Kar85, KC73, IEC94, LO82, Lin74, Lin75, Lin81, Lin70, Lon67, Mal71, Mal72, MI81, Mat70, McK98, Mol67, Møl65b, MLK98, Ove01, Pla88, Pri91, Pri92, RK75, RS88, Rus98, Sch81, Sch84, Sev98, SB94, She97, Smi91, SW90, Ste74a, Stu82, Swe65, Tan89, Tan90b, Tan92, Tie70, Tis01, Ura68, VCV01a, VCV01b, Vit68, Wal90, Wic89, Wic92, Wil60, Wil63a, Wri97, NAG86, IEE85, IEE87, Boh77]. Points [Rei90]. Poisson’s [BGN70]. Polar [Hig86a, Hig94a, HP94a]. Polynomial [Ada67, BP94b, EM95b, Fik67, New74, Oli79, Olv86, PW71, Rei91, SR80, Smi65, Ste71, Wer84, Wil84]. Polynomials [CR93, Gau83, Hig88b, HT01, Kit95, Van79, Pan01, PS73, TT94, Tra66, Knu62]. Polynoms [Hil94]. PORT [FHS78]. Portable [Blu78, Cod93, DDK92, DHSS94, Yoh79b, Yoh80]. Posed [Dem87a, DK88]. Positive [And87, BR71a, Bec00, dBP77, Cry73, Cry76, CV86, Doy57, GL81a, GV79, Hig88a, Hig98a, Ker90, Mak91, MPW66, Mat92]. Positive-Definite [Doy57]. Positivity [GP92, Kar68]. Possible [Sto62]. posteriori} [OW82]. Power [Boy74, KW92]. PowerPC [Com94]. Powers [FS80, Gau53, HK95, MP77]. Practical [BCP^{+}97, EGK90, Fle87, GMW81, HW76, Kap99, LPS92, PW71, vR01]. Practice [Cod88b, Fos94, KM81]. PRECISE [KFC^{+}01]. Precision [BKP93, Bre73, BHY80, Bre78b, Bre78a, CCF96, CHH83, Dek71, Dem87b, GS92, Gre94, HLB01, Hig91b, HK93b, HK95, Hul82, JW85, Kah87b, KM97, Kie81, LM01, LDB^{+}00, Lin81, MK85, Mil88a, Møl65a, Møl65b, Pri91, She97, Smi91, Smi98, VCV01a, VCV01b]. Precompiler [Cra79, Kub91]. Preconditioned [HS95, vdV90]. Preconditioners [GMPS92]. Preconditioning [CGM85]. Predicates [KFC^{+}01, She97]. Predicting [GS92]. Prefix [Dem92c, Mat95b]. Prescribed [CL83a, Chu99, Chu00, KS00]. Presence [Neu85, Not93, vdV90]. Present [Woo58]. Prespecified [DK85]. Preventing [Act96]. Primal [Wri97]. Primal-Dual [Wri97]. Print [SW90]. Priori [Wil68]. Probabilistic [Bar81b, BB85b, CB90, Hen66, HS66, Kah96a, YC97]. Problem [Boy94, Bur97, CB90, CH99a, CH99c, Dem87a, Eld80, GS90, GNV79, Gul95, JW86, Kar74, KW97, Lu95, Mal01, Mol90a, Ort63, Par64, Par98, PW70, Ple76, PC83, Rai76, Sau78, Sun95a, Sun96, Sun97, TM01, Tuc02, WKS95, WT94, Wei90, Wil65a, Wil70, Zha93b, Uni92]. Problem-Solving [JW86]. Problems [Abd71, ADdR89, BDDD97, BR77, BI87, Bar88, BNP88, BV92a, Bjö67b, Bjö87, Bjö96, Che98, Cli73, CH98b, CH99b, DK88, Dem92b, DGL89, Gol65, GDD94, Gu98, HHKR93, Hea84, Hig89b, Hig90c, Hof89a, JW85, Kau83, Kni93, LH95, LZ92, Van85, MS02, Man81, Mol78, Par65, PW71, PR69, Qui83, Rei00, Roh95, Rut58, vdS75, Ste73a, Ste77a, Tis01, Tod61, Tra66, Var90, Wri93]. Procedures [Jor68]. Proceedings [Har48]. Procesach [Wil67]. Process [Gil51]. Processes [BR71b, Dax90, Des63, Tur48, Wil63b, Wil68, Bod49]. Processor [BKP93, Rus98, SB94]. Product [GA96, HPS83, Neu85, Win68]. Products [BV87, GF75, HS81b, Mat96, SV89]. Profit [GS66]. Program [BBG92, Blu78, Hor79, MLK98, Pax91]. Programming [CK81, ES92, GMSW86, GMSW91, HJ88, HC88, Kah01, MD64, Sau72, Wal90, Wol64]. Programs [Bai93, DCHH88b, DCDH90b, HHKR93, Hau96, Lon67, NT58]. Progress [Wil48b]. Progressions [CW87]. Project [WPD01]. Projection [AL86]. Projections [Ste77a]. Proof [MLK98, Rus98, Yuv78]. Propagation [CB90, Hen63, Hen66, Hen80, HS66, Oli79]. Properties [AN63, Cry76, GM74, LRWW98, MN79, Mal72, Men56, Pla88, Pri92, SS74]. Proposal [Str86, Tur45]. Proposed [CCG^{+}84, Fat82, IEE81, Hou81, KP79, Kah80]. Provably [KW00]. provenant [Ben24]. Proving [Fro01]. Pseudo [PW70, RT92b, Ste77a, Wed73]. Pseudo-Eigenvalues [RT92b]. Pseudo-Inverses [PW70, Ste77a, Wed73]. Pseudospectra [ETxx, HT00, TT94, Tre92b, Tre99]. Pseudozeros [TT94]. Publication [CT93]. Punch [Ver49]. Pure [Har67, Bri92].

Puzzles [Gar61]. Pythagorean [Dub83, Inc85, MM83a].

QR [ABD92, CH92a, CPS97, CP01, CH98b, DGKS76, DG94, GW92, Hig00, Pai90, Rei91, SB92, Ste93b, Sun95b, War76]. QR-Decomposition [GW92]. Quad [HLB01]. Quad-Double [HLB01]. Quadraten [Läu61]. Quadratic [Doy57, For69a, For69b, GMSW91, HK00, Lic87, Roh95, Syl51, Tao84b, TM01]. Quadrature [LK76]. Quantitative [HP96]. Quartically [Bai88a]. Quasi [BW93, DW84, Møl65a, Møl65b, Van95]. Quasi-Definite [Van95]. Quasi-Newton [DW84]. Quasi-Toeplitz [BW93]. Quasidefinite [GSS96]. Queuing [Kau83].

Radii [KL88]. Radius [Gem86]. Radix [CCG^{+}84, IEE87]. Radix- [CCG^{+}84]. Radix-Independent [IEE87]. Ramanujan [BBB89]. Random [AOU87, And90, Ede88, Ede91a, Ede93a, EKS94, Gem86, Hou89, KW92, Sch89, Ste80, VT98]. Range [VCV01a, VCV01b]. Rank [BV92b, CH92a, EY36, Man81, Pan00, SB92, Yip86]. Rank- [Yip86]. Rank-One [SB92]. Rank-Revealing [Pan00]. Raphson [Lan66, Ypm95]. Rapid [Jan97, Per02]. Rate [Not93]. Rational [MK85]. Rationals [ADW84]. Ratios [Sto62]. Read [Cli90]. Readings [BN71]. Real [Act96, BD93b, Bec00, BMPW66, CL83a, CL83b, Cro41, EKS94, Giv54, Ham87a, Hig87a, Hig88c, Hig92b, Hig98a, JP93, Kah86, Tur91]. Realistic [Bro81]. Reanalysis [Jan97]. Reasonably [Yoh79b]. Recapture [BP92]. Rechenanlagen [SB53]. Rechengenauigkeit [SB53]. Recipes [PTVF92]. Recommendations [CGL83]. Records [Hun97]. Rectangular [Kni95]. Recurrences [Gre89]. Recursion [Gus97]. Recursive [Chu00, DEL00, FGP91]. Recursively [Gre97a]. Reducing [Kah65, Ros65, Wol64]. Reduction [Aas71, AM87, AM94, BG94, Fer95, Rob80, Tor65]. Reference [CGG^{+}91, HW76, IBM88, Inc89, KKN^{+}92, Tol97]. Refined [Mei83]. Refinement [BG67, Bjö67a, Bjö68, Bjö78, Bjö90a, GL95a, Gul94, Hig90g, Hig91a, Hig97a, JW77, Kie81, MPW66, Mol67, Ske80, SS84, Tis01, GW66, GP90]. Reflections [Par71]. Reflectors [SP88]. Refutation [Kah91]. Regression [BRB76, Hig90c]. Regular [KP96a, KP96b, Ple76]. Rehabilitation [DH89]. Related [Ste97c]. Relation [Hig94a, PC83, Syl51]. Relative [Dem87b, DGE^{+}99, GS90, GFB85, LS80, LPW83, Pry84, Ziv82]. Relatively [PD00]. Relaxation [For51, Lyn64]. Release [BDDD97, CGT92, DGL92, IBM88, NAG86]. Reliability [Dem84, Fra92, Hor79, MV99]. Reliable [Bjö90a, CH90, Dhi98, KFC^{+}01]. Remainder [VCV01a]. Remark [Ikr98, Tau49, Ura68]. Remarks [Bau69, Kah65, Sma90, Zie88]. Reminiscences [Sev98]. Reorthogonalization [DGKS76]. Replacing [MM83a]. Report [CD86, ILA93, Wil48b, Zie86]. Representation [BV87, Buc59, Ham87a, MI81, Pug92, SV89, SB95]. Representations [PD00]. Reprint [For59]. Require [Pow75]. Requirements [Kre76]. Research [Mol78, Ste77c]. Residual [GL95b, Gre97a, Ske81, Stu85b]. résolution [Ben24]. Resource [BPR^{+}97]. Resources [BDG^{+}94, Kea96b]. Respectable [Wri96]. Result [Yuv78]. Résultats [PV74]. Results [Roh95]. Reveal [GM74, Mal72]. Revealing [CH92a, Pan00]. Revelations [Tur84]. Review [Bel31, Bod49, CK84, HS81b, Meu92]. Revised [Tho42]. Revisited [Wil86]. Revisiting [Fri91]. Reziproken [Sch33]. Riccati [AL84, Bye85, Bye87, GL91, GL95a, KH90, KLP92, Lau79, Rob80]. Richardson [Gol62, Zaw91]. Ridge [HGD90]. Riesz [Gil91]. Right [HH92b, OP64, Sun96]. Right-Hand [HH92b, OP64, Sun96]. Rigorous [CKP95, Tuc02]. Robust [And91, PD00, PR93, She97]. Robustness [Hof89a]. Role [CLP01]. Roll [BKP93]. Roll-Up [BKP93]. Rook [Fos97, PN00]. Room [WSW72]. Root [Ada67, Hig97c, KM97, Rus98, VCV01a]. Roots [EM95b, Gen73, Hig87a, Jos91, KFC^{+}01, MM83a, Tau49, Wil55]. Rotation [Ham74]. Rotations [AP94, BI87, Rat82, Wil63a]. Roulette [Ede93a]. Round [Abd71, AB79, BC86, Cap71, CB90, Des63, Dou59, Gol62, GF75, Lyn64, Pry81, Qui94, VA86, Ziv82]. Round-Off [AB79, Cap71, Des63, Dou59, Gol62, GF75, Lyn64, Pry81, Qui94, VA86, Ziv82, BC86, CB90]. Rounded [Dia78, Gay90, LMT98, Ziv91]. Rounding [Bau74, Boh75, For59, Hen80, Hun97, Iri91, Kie87, Lan01, LM01, LO82, Lin75, Lin76, Mül83, Neu85, Not93, Oli79, Qui83, Roh90, SZ80, SvdVM00, Stu80, Sun92d, Tur48, Ura68, Wil63b, Ypm83a, vdV90, Bod49, Mol90b]. Rounding-Error [Kie87, Sun92d]. Rounding-Off [For59, Tur48, Bod49]. Roundoff [Bar81a, BB85a, BR71b, Hen66, HS66, LS78, LS80, Mil75b, Mil76, MW80, MS78b, MS78a, QT90, Ram71, TZ01, Tie70, Ura68, Woź78, Woź80]. Roundoff-Error [Woź80, Woź78]. Routines [Fik67, GBDM77, JW86, LK76, SBD^{+}76]. Row [CH99c, Gra92]. Row-Wise [CH99c]. Rows [Peñ98a]. Rule [Mol74, Ove01, Ura68]. Rundungsfehler [Wil69]. Rundungsfehleranalyse [Neu74]. Runge [But87].

S [Kah73, vdV90]. S.O.R. [HW76]. Saddles [SSL93]. Safety [Ske89]. Sample [Bar91a, CGL83, GKL95, KL94]. Sand [Kah83]. Satisfactory [For69b]. Satisfy [Asp59]. Saudi [Uni92]. Scalable [Ede94a, LD98]. ScaLAPACK [BCC^{+}97]. Scale [CGT92, ES92, Sau72, Har48]. Scaled [BI87, Bau63b, Bau69, Bus68, Ede92b, Sha82]. Scaling [AP94, Bau63a, CR72, ES93, GV74, Rei00, Rum02, Sha85, Sha91, Ske79, Wat91]. SCALP [FM91]. Scheme [Mül83]. Schemes [Fis74, Oli79]. Schmidt [Abd71, Bjö67b, BP92, Bjö94, DGKS76, Hof89b, JP91, Ric66a, Ruh83]. Schur [BD93b, Cot74, GNV79, KW89, Lau79]. Science [MD64, vW66]. Sciences [BP94a, Hig98b]. Scientific [Act96, BRK^{+}87, BC86, Fox71, IBM88, Inc89, Kah90, KKL^{+}93, KM83, PTVF92, Smi75, Var90]. Scientist [Gol91]. Scientists [Ham73]. Scorned [Wri96]. Search [Ber91, DT91, Hig93b, LTT00, Pow98, Swa72, Swa74, Tor89, Tor91, Tor97, Wri96]. Searching [GA96]. Secant [DW93, GS76]. Second [ABCR93, Gol62, Kub91, Ort72]. Second-Kind [ABCR93]. Segments [SL54, Tau49, Tod54]. Selected [Ran75]. Self [Han97, Rum01]. Self-Adjoint [Han97]. Self-Validating [Rum01]. Semi [Hig90a]. Semi-definite [Hig90a]. Semidefinite [Hig88a, Kel65, Ste97d]. Semidefiniteness [Ker90, Mak91]. Seminormal [Bjö87]. Seminumerical [Knu98]. Sensitivity [Bar91c, BH92, CI95, CP98, GLKH90, GL95a, HK88, KH90]. Separation [KP96b, Var79]. Sequence [Har46]. Series [And71, Boy94, Cle55, CT65, New73, Oli77, Smi65, Hew93]. Services [BDG^{+}94]. Set [CR83, DCHH88b, DCHH88a, DCDH90a, DR82, Har80, Oet65, Sny55, BDD^{+}02, DCDH90b]. Sets [DRMN79, DR82, WSW72]. Severance [Sev98]. Severe [FT86]. Shadowing [CKP95]. Shadows [SSL93]. Shared [KP92]. Sharply [OPW65]. Sheet [BKP93]. Sherman [Yip86]. Shifting [Ste71]. Shifts [DG94]. Short [Cro41]. Shorthand [SZ80]. Should [Gol91]. Sides [HH92b, OP64, Sun96]. Sign [Bye87, Hig94a, Kah87a, KLP92, PY01, Rob80, SA75]. Sign/Logarithm [SA75]. Significance [FG82, FT86, Mei69, Met77]. Significant [Boy94, Tur82]. Signs [Dan76]. SIMD [BMSV92]. Similarity [Rat82]. Simple [Bro81, Gri85, GEJ95, Tsa83, Yuv78]. Simplest [Mol90a]. Simplex [GM73, LRWW98, McK98, NM65, RSD77]. Simultaneous [BSD60, FHW48, Sny55, Ver49]. Single [JW85]. Singular [And90, BD93a, BN87, CI94, Chu99, Chu00, Dax90, DK90, Dem99, DGE^{+}99, Gri85, GDD94, HK93b, Jen82, Kar74, Kel65, Kit95, KS00, MP77, MM83b, Ste93a, Var75, Dem92a]. Singularities [Gri85]. Singularity [Rum99b, Rum02]. Sir [Bel31]. Size [Coh74, Cry68, Ske81]. Sizing [DW93]. Skew [Bun82]. Skew-Symmetric [Bun82]. Slash [MK85]. Slightly [Gre89]. Sluggish [Ste97a]. Smale’s [Tuc02]. Small [GKL95, KW00, KL94, Peñ96]. Small-Sample [GKL95, KL94]. Smallest [Ede91a, Var75]. Snippet [Fox78]. Software [AL84, CW80, Cod82, Dem84, DG87, Don01, For78, FJ98, GWL^{+}92, Hul79, JW86, KP96a, KP96b, KMN89, Lin81, MV99, Mil75b, MW80, Mil84, MS78a, PH98, WPD01, Yoh79b, Yoh80, ZWS86, LPW83, MS78b, Uni92]. Solution [Abd71, ABCR93, APS93, AM94, AR92, Ata40, Axe94, BLS91, BMvN46, BI87, BBC^{+}94, BSD60, BP70, BP94c, BMPW66, CI95, CG67, Chu87, Cli73, Dou59, DRMN79, Ept80, FM67, FHW48, GLAM92, GL81a, Ham82, Ham87b, Har80, Hig88b, HP94b, JW85, Kel65, Lu94, Lu96, MPW66, Mil88b, OP64, Oet65, Ost73, Pen98b, PW71, RG67, Rob80, Rut58, Sha94, Tra66, Var72, Ver49, Vit69, Wac88, Wes68, Wil74a, Woź77b, You71, DR95, GW66, BS72]. Solutions [BRB76, BG67, Bjö67a, Bjö68, Bjö78, Bjö91, BG65, CP76, CR79, DK88, Fra92, GL95a, Gri85, Hea77, Hig89a, Kov76, Lan70, OPW65, vdS75, Sny55]. Solve [BR97, Kah86]. Solver [CR92, For69b, Mol2b, Tuc02]. Solvers [ADR92, AGL98, Che98, DH93, Enr89, Hig95b, SvdVM00]. Solves [And91, GP90]. Solving [AL85, ADD89, BG76, Bjö67b, Bol84, BKO94, BP71, BK77, BDV89, BGN70, Bye87, CH99a, Cro41, Dax90, DK85, DR82, For53, For69a, GWL^{+}92, GS90, Gol65, GS76, Gre97b, Gri85, HW91, Hig87b, Hig90g, Hig90h, Hig91a, HK93b, Hof87, JW86, KW89, KP96b, Lau79, LH95, LYK89, LZ92, MP74, Pai73, SB77, Smo95, WWWS86, Woź77a, Yip86]. Some [BS68, BK77, CH92a, Cry76, DDR95, FH55, FB91, Gar76, Hot43, JW86, Jor68, Lin74, Mac89, Men56, Pai90, Roh95, Rut47, Rut52, Sma90, Sun92c, Tao84b, Tyr91, Wil71b, Zie88, Kie73, MN79]. Somme [Pic72]. SOR [SN91]. Sorcerer [Whi97]. Soviet [War60]. Space [CH99a, FH55, Pow86, SW62]. Spaces [Ost73]. SPARCompiler [Sun92a, Sun92b]. Sparse [APS93, ADdR89, ADD89, BV92b, BLP90, DEL00, DRMN79, DR82, DER86, DGL89, DGR^{+}91, DGL92, ER74, EGL^{+}87, GL81a, GL81b, Hea84, HP94b, LD98, Liu87, MS02, Rei87, ZWS86, Bar86, DR95]. Special [BG94, MH57]. Specification [Dem92c, Wic89]. Specified [CHKL01, Ikr98, KP96a, ZZ95]. Spectra [BW93, Kos92, Tre99]. Spectral [BDG97, BDD^{+}97, Gem86, Hen62a, TT87]. Speed [Bai88b, BF88, Pan84]. Spline [AN63]. Splittings [Ple76]. Square [BR71b, Gen73, Hig87a, Hig97c, KM97, MM83a, Rus98, VCV01a]. Squares [Abd71, ADdR89, BI87, Bar88, BNP88, BV92a, BG67, Bjö67a, Bjö67b, Bjö68, Bjö78, Bjö87, Bjö90b, Bjö91, Bjö96, BG65, Cli73, CH98b, CH99a, CH99b, CH99c, Eld80, Far88a, Gen73, GS90, Gol65, GDD94, Gu98, GW92, Gul94, Gul95, Hea84, HV97, Jan97, JO74, Jor68, Kar74, KW97, KLR98a, LH95, Van85, Lon67, MS02, Man81, PW70, Ple74, PR69, Rei91, Rei00, Sau78, vdS75, Ste77a, Sun96, Sun97, WKS95, WSW72, Wei90, GW66]. Squaring [Fri91]. Stability [AR92, Bab69, BL80, Bjö87, Bjö92, Bol84, Bun87, BDV89, CH98b, CH99a, CH92b, DH92, Dem93, DHS95, Dia78, DGRS95, ER74, FGS96, GSS96, Hig90g, Hig90h, Hig91a, Hig92b, HP94b, Hig95b, Hig97b, Hig99a, Hig99b, Hig02, IM94, JP91, JW77, JT74, dJ77, KL88, Lin92, Mal00, Mat84a, Mat84b, MRW01, Mil75a, PW75, Pri92, RT92a, Rei71, Row90, Ske79, Ske80, vdS75, Smo95, TS90, Vit69, Woź77b, Woź77a, Yal95, Yal00, Yip86, Zaw91, Bar86, Bus71, Stu85c]. Stable [ABO88, BK77, Bun82, Bye88, CH99c, DGKS76, DH00, DK87, Fos91, GLKH90, GM73, GS76, Ham82, HK88, Hig97c, Gau90]. Stand [Tur91]. Standard [CL79, CCG^{+}84, CC93, CHN99, Don01, Fat82, GB91, IEE81, Hou81, KP79, Kah80, Kah81, Kah87b, IEE85, IEE87, Bas02, Kah96b]. Standards [Cod88b]. Start [KW92]. State [Wil85]. Static [Cod73, LD98]. Stationary [AR92, Dax90, HK93a, HK93b, Kni93]. Statistical [And71, GKL95, HS87, KG80, KL94, KLR98b, KLR98a, KC73, Lin75, Mai84, Mui82, SB94, Thi88, Tie70]. Statistics [Mac89, MN99, Ste95]. Statistique [Bru89, PV74]. Status [Kah96b]. Step [Gil51]. Step-By-Step [Gil51]. Stevenson [IEE81]. Stiffness [HT93]. Stochastic [BC86, Ste90, VA86]. Stock [Qui83]. Stopping [Ada67, ADR92, Ari00]. Storage [SV89]. Store [BSD60]. Story [Hou89, Lav80, Sch89]. Strassen [Pan78, BF88, BHH93]. Strassen-Type [BHH93]. Strassen’s [BLS91, DHSS94, Dum98, Kre76, Yuv78]. Strategies [Bun74, Dax82, Peñ96, Peñ98a]. Strategy [Liu87, MRW01, PN00]. Strict [Stu85a]. Strong [Bun87, BDV89, Smo95, Yal00]. Structured [BH92, Dem99, GK93, HH92a, PBRZ99, Pan01, Rum98, Sun98, Bar91c]. Structures [BN71, Jan97]. Study [BRK^{+}87, KC73]. Style [Bye84, KP96b]. Subgradient [Bar91c, Tao84a]. Subject [Gau95]. Subprograms [Bas02, BDD^{+}02, DCHH88b, DCDH90b, DCHH88a, DCDH90a, LHKK79]. Subroutine [CL83a, Cod88a, FHS78, IBM88]. Subroutines [DR82]. Subspace [HS95, HR92]. Subspaces [Ste73a]. Subtraction [FG82, FT86]. Successive [Lyn64]. Sufficient [Sha82]. Suite [DM89, SR97]. Sum [Fer95, HS81a, Per02, WSW72]. Sum-Free [WSW72]. Summation [And99, Bar91a, Cap71, Cle55, Dem01, Esp95, Hig93a, JSW83, KW00, Kie73, LO82, Lin70, Mal71, Neu74]. Summen [Neu74]. Summierungsverfahren [Nic70]. Summing [Smi65]. Sums [Dub83, Inc85, Lin74, MM83a]. Super [RSD77]. Supercomputers [BSBF89]. Superfast [Pan01]. Supplement [Gau95]. Support [CC93, Far88b]. Supremum [SS62]. Survey [BL77, COT89, Ede91b, Gar76, Hig87c, Hig94b, Kah72, Pow70, YG72]. SVD [Hig00]. Swapping [BD93b]. SWEEP [Goo79]. Sylvester [Syl04, GWL^{+}92, GLAM92, HR92, KW89, KP92, Kåg94, KP96b, Ste92]. Symbolic [Matc]. Symmetric [Aas71, AGL98, BG76, BP71, Bun71b, Bun74, BKP76, BK77, Bun82, BDV89, CL83a, CL83b, CH98a, Che98, CV86, DK77, Dax82, Doy57, DRMN79, DR82, DGR^{+}91, Fle76, FGS96, GSS96, Giv54, Hig88a, Hig98a, Hig99b, Ikr98, IK00, JP93, JP94, Liu87, Meu92, Ort63, Par98, PD00, Rum98, SvdVM00, Smo95, Van95, ZZ95, DR95]. Symposium [Har48]. Synthetic [Ste71]. System [AL85, ADdR89, Bai95, Bar89, Bjö92, Che98, DH93, Ede91b, GEJ95, Hig95b, MPW66, Oet65, RG67, SA75, Yip86, Uni92]. système [Ben24]. Systems [AM94, ADD89, AR92, Ata40, BLS91, BMvN46, BBC^{+}94, Bar91c, BH92, BG76, BP70, BE73a, BP94c, Boh75, BG94, dBP77, BKO94, BMPW66, Bre73, BP71, BK77, BDV89, CF88, CN96, CI95, CP76, CR79, Cro41, Dax90, DK85, FGS96, FM67, Gau90, GL81a, GMPS92, GSS96, GV79, Gre97b, Han97, Hig87b, Hig88b, Hig89a, Hig90b, Hig90h, HH92a, HH92b, HK93b, HP94b, Hig97a, Hof87, Kel65, KL88, KLR98b, Kie81, KC73, Lu94, Lu95, Lu96, LM66, MP74, Mat92, MK85, MP77, Mol74, NP84, OPW65, Peñ96, RO91, vR01, Roh89a, Roh89b, SB77, SvdVM00, vdS70, Smo95, SS97, Sun98, Var72, Wic92, Wil74a, Woź77b, You71, DR95, GP90, Woź78]. Systolic [BI87].

Table [Tan89, Tan90b, Tan91, Tan92]. Table-Driven [Tan89, Tan90b, Tan92]. Table-Lookup [Tan91]. Tables [Leh43, SL54]. Tale [Mol95]. Tape [BSD60]. Targets [Hun97]. Task [IEE81]. Taylor [Lin76]. Tears [Pry81]. Technical [Bas02, Mol90b]. Technique [Dek71, Pan78]. Techniques [DK77, Wil62, Ral81]. Technology [IEC94, War60]. Templates [BBC^{+}94]. Term [Qui94, QT90, Tro84b]. Terms [JSW83]. Test [BDDD97, BG79, BPR^{+}97, Caf63, Cod93, DM89, DCHH88b, DCDH90b, DGL89, Ede93a, Hen66, Rut68, Sch81, Zie86, Hig95c]. Testing [Cod82, GK69, Ker90, LDB^{+}00, Pax91, Tan90a, VCV01a, VCV01b]. Tests [HS66, JW86]. Tetrahedron [Kah01]. Their [Hig95a, Tan91, DH00, Smi65]. Them [Lav80]. Theorem [FV62, Ove01, WT94, Syl50]. Theorems [DW84, Dun90, Gil91, SS62, Stu85c, Syl50]. Theoretical [Wen66]. Theorie [Hil94]. Theory [BCS97, CF92, DK88, Eld80, Gan59a, Gan59b, Gau95, Geu82, Hal67, HHKR93, HH92b, Hig93c, Hig94b, Hod92, Hou64, Kat76, KM81, Mui82, O’C93, Pet81, Ric66b, SS90, Ste90, Stu81, Tau88, Tre90, Tur29, TA32, Usp48, Wed73, Cod88b, HSS99, KP96a, SP88]. Third [Tor65]. Thomas [Bel31]. Three [CR83, Hig91b, Hig92b, Mol78, QTD91, Tre85]. Thumb [Ove01]. Time [And71, Coo94, Dhi98, GLKH90, GL95b, KW00]. Today [For67]. Toeplitz [BW93, CN96, CV86, Mak91, RT92b, Ste97d, Str86, Tyr91]. Tolerance [BGM^{+}94]. Tool [BRR85, BC86, VCV01a, VCV01b]. Toolbox [HHKR93, Matc, Hig95c]. Tools [Ral91]. Topics [Bar91c, HJ91]. Torchbearer [Tau88]. Total [GP92, Kar68]. Totally [And87, dBP77, Cry73, Cry76, Fal01]. Trade [Sto91]. Trading [Dem93]. Traditional [LYK89]. Transcendental [FB91]. Transcendentals [LMT98]. Transform [BH95, Lin92, Ram71, Sch96, Van92]. Transform-Based [Lin92]. Transformation [BG67, Pry85, Rut58, Tsa75]. Transformations [BG65, CM79, Dem83, Gen73, Gen75, Mil76, PR69, Rat82, SV89, SW62, Wil62, Wil65b, Ypm83b]. Transforms [GS66]. Transients [Ste97a]. Translation [Bai93]. Transportable [For78]. Trapezoid [Lem75]. Treats [Cho83]. Triangle [Gra92, Kah97]. Triangular [APS93, And91, Bre70b, Cho68, ER74, Hig87c, Hig89a, HP94b, Hig95b, KP92, KS00, Lem75, SB77, Ste97c, VT98, Bar86]. Triangularization [Hou58]. Tricks [Cho83]. Tridiagonal [Aas71, AM94, BG94, BE73b, Dhi98, Hig86b, Hig90b, Hig99b, MP74, Mat84b, Meu92, Var72]. Tridiagonals [PD00]. Trilinear [Pan78]. Triplex [Nic70]. Triplex-ALGOL [Nic70]. Trummer’s [Lu95]. Truncation [Kah65, Kar74, Ros65, Wol64]. Turing [CK84, Wil80, HAC^{+}85, Bod49, Bri92, Hod83, HC88, Inc92, ACM87]. Turing’s [CD86]. Tutorial [Goo79]. Twenty [ACM87]. Two [Bar91c, Bir75, Gau95, Knu92, Mol95, Var79]. Type [BHH93, Ede91a, Kea96a]. Typical [Bow53]. Typography [Knu99].

Unconstrained [DS83, Pow70, Pow75]. Underdetermined [AL85, CP76, DH93, SS97]. Underflow [Coo81, Dem84, FT86, Neu85]. Underflow-Free [MI81]. UNICOS [Inc89]. Unified [Pan01]. Uniform [IM94, Tra66]. Unit [CHH83, KS00, Sch81, Tao84b]. Unitary [BM94, Leh96, Mat96, Hou58]. Uniting [Pan78]. Units [GM74]. Unstable [Bye88, Wri93]. Unstructured [Dax90]. Unsymmetric [EGL^{+}87, GV79]. Untersuchungen [Spi76]. Unzalässige [PN85]. Update [ES92]. updated [BDD^{+}02]. Updates [SB92]. Updating [DGKS76]. Upper [MAV^{+}97, Sha91]. Usage [LHKK79]. Use [And91, Fro01, New81, Rob80, Wic92, Wil55, Wil65b, Woo58]. User [ABB^{+}99, BBD^{+}01, BCC^{+}97, DBMS79, DGL92, GMSW86, Lon67, Hew93]. User’s [NAG86, Matc, Sun92b]. Uses [Kea96b]. Using [Bai88a, BLS91, BSD60, BP94c, Bre70b, Don01, FGM95, Gul95, HFT94, JW85, Lad76, Mil88a, Qui94, Ros65, Sau72, Sto77, Sny55].

V [CGG^{+}91]. Validating [Rum01]. Validation [NAG86]. Value [BD93a, BR77, CI94, Dem99, DGE^{+}99, Enr89, GDD94, MM83b, Ste93a, Var75]. Values [And90, Chu99, Chu00, DK90, Giv54, Hen62a, Jen82, Kit95, KS00]. Vandermonde [Bar91c, BH92, Bec00, BP70, BE73a, BKO94, CR92, CR93, Dem89a, Dem90, Gau62, Gau75, Gau78, Gau83, Gau90, GO94, Hig87b, Hig88b, Hig90h, Lu94, Lu95, Lu96, RO91, Rei91, Sun98, TG81]. Vandermonde-Like [Bar91c, CR93, Gau83, Lu96, Rei91, BH92, Hig88b, Hig90h]. Variable [Gus97, Hen62b, Kie81]. Variable-Precision [Kie81]. Variance [Bar91a, CGL83]. Variant [DHSS94, GS76]. Variants [DS94]. Variation [BM94, Hen62a]. Variations [Gil91]. Various [Bre73, Don01]. Varying [Dem87b, Tie70]. Vec [HS81b]. Vec-Permutation [HS81b]. Vector [Blu78, DGK84, Hof87, War76]. Vectors [BN87, Pry84, Ruh83, Wil55]. Verfahren [Neu74]. Verifiable [Mil88a]. Verified [HHKR93]. Vermutung [PN85]. Versatile [Cra79]. Version [GMSW86, Hig95c, Matc]. Verstärkungen [PN85]. Versus [Kah94, Wer84]. Very [GS90, JSW83]. Via [DG87, CW87, DL94]. View [Lon67, Pow75]. Vintage [Lav80]. Virgule [Pic72]. Vol [Tod77]. Volume [BP94b, Kah01, Ste01]. Vortex [BKP93].

w [Wil67]. W. [Kah73]. Watt [Gen69]. Wavelet [ABCR93]. Wavelet-Like [ABCR93]. Ways [MV78]. Weak [Bun87]. Web [BPR^{+}97]. Weighted [CH98b, GW92, Gul94, Gul95, HV97]. Weighting [Van85]. Well [CF88]. Well-Conditioned [CF88]. where [Har80]. Which [Asp59, Bus68, SW98, Wri93]. Who [Lav80]. Whose [DRMN79]. Wielandt [Fra58]. Wilkinson [Eva76, Ano87, Fox78, Fox87, Par90, PC83, PN85]. Will [Tur91]. William [Sev98, Woe97]. Winograd [DHSS94]. Winograd’s [Bre70b, Har72]. Wise [Bjö91, CH99c]. Wishart [Ede91a]. within [Hig90e]. Without [CG85, DG94, Gen73, YC97]. Woodbury [Yip86]. Word [Buc81, CCG^{+}84]. Word-Length-Independent [CCG^{+}84]. Work [Act70, SW89, Wil80]. Works [Bri92, Inc92, Tau63]. World [Mol90a]. Worst [LM01, TZ01]. Wprowadzenie [KS92]. Writing [Hig98b]. Written [Kah83]. WY [BV87, Pug92, SV89].

XSC [HHKR93, KKN^{+}92, KKL^{+}93].

Y-MP [Car91]. Year [QTD91]. Years [EG91, ACM87]. Yielding [Kub91].

Zaokragleń [Wil67]. Zaphod [Gra92]. Zautomatyzowanych [KS92]. Zeitgewinns [Spi76]. Zero [Ypm83b]. Zeros [Kit95, Ste71]. zur [Hil94, Neu74, Spi76]. Zwei [PN85].

References

Aasen:1971:RSM

[Aas71] Jan Ole Aasen. On the reduction of a symmetric matrix to tridiagonal form. BIT, 11: 233–242, 1971. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Aggarwal:1979:REM

[AB79] Vijay B. Aggarwal and James W.

Burgmeier. A round-off error model with applications to arithmetic expressions. SIAM J. Comput., 8(1):60–72, 1979.

Anderson:1999:LUG

[ABB^{+}99] E. Anderson, Z. Bai, C. H.

Bischof, S. Blackford, J. W. Demmel, J. J. Dongarra, J. J. Du Croz, A. Greenbaum, S. J. Hammarling, A. McKenney, and D. C. Sorensen. LA-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, third edition, 1999. ISBN 0-89871-447-8. xxvi + 407 pp.

Alpert:1993:WLB

[ABCR93] B. Alpert, G. Beylkin, R. Coifman, and V. Rokhlin. Waveletlike bases for the fast solution of second-kind integral equations. SIAM J. Sci. Comput., 14(1): 159–184, 1993.

Abdelmalek:1971:REA

[Abd71] Nabih N. Abdelmalek. Round off error analysis for Gram– Schmidt method and solution of linear least squares problems. BIT, 11:345–368, 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Anderson:1992:GQF

[ABD92] E. Anderson, Z. Bai, and J. Dongarra. Generalized QR factorization and its applications. Linear Algebra Appl., 162–164:243–271, 1992.

Ahac:1988:SFM

[ABO88] Alan A. Ahac, John J. Buoni, and D. D. Olesky. Stable LU factorization of H-matrices. Linear Algebra Appl., 99:97– 110, 1988.

acm87

[ACM87] ACM Turing Award Lectures: The First Twenty Years, 1966– 1985. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-54885-2. xviii + 483 pp.

Acton:1970:NMW

[Act70] Forman S. Acton. Numerical Methods That Work. Harper and Row, New York, 1970. ISBN 0-88385-450-3. xviii + 541 pp. Reprinted by Mathematical Association of America, Washington, D.C., with new preface and additional problems, 1990.

Acton:1996:RCM

[Act96] Forman S. Acton. Real Computing Made Real: Preventing Errors in Scientific and Engineering Calculations. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-03663-2. xv + 259 pp.

Adams:1967:SCP

[Ada67] Duane A. Adams. A stopping criterion for polynomial root finding. Comm. ACM, 10 (10):655–658, 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arioli:1989:SSL

[ADD89] M. Arioli, J. W. Demmel, and I. S. Duff. Solving sparse linear systems with sparse backward error. SIAM J. Matrix Anal. Appl., 10(2):165–190, 1989.

Arioli:1989:ASA

[ADdR89] M. Arioli, I. S. Duff, and P. P. M. de Rijk. On the augmented system approach to sparse least-squares problems. Numer. Math., 55:667– 684, 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1992:SCI

[ADR92] Mario Arioli, Iain S. Duff, and Daniel Ruiz. Stopping criteria for iterative solvers. SIAM J. Matrix Anal. Appl., 13(1):138– 144, 1992.

Almacany:1984:DCA

[ADW84] M. Almacany, C. B. Dunham, and J. Williams. Discrete Chebyshev approximation by interpolating rationals. IMA J. Numer. Anal., 4:467–477, 1984.

Ashcraft:1998:ASI

[AGL98] Cleve Ashcraft, Roger G. Grimes, and John G. Lewis. Accurate symmetric indefinite linear equation solvers. SIAM J. Matrix Anal. Appl., 20(2):513– 561, 1998.

Alonso:1997:BEA

[AGP97] P. Alonso, M. Gasca, and J. M. Peña. Backward error analysis of Neville elimination. Appl. Numer. Math., 23:193–204, 1997.

Alefeld:1983:IIC

[AH83] Göltz Alefeld and Jürgen Herzberger. Introduction to Interval Computations. Academic Press, New York, 1983. ISBN 0-12-049820-0. xviii + 333 pp.

Arnold:1984:GEA

[AL84] William F. Arnold and Alan J. Laub. Generalized eigenproblem algorithms and software for algebraic Riccati equations. Proceedings of the IEEE, 72 (12):1746–1754, 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Arioli:1985:EAA

[AL85] M. Arioli and A. Laratta. Error analysis of an algorithm for solving an underdetermined linear system. Numer. Math., 46:255– 268, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1986:EAA

[AL86] M. Arioli and A. Laratta. Error analysis of algorithms for computing the projection of a point onto a linear manifold. Linear Algebra Appl., 82:1–26, 1986.

Ewing:1991:PHP

[Alb91] Paul Halmos by parts (interviews by Donald J. Albers). In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 3–32. Springer-Verlag, Berlin, 1991.

Ashenhurst:1965:EEC

[AM65] R. L. Ashenhurst and N. Metropolis. Error estimation in computer calculation. Amer. Math. Monthly, 72(2):47–58, 1965.

Althoen:1987:GJR

[AM87] Steven C. Althoen and Renate Mclaughlin. Gauss–Jordan reduction: A brief history. Amer. Math. Monthly, 94(2):130–142, 1987.

Amodio:1994:BEA

[AM94] Pierluigi Amodio and Francesca Mazzia. Backward error analysis of cyclic reduction for the solution of tridiagonal systems. Math. Comp., 62(206):601–617, 1994.

Ahlberg:1963:CPS

[AN63] J. H. Ahlberg and E. N. Nilson. Convergence properties of the spline fit. J. Soc. Indust. Appl. Math., 11(1):95–104, 1963.

Anderson:1971:SAT

[And71] T. W. Anderson. The Statistical Analysis of Time Series. Wiley, New York, 1971. ISBN 0-471-02900-9. xiv + 704 pp.

Ando:1987:TPM

[And87] T. Ando. Totally positive matrices. Linear Algebra Appl., 90: 165–219, 1987.

Andrew:1990:ESV

[And90] Alan L. Andrew. Eigenvalues and singular values of certain random matrices. J. Comput. Appl. Math., 30:165–171, 1990.

Anderson:1991:RTS

[And91] Edward Anderson. Robust triangular solves for use in condition estimation. Technical Report CS-91-142, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 1991. 35 pp. LAPACK Working Note 36.

Anderson:1999:DAF

[And99] I. J. Anderson. A distillation algorithm for floating-point summation. SIAM J. Sci. Comput., 20(5):1797–1806, 1999.

Anonymous:1922:CC

[Ano22] Anonymous. Le Commandant Cholesky. Bulletin Géodésique (Toulouse), pages 159–161, 1922. CODEN BGDQAG. ISSN 0007-4632. Translation by Richard W. Cottle (“Major Cholesky”) appears in [Mai84, Appendix] and in NA Digest, Volume 90, Issue 7, 1990.

Anonymous:1987:JW

[Ano87] Anonymous. James Wilkinson (1919–1986). Annals of the History of Computing, 9(2):205– 210, 1987. From the introduction: “A series of lightly edited extracts from messages that were sent over various computer networks during the period October 5, 1986–February 13, 1987”.

Anderson:1987:GRO

[AOU87] T. W. Anderson, I. Olkin, and L. G. Underhill. Generation of random orthogonal matrices. SIAM J. Sci. Statist. Comput., 8(4):625–629, 1987.

Anda:1994:FPR

[AP94] Andrew A. Anda and Haesun Park. Fast plane rotations with dynamic scaling. SIAM J. Matrix Anal. Appl., 15(1):162–174,

Alvarado:1993:HPS

[APS93] Fernando L. Alvarado, Alex Pothen, and Robert S. Schreiber. Highly parallel sparse triangular solution. In J. Alan George, John R. Gilbert, and Joseph W. H. Liu, editors, Graph Theory and Sparse Matrix Computations, volume 56 of IMA Volumes in Mathematics and Its Applications, pages 141– 158. Springer-Verlag, New York, 1993.

Arioli:1992:SCC

[AR92] Mario Arioli and Francesco Romani. Stability, convergence, and conditioning of stationary iterative methods of the form x^{(i+1)} = P x^{(i)} + q for the solution of linear systems. IMA J. Numer. Anal., 12:21–30, 1992.

Arioli:2000:SCC

[Ari00] Mario Arioli. A stopping criterion for the conjugate gradient algorithm in a finite element method framework. Technical Report 1179, Istituto di Analisi Numerica - C.N.R., Pavia, Italy, 2000. 12 pp.

Asplund:1959:IMW

[Asp59] Edgar Asplund. Inverses of matrices {a_{ij}} which satisfy a_{ij} = 0 for j > i + p. Mathematica Scandinavica, 7:57–60, 1959. CODEN MTSCAN. ISSN 0025-5521 (print), 1903-1807 (electronic).

Atanasoff:1940:CMS

[Ata40] John V. Atanasoff. Computing machine for the solution of large systems of linear algebraic equations. Unpublished manuscript, Iowa State College, Ames, IA, USA, August 1940. Reprinted in [Ran75, pp. 315–335].

Alt:1981:CBC

[AvL81] H. Alt and J. van Leeuwen.

The complexity of basic complex operations. Computing: Archiv für informatik und numerik, 27:205–215, 1981. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Axelsson:1994:ISM

[Axe94] Owe Axelsson. Iterative Solution Methods. Cambridge University Press, 1994. ISBN 0-521-44524-8. xiii + 654 pp.

Babuska:1969:NSM

[Bab69] Ivo Babuška. Numerical stability in mathematical analysis. In Proc. IFIP Congress, Information Processing 68, pages 11– 23. North-Holland, Amsterdam, The Netherlands, 1969.

Bailey:1988:CDD

[Bai88a] David H. Bailey. The computation of π to 29,360,000 decimal digits using Borweins’ quartically convergent algorithm. Math. Comp., 50(181):283–296,

Bailey:1988:EHS

[Bai88b] David H. Bailey. Extra high speed matrix multiplication on the Cray-2. SIAM J. Sci. Statist. Comput., 9(3):603–607, 1988.

Bailey:1993:AMT

[Bai93] David H. Bailey. Algorithm 719: Multiprecision translation and execution of FORTRAN programs. ACM Trans. Math. Software, 19(3):288–319, 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bailey:1995:FBM

[Bai95] David H. Bailey. A Fortran 90-based multiprecision system. ACM Trans. Math. Software, 21(4):379–387, 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bard:1974:NPE

[Bar74] Yonathan Bard. Nonlinear Parameter Estimation. Academic Press, New York, 1974.

Barlow:1981:DAR

[Bar81a] Jesse L. Barlow. On the distribution of accumulated roundoff error in floating point arithmetic. In Proc. 5th IEEE Symposium on Computer Arithmetic, Ann Arbor, Michigan, pages 100–105, 1981.

Barlow:1981:PEA

[Bar81b] Jesse L. Barlow. Probabilistic Error Analysis of Floating Point and CRD Arithmetics. PhD thesis, Northwestern University, Evanston, IL, USA, June 1981.

barl86

[Bar86] Jesse L. Barlow. A note on monitoring the stability of triangular decomposition of sparse matrices. SIAM J. Sci. Statist. Comput., 7(1):166–168, 1986.

Barlow:1988:EAI

[Bar88] Jesse L. Barlow. Error analysis and implementation aspects of deferred correction for equality constrained least squares problems. SIAM J. Numer. Anal., 25(6):1340–1358, 1988.

Barrett:1989:FMA

[Bar89] Geoff Barrett. Formal methods applied to a floating-point number system. IEEE Trans. Software Engrg., 15(5):611–621, 1989.

Barlow:1991:EAP

[Bar91a] Jesse L. Barlow. Error analysis of a pairwise summation algorithm to compute the sample variance. Numer. Math., 58:583–590, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrlund:1991:PBD

[Bar91b] Anders Barrlund. Perturbation bounds for the LDL^{H} and LU decompositions. BIT, 31: 358–363, 1991. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bartels:1991:TTM

[Bar91c] Sven G. Bartels. Two topics in matrix analysis: Structured sensitivity for Vandermondelike systems and a subgradient method for matrix norm estimation. M.Sc. Thesis, Department of Mathematics and Computer Science, University of Dundee, Dundee, Scotland, September 1991.

Barrlund:1994:PBG

[Bar94] Anders Barrlund. Perturbation bounds for the generalized QR factorization. Linear Algebra Appl., 207:251–271, 1994.

BLAST

[Bas02] Basic Linear Algebra Subprograms Technical (BLAST) Forum Standard. Int. J. High Performance Applications and Supercomputing, 2002. To appear.

Bauer:1963:OSMb

[Bau63a] F. L. Bauer. Optimal scaling of matrices and the importance of the minimal condition. In Cicely M. Popplewell, editor, Proc. IFIP Congress 1962, Information Processing 62, pages 198–201. North-Holland, Amsterdam, The Netherlands, 1963.

Bauer:1963:OSMa

[Bau63b] F. L. Bauer. Optimally scaled matrices. Numer. Math., 5:73– 87, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1966:GLL

[Bau66] F. L. Bauer. Genauigkeitsfragen bei der Lösung linearer Gleichungssysteme. Z. Angew. Math. Mech., 46(7):409–421, 1966.

Bauer:1969:ROS

[Bau69] F. L. Bauer. Remarks on optimally scaled matrices. Numer. Math., 13:1–3, 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1974:CGR

[Bau74] F. L. Bauer. Computational graphs and rounding errors. SIAM J. Numer. Anal., 11(1): 87–96, 1974.

Barlow:1985:RED

[BB85a] Jesse L. Barlow and E. H.

Bareiss. On roundoff error distributions in floating point and logarithmic arithmetic. Computing: Archiv für informatik und numerik, 34:325–347, 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barlow:1985:PEA

[BB85b] Jesse L. Barlow and E. H.

Bareiss. Probabilistic error analysis of Gaussian elimination in floating point and logarithmic arithmetic. Computing: Archiv für informatik und numerik, 34:349–364, 1985. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Borwein:1989:RME

[BBB89] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to Pi or how to compute one billion digits of Pi. Amer. Math. Monthly, 96(3): 201–219, 1989.

Barrett:1994:TSL

[BBC^{+}94] Richard Barrett, Michael Berry, Tony F. Chan, James Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-328-5. xiii + 112 pp.

Barker:2001:LUG

[BBD^{+}01] V. A. Barker, L. S. Blackford, J. J. Dongarra, J. J. Du Croz, S. J. Hammarling, M. Marinova, J. Waśniewski, and P. Yalamov. LAPACK95 Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-504-0. xviii + 258 pp.

Bliss:1992:API

[BBG92] B. Bliss, M.-C. Brunet, and E. Gallopoulos. Automatic program instrumentation with applications in performance and error analysis. In E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors, Expert Systems for Scientific Computing, pages 235–260. North-Holland, Amsterdam, The Netherlands, 1992.

Brunet:1986:CTS

[BC86] Marie-Christine Brunet and Françoise Chatelin. CESTAC, a tool for a stochastic roundoff error analysis in scientific computing. In R. Vichnevetsky and J. Vignes, editors, Numerical Mathematics and Applications, pages 11–20. Elsevier Science Publishers B.V. (North-Holland), Amsterdam, The Netherlands, 1986.

Blackford:1997:SUG

[BCC^{+}97] L. S. Blackford, J. Choi, A. Cleary, E. D’Azevedo, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, G. Henry, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-397-8. xxvi + 325 pp.

Blackford:1997:PEN

[BCP^{+}97] L. S. Blackford, A. Cleary, A. Petitet, R. C. Whaley, J. Demmel, I. Dhillon, H. Ren, K. Stanley, J. Dongarra, and S. Hammarling. Practical experience in the numerical dangers of heterogeneous computing. ACM Trans. Math. Software, 23(2):133–147, 1997. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Burgisser:1997:ACT

[BCS97] Peter Bürgisser, Michael Clausen, and M. Amin Shokrollahi. Algebraic Complexity Theory. Springer-Verlag, Berlin, 1997. ISBN 0-540-60582-7. xxiii + 618 pp.

Bai:1993:CGS

[BD93a] Zhaojun Bai and James W. Demmel. Computing the generalized singular value decomposition. SIAM J. Sci. Comput., 14 (6):1464–1486, 1993.

Bai:1993:SDB

[BD93b] Zhaojun Bai and James W. Demmel. On swapping diagonal blocks in real Schur form. Linear Algebra Appl., 186:73–95, 1993.

Bai:1997:SDN

[BDD^{+}97] Zhaojun Bai, James W. Demmel, Jack J. Dongarra, Antoine Petitet, Howard Robinson, and K. Stanley. The spectral decomposition of nonsymmetric matrices on distributed memory parallel computers. SIAM J. Sci. Comput., 18(5):1446–1461, 1997.

Blackford:2002:USB

[BDD^{+}02] L. Susan Blackford, James Demmel, Jack Dongarra, Iain Duff, Sven Hammarling, Greg Henry, Michael Heroux, Linda Kaufman, Andrew Lumsdaine, Antoine Petitet, Roldan Pozo, Karin Remington, and R. Clint Whaley. An updated set of Basic Linear Algebra Subprograms (BLAS). ACM Trans.

Math. Software, 28(2):135–151, June 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bai:1997:TMC

[BDDD97] Zhaojun Bai, David Day, James Demmel, and Jack Dongarra. A test matrix collection for non-Hermitian eigenvalue problems (release 1.0). Technical Report CS-97-355, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, March 1997. 45 pp. LAPACK Working Note 123.

Browne:1994:NSR

[BDG^{+}94] Shirley V. Browne, Jack J. Dongarra, Stan C. Green, Keith Moore, Thomas H. Rowan, and Reed C. Wade. Netlib services and resources. Report ORNL/TM-12680, Oak Ridge National Laboratory, Oak Ridge, TN, USA, April 1994. 42 pp.

Bai:1997:IFP

[BDG97] Zhaojun Bai, James W. Demmel, and Ming Gu. Inverse free parallel spectral divide and conquer algorithms for nonsymmetric eigenproblems. Numer. Math., 76:279–308, 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bai:1993:CCN

[BDM93] Zhaojun Bai, James W. Demmel, and Alan McKenney. On computing condition numbers for the nonsymmetric eigenproblem. ACM Trans. Math. Software, 19(2):202–223, 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bunch:1989:SSA

[BDV89] James R. Bunch, James W. Demmel, and Charles F. Van Loan. The strong stability of algorithms for solving symmetric linear systems. SIAM J. Matrix Anal. Appl., 10(4):494–499, 1989.

Bjorck:1973:ACV

[BE73a] Åke Björck and Tommy Elfving. Algorithms for confluent Vandermonde systems. Numer. Math., 21:130–137, 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bukhberger:1973:MIT

[BE73b] B. Bukhberger and G. A. Emel’yanenko. Methods of inverting tridiagonal matrices. U.S.S.R. Computational Mathematics and Mathematical Physics, 13:10–20, 1973. CO-DEN CMMPA9. ISSN 0041-5553, 0502-9902.

Beckermann:2000:CNR

[Bec00] Bernhard Beckermann. The condition number of real Vandermonde, Krylov and positive definite Hankel matrices. Numer. Math., 85:553–577, 2000. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bell:1931:RCH

[Bel31] E. T. Bell. Review of “Contributions to the History of Determinants, 1900-1920”, by Sir Thomas Muir. Amer. Math. Monthly, 38:161–164, 1931.

Bellman:1970:IMA

[Bel70] Richard Bellman. Introduction to Matrix Analysis. McGraw-Hill, New York, second edition, 1970. xxiii + 403 pp. Reprinted by Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-399-4.

Benoit:1924:NMR

[Ben24] Commandant Benoit. Note sur une méthode de résolution des équations normales provenant de l’application de la méthode des moindres carrés à un système d’équations linéaires en nombre inférieur à celui des inconnues. Application de la méthode à la résolution d’un système défini d’équations linéaires (Procédé du Commandant Cholesky). Bulletin Géodésique (Toulouse), 2:67–77, 1924. Cited in [Ano22, Cottle’s translation].

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 78(4):551–572, March 1938. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://links.
jstor.org/sici?sici=0003-
049X%2819380331%2978%3A4%
3C551%3ATLOAN%3E2.0.CO%3B2-

G. See comments about Benford’s biased rounding practices [?].

Berenbaum:1991:DSM

[Ber91]  M. C. Berenbaum.  Direct

search methods in the optimisation of cancer chemotherapy. British Journal of Cancer, 61: 101–109, 1991. CODEN BJ-CAAI. ISSN 0007-0920 (print), 1532-1827 (electronic).

Bailey:1988:SNA

[BF88] D. H. Bailey and H. R. P. Ferguson. A Strassen–Newton algorithm for high-speed parallelizable matrix inversion. In Proceedings of Supercomputing ’88, pages 419–424. IEEE Computer Society Press, New York, 1988.

Businger:1965:LLS

[BG65] P. Businger and G. H. Golub. Linear least squares solutions by Householder transformations. Numer. Math., 7:269–276, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 111– 118], Contribution I/8.

Bjorck:1967:IRLa

[BG67] Åke Björck and Gene H. Golub. Iterative refinement of linear least squares solutions by Householder transformation. BIT, 7:322–337, 1967. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Barwell:1976:CAS

[BG76] Victor Barwell and Alan George.

A comparison of algorithms for solving symmetric indefinite systems of linear equations. ACM Trans. Math. Software, 2(3): 242–251, 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Birkhoff:1979:IDT

[BG79] Garrett Birkhoff and Surender Gulati. Isotropic distributions of test matrices. J. Appl. Math. Phys. (ZAMP), 30:148– 158, 1979.

Bondeli:1994:CRS

[BG94] S. Bondeli and W. Gander.

Cyclic reduction for special tridiagonal systems. SIAM J. Matrix Anal. Appl., 15(1):321– 330, 1994.

Boley:1994:FPF

[BGM^{+}94] Daniel Boley, Gene H. Golub, Samy Makar, Nirmal Saxena, and Edward J. McCluskey. Floating point fault tolerance with backward error assertions. IEEE Trans. Comput., 44(2): 302–311, 1994.

Buzbee:1970:DMS

[BGN70] B. L. Buzbee, G. H. Golub, and C. W. Nielson. On direct methods for solving Poisson’s equations. SIAM J. Numer. Anal., 7 (4):627–656, 1970.

Bartels:1992:SSV

[BH92] Sven G. Bartels and Desmond J. Higham. The structured sensitivity of Vandermonde-like systems. Numer. Math., 62:17– 33, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Briggs:1995:DOM

[BH95] William L. Briggs and Van Emden Henson. The DFT: An Owner’s Manual for the Discrete Fourier Transform. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-342-0. xv + 434 pp.

Bhatia:1997:MA

[Bha97] Rajendra Bhatia. Matrix Analysis, volume 169 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94846-5 (print), 1-4612-0653-7 (e-book). xii + 347 pp. LCCN QA188 .B485 1996.

URL http://catdir.loc.gov/
catdir/enhancements/fy0818/
96032217-d.html;   http://
www.gbv.de/dms/goettingen/
213798093.pdf.

Balle:1993:STM

[BHH93] Susanne M. Balle, Per Christian Hansen, and Nicholas J. Higham. A Strassen-type matrix inversion algorithm for the Connection Machine. Technical Report CNC/1993/028, Centre for Novel Computing, University of Manchester, Manchester, England, October 1993. 29 pp.

Brent:1980:AIB

[BHY80] Richard P. Brent, Judith A.

Hooper, and J. Michael Yohe. An AUGMENT interface for Brent’s multiple precision arithmetic package. ACM Trans. Math. Software, 6(2):146–149, 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barlow:1987:SGR

[BI87] Jesse L. Barlow and Ilse C. F.

Ipsen. Scaled Givens rotations for the solution of linear least squares problems on systolic arrays. SIAM J. Sci. Statist. Comput., 8(5):716–733, 1987.

Birkhoff:1975:THN

[Bir75] Garrett Birkhoff. Two Hadamard numbers for matrices. Comm. ACM, 18(1):25–29, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bischof:1990:ICEa

[Bis90] Christian H. Bischof. Incremental condition estimation. SIAM J. Matrix Anal. Appl., 11(2): 312–322, 1990.

Bjorck:1967:IRLb

[Bjö67a] Åke Björck. Iterative refinement of linear least squares solutions I. BIT, 7:257–278, 1967. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjorck:1967:SLL

[Bjö67b]  Åke Björck.  Solving lin-

ear least squares problems by Gram–Schmidt orthogonalization. BIT, 7:1–21, 1967. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjorck:1968:IRL

[Bjö68] Åke Björck. Iterative refinement of linear least squares solutions II. BIT, 8:8–30, 1968. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjorck:1978:CIR

[Bjö78] Åke Björck. Comment on the iterative refinement of leastsquares solutions. J. Amer. Statist. Assoc., 73(361):161– 166, 1978.

Bjorck:1987:SAM

[Bjö87] Åke Björck. Stability analysis of the method of seminormal equations for linear least squares problems. Linear Algebra Appl., 88/89:31–48, 1987.

Bjorck:1990:IRR

[Bjö90a] Åke Björck. Iterative refinement and reliable computing. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 249– 266. Oxford University Press, 1990.

Bjorck:1990:LSM

[Bjö90b] Åke Björck. Least squares methods. In P. G. Ciarlet and J. L. Lions, editors, Handbook of Numerical Analysis, volume I: Finite Difference Methods— Solution of Equations in R^{n}. Elsevier/North-Holland, Amsterdam, The Netherlands, 1990.

Bjorck:1991:CWP

[Bjö91] Åke Björck. Component-wise perturbation analysis and error bounds for linear least squares solutions. BIT, 31: 238–244, 1991. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bjorck:1992:PSA

[Bjö92] Åke Björck. Pivoting and stability in the augmented system method. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 1–16. Longman Scientific and Technical, Essex, UK, 1992.

Bjorck:1994:NGS

[Bjö94] Åke Björck. Numerics of Gram– Schmidt orthogonalization. Linear Algebra Appl., 197/198:297– 316, 1994.

Bjorck:1996:NML

[Bjö96] Åke Björck. Numerical Methods for Least Squares Problems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-360-9. xvii + 408 pp.

Bunch:1977:SSM

[BK77] James R. Bunch and Linda Kaufman. Some stable methods for calculating inertia and solving symmetric linear systems. Math. Comp., 31(137):163–179, 1977.

Baksalary:1979:ME

[BK79] J. K. Baksalary and R. Kala. The matrix equation AX − Y B = C. Linear Algebra Appl., 25:41–43, 1979.

Baksalary:1980:ME

[BK80] J. K. Baksalary and R. Kala. The matrix equation AXB + CY D = E. Linear Algebra Appl., 30:141–147, 1980.

Boros:1994:FAS

[BKO94] T. Boros, T. Kailath, and V. Olshevsky. Fast algorithms for solving Vandermonde and Chebyshev–Vandermonde systems. Manuscript, 1994.

Bunch:1976:DSM

[BKP76] James R. Bunch, Linda Kaufman, and Beresford N. Parlett. Decomposition of a symmetric matrix. Numer. Math., 27:95– 109, 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bailey:1993:MPM

[BKP93] David H. Bailey, Robert Krasny, and Richard Pelz. Multiple precision, multiple processor vortex sheet roll-up computation. In Richard F. Sincovec, David E. Keyes, Michael R. Leuze, Linda R. Petzold, and Daniel A. Reed, editors, Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, Volume I, pages 52–56. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993.

Birkhoff:1977:SMA

[BL77] Garrett Birkhoff and Saunders Mac Lane. A Survey of Modern Algebra. Macmillan, New York, fourth edition, 1977. ISBN 0-02-310070-2. xi + 500 pp.

Bini:1980:SFA

[BL80] Dario Bini and Grazia Lotti.

Stability of fast algorithms for matrix multiplication. Numer. Math., 36:63–72, 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blanch:1964:NEC

[Bla64] G. Blanch. Numerical evaluation of continued fractions. SIAM Rev., 6(4):383–421, 1964.

Bloch:1948:MC

[Blo48] Richard M. Bloch. Mark I calculator. In Harvard [Har48], pages 23–30. Reprinted, with a new introduction by William Aspray, Volume 7 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Bischof:1990:ICEb

[BLP90] Christian H. Bischof, John G. Lewis, and Daniel J. Pierce. Incremental condition estimation for sparse matrices. SIAM J. Matrix Anal. Appl., 11(4):644– 659, 1990.

Bailey:1991:USA

[BLS91] David H. Bailey, King Lee, and Horst D. Simon. Using Strassen’s algorithm to accelerate the solution of linear systems. The Journal of Supercomputing, 4:357–371, 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Blue:1978:PFP

[Blu78] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. ACM Trans. Math. Software, 4(1):15– 23, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bhatia:1994:VUP

[BM94] Rajendra Bhatia and Kalyan Mukherjea. Variation of the unitary part of a matrix. SIAM J. Matrix Anal. Appl., 15(3):1007– 1014, 1994.

Bowdler:1966:SRC

[BMPW66] H. J. Bowdler, R. S. Martin, G. Peters, and J. H. Wilkinson. Solution of real and complex systems of linear equations.

Numer. Math., 8:217–234, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 93– 110], Contribution I/7.

Bjorstad:1992:EMM

[BMSV92] P. Bjørstad, F. Manne, T. Sørevik, and M. Vajteršic. Efficient matrix multiplication on SIMD computers. SIAM J. Matrix Anal. Appl., 13(1):386–401, 1992.

Bargmann:1946:SLS

[BMvN46] V. Bargmann, D. Montgomery, and J. von Neumann. Solution of linear systems of high order. Report prepared for Navy Bureau of Ordnance, 1946. Reprinted in [Tau63, pp. 421– 477].

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell. Computer Structures: Readings and Examples. Mc-Graw-Hill, New York, 1971.

Byers:1987:SVL

[BN87] Ralph Byers and Stephen Nash.

On the singular “vectors” of the Lyapunov operator. SIAM J. Alg. Discrete Methods, 8(1):59– 66, 1987.

Barlow:1988:IME

[BNP88] J. L. Barlow, N. K. Nichols, and R. J. Plemmons. Iterative methods for equalityconstrained least squares problems. SIAM J. Sci. Statist. Comput., 9(5):892–906, 1988.

Bodewig:1949:RRE

[Bod49]  E. Bodewig.  Review of

“Rounding-Off Errors in Matrix Processes” by A. M. Turing. Mathematical Reviews, 10: 405, 1949. CODEN MAREAR. ISSN 0025-5629.

Bohte:1975:BRE

[Boh75] Z. Bohte. Bounds for rounding errors in the Gaussian elimination for band systems. J. Inst. Maths. Applics., 16:133– 142, 1975.

Bohlender:1977:FPC

[Boh77]  Gerd Bohlender.  Floating-

point computation of functions with maximum accuracy. IEEE Trans. Comput., C-26(7):621– 632, 1977.

Bollen:1984:NSD

[Bol84] Jo A. M. Bollen. Numerical stability of descent methods for solving linear equations. Numer. Math., 43:361–377, 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bowden:1953:OTM

[Bow53] B. V. Bowden. The organization of a typical machine. In B. V. Bowden, editor, Faster than Thought: A Symposium on Digital Computing Machines, pages 67–77. Pitman, London, 1953.

Boyd:1974:PMN

[Boy74] David W. Boyd. The power method for ℓ^{p} norms. Linear Algebra Appl., 9:95–101, 1974.

Boyer:1989:HM

[Boy89] Carl B. Boyer. A History of Mathematics. Wiley, New York, second edition, 1989. ISBN 0-471-50357-6. xviii + 762 pp. Revised by Uta C. Merzbach.

Boyle:1994:AFS

[Boy94] Jeff Boyle. An application of Fourier series to the most significant digit problem. Amer. Math. Monthly, 101(11):879–886, 1994.

Ballester:1967:CDA

[BP67] C. Ballester and V. Pereyra. On the construction of discrete approximations to linear differential expressions. Math. Comp., 21:297–302, 1967.

Bjorck:1970:SVS

[BP70] Åke Björck and Victor Pereyra.

Solution of Vandermonde systems of equations. Math. Comp., 24(112):893–903, 1970.

Bunch:1971:DMS

[BP71] James R. Bunch and Beresford N. Parlett. Direct methods for solving symmetric indefinite systems of linear equations. SIAM J. Numer. Anal., 8(4): 639–655, 1971.

Bjorck:1992:LRO

[BP92] Åke Björck and C. C. Paige.

Loss and recapture of orthogonality in the modified Gram– Schmidt algorithm. SIAM J. Matrix Anal. Appl., 13(1):176– 190, 1992.

Berman:1994:NMM

[BP94a] Abraham Berman and Robert J. Plemmons. Nonnegative Matrices in the Mathematical Sciences, volume 9 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-321-8. xx + 340 pp. LCCN QA188 .B47 1994. URL

http://www.loc.gov/catdir/
enhancements/fy0708/94037449-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0708/
94037449-t.html.  Revised

reprint of [?].

Bini:1994:PMC

[BP94b] Dario Bini and Victor Y. Pan. Polynomial and Matrix Computations. Volume 1: Fundamental Algorithms. Birkhäuser, Boston, MA, USA, 1994. ISBN 0-8176-3786-9. xvi + 415 pp.

Bjorck:1994:SAL

[BP94c] Åke Björck and C. C. Paige. Solution of augmented linear systems using orthogonal factorizations. BIT, 34:1–24, 1994. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Boisvert:1997:MMW

[BPR^{+}97] Ronald F. Boisvert, Roldan Pozo, Karin Remington, Richard F.

Barrett, and Jack J. Dongarra. Matrix Market: A Web resource for test matrix collections. In Ronald F. Boisvert, editor, Quality of Numerical Software: Assessment and Enhancement, pages 125–136. Chapman and Hall, London, 1997.

Bauer:1971:IPD

[BR71a] F. L. Bauer and C. Reinsch.

Inversion of positive definite matrices by the Gauss–Jordan method. In J. H. Wilkinson and C. Reinsch, editors, Linear Algebra, volume II of Handbook for Automatic Computation, pages 45–49. Springer-Verlag, Berlin, 1971. Contribution I/3.

Benschop:1971:MSE

[BR71b] N. F. Benschop and H. C. Ratz.

A mean square estimate of the generated roundoff error in constant matrix iterative processes. J. Assoc. Comput. Mach., 18(1): 48–62, 1971.

Bank:1977:MAE

[BR77] Randolph E. Bank and Donald J. Rose. Marching algorithms for elliptic boundary value problems. I: The constant coefficient case. SIAM J. Numer. Anal., 14(5):792–829, 1977.

Bhatia:1997:HWS

[BR97] Rajendra Bhatia and Peter Rosenthal. How and why to solve the operator equation AX − XB = Y. Bulletin of the London Mathematical Society, 29:1–21, 1997. CODEN LMS-BBT. ISSN 0024-6093 (print), 1469-2120 (electronic).

Beaton:1976:ARS

[BRB76] Albert E. Beaton, Donald B. Rubin, and John L. Barone. The acceptability of regression solutions: Another look at computational accuracy. J. Amer. Statist. Assoc., 71(353): 158–168, 1976.

Brent:1970:AMM

[Bre70a] R. P. Brent. Algorithms for matrix multiplication. Technical Report CS 157, Department of Computer Science, Stanford University, Stanford, CA, USA, March 1970. ii + 52 pp.

Brent:1970:EAA

[Bre70b] R. P. Brent. Error analysis of algorithms for matrix multiplication and triangular decomposition using Winograd’s identity. Numer. Math., 16:145– 156, 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brent:1973:PAV

[Bre73] Richard P. Brent. On the precision attainable with various floating-point number systems. IEEE Trans. Comput., C-22(6): 601–607, 1973.

Brent:19xx:FMP

[Bre78a] Richard P. Brent. A Fortran multiple-precision arithmetic package. ACM Trans. Math.

Software, 4(1):57–70, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brent:19xx:AMF

[Bre78b] Richard P. Brent. ALGO-RITHM 524 MP, A Fortran multiple-precision arithmetic package. ACM Trans. Math. Software, 4(1):71–81, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brezinski:1996:ALC

[Bre96] Claude Brezinski. André Louis Cholesky. In A Numerical Analysis Conference in Honour of Jean Meinguet, pages 45–50. 1996. Published as a supplement to the Bulletin of the Belgian Mathematical Society. Also available as Publication ANO-347, Laboratoire d’Analyse Numérique et d’Optimisation, Université des Sciences et Technologies de Lille, Nov. 1995,

http://ano.univ-lille1.fr.

Britton:1992:CWM

[Bri92] J. L. Britton, editor. Collected Works of A. M. Turing: Pure Mathematics. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88059-3. xxii + 287 pp.

Bleher:1987:SFE

[BRK^{+}87] J. H. Bleher, S. M. Rump, U. Kulisch, M. Metzger, Ch. Ullrich, and W. Walter. A study of a FORTRAN extension for engineering/scientific computation with access to ACRITH. Computing: Archiv für informatik und numerik, 39:93–110, 1987. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Brown:1981:SRM

[Bro81] W. S. Brown. A simple but realistic model of floating-point arithmetic. ACM Trans. Math. Software, 7(4):445–480, 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bleher:1985:AHA

[BRR85] J. H. Bleher, A. E. Roeder, and S. M. Rump. ACRITH: Highaccuracy arithmetic. An advanced tool for numerical computation. In Kai Hwang, editor, Proceedings of the 7th Symposium on Computer Arithmetic, pages 318–321. IEEE Computer Society Press, Silver Spring, MD, USA, 1985.

Brunet:1989:CFL

[Bru89] Marie-Christine Brunet. Contribution à la Fiabilité de Logiciels Numériques et à L’analyse de Leur Comportement: Une Approche Statistique. PhD thesis, Université de Paris IX Dauphine, U.E.R. Mathématiques de la Décision, January 1989. viii + 214 pp.

Barnett:1968:SAL

[BS68] S. Barnett and C. Storey. Some applications of the Lyapunov matrix equation. J. Inst. Maths. Applics., 4:33–42, 1968.

Bartels:1972:ASM

[BS72] R. H. Bartels and G. W. Stewart. Algorithm 432: Solution of the matrix equation AX+XB = C. Comm. ACM, 15(9):820– 826, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bailey:1989:FPA

[BSBF89] David H. Bailey, Horst D. Simon, John T. Barton, and Martin J. Fouts. Floating point arithmetic in future supercomputers. Internat. J. Supercomputer Appl., 3(3):86–90, 1989. CODEN IJSAE9. ISSN 0890-2720.

Barron:1960:SSL

[BSD60] D. W. Barron and H. P. F.

Swinnerton-Dyer. Solution of simultaneous linear equations using a magnetic-tape store. Comput. J., 3(1):28–33, 1960. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Bauer:1961:AMN

[BSW61] F. L. Bauer, J. Stoer, and C. Witzgall. Absolute and monotonic norms. Numer. Math., 3:257–264, 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buchanan:1992:NMA

[BT92] James L. Buchanan and Peter R. Turner. Numerical Methods and Analysis. McGraw-Hill, New York, 1992. ISBN 0-07-008717-2, 0-07-112922-7 (international paperback edition). xv + 751 pp.

Buchholz:1959:FFC

[Buc59] W. Buchholz. Fingers or fists? (The choice of decimal or binary representation). Comm. ACM, 2(12):3–11, 1959. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buchholz:1981:OWE

[Buc81] W. Buchholz. Origin of the word byte. Annals of the History of Computing, 3(1):72, 1981.

Bunch:1971:ADP

[Bun71a] James R. Bunch. Analysis of the diagonal pivoting method. SIAM J. Numer. Anal., 8(4): 656–680, 1971.

Bunch:1971:ESM

[Bun71b] James R. Bunch. Equilibration of symmetric matrices in the max-norm. J. Assoc. Comput. Mach., 18(4):566–572, 1971.

Bunch:1974:PPS

[Bun74] James R. Bunch. Partial pivoting strategies for symmetric matrices. SIAM J. Numer. Anal., 11(3):521–528, 1974.

Bunch:1982:NSD

[Bun82] James R. Bunch. A note on the stable decomposition of skew-symmetric matrices. Math. Comp., 38(158):475–479, 1982.

Bunch:1987:WSS

[Bun87] James R. Bunch. The weak and strong stability of algorithms in numerical linear algebra. Linear Algebra Appl., 88/89:49–66, 1987.

Burdakov:1997:GAO

[Bur97] Oleg Burdakov. A greedy algorithm for the optimal basis problem. BIT, 37(3):591– 599, 1997. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Businger:1968:MWC

[Bus68] P. A. Businger. Matrices which can be optimally scaled. Numer. Math., 12:346–348, 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Businger:1971:MNS

[Bus71] P. A. Businger. Monitoring the numerical stability of Gaussian elimination. Numer. Math., 16:360–361, 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Butcher:1987:NAO

[But87] J. C. Butcher. The Numerical Analysis of Ordinary Differential Equations: Runge–Kutta and General Linear Methods.

Wiley, Chichester, UK, 1987. ISBN 0-471-91046-5. xv + 512 pp.

Bischof:1987:WRP

[BV87] Christian H. Bischof and Charles F. Van Loan. The WY representation for products of Householder matrices. SIAM J. Sci. Statist. Comput., 8(1):s2– s13, 1987.

Barlow:1992:NDC

[BV92a] Jesse L. Barlow and Udaya B. Vemulapati. A note on deferred correction for equality constrained least squares problems. SIAM J. Numer. Anal., 29(1):249–256, 1992.

Barlow:1992:RDM

[BV92b] Jesse L. Barlow and Udaya B. Vemulapati. Rank detection methods for sparse matrices. SIAM J. Matrix Anal. Appl., 13 (4):1279–1297, 1992.

Beam:1993:ASB

[BW93] Richard M. Beam and Robert F. Warming. The asymptotic spectra of banded Toeplitz and quasi-Toeplitz matrices. SIAM J. Sci. Comput., 14(4):971– 1006, 1993.

Byers:1984:LSC

[Bye84] Ralph Byers. A LINPACKstyle condition estimator for the equation AX−XB^{T} = C. IEEE Trans. Automat. Control, AC-29 (10):926–928, 1984.

Byers:1985:NCA

[Bye85] Ralph Byers. Numerical condition of the algebraic Riccati equation. In B. N. Datta, editor, Linear Algebra and Its Role in Systems Theory, volume 47 of Contemporary Math., pages 35–49. American Mathematical Society, Providence, RI, USA, 1985.

Byers:1987:SAR

[Bye87] Ralph Byers. Solving the algebraic Riccati equation with the matrix sign function. Linear Algebra Appl., 85:267–279, 1987.

Byers:1988:BMM

[Bye88]  Ralph Byers.  A bisection

method for measuring the distance of a stable matrix to the unstable matrices. SIAM J. Sci. Statist. Comput., 9(5):875–881, 1988.

Barlow:1998:GGE

[BZ98] Jesse L. Barlow and Hongyuan Zha. Growth in Gaussian elimination, orthogonal matrices, and the 2-norm. SIAM J. Matrix Anal. Appl., 19(3):807– 815, 1998.

Caffney:1963:ATM

[Caf63] John Caffney. Another test matrix for determinants and inverses. Comm. ACM, 6(6): 310, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Caprani:1971:ILR

[Cap71]  Ole Caprani.  Implementa-

tion of a low round-off summation method. BIT, 11:271– 275, 1971. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Carr:1959:EAF

[Car59] John W. Carr III. Error analysis in floating point arithmetic. Comm. ACM, 2(5):10–15, 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carter:1991:MFP

[Car91] Russell Carter. Y-MP floating point and Cholesky factorization. International Journal of High Speed Computing, 3(3/4): 215–222, 1991. CODEN IH-SCEZ. ISSN 0129-0533.

Cauchy:1841:EAP

[Cau41] A. Cauchy. Exercices d’Analyse et de Phys. Math., volume 2. Paris, 1841. Cited in [SL54].

Chatelin:1990:PRE

[CB90] Françoise Chatelin and Marie-Christine Brunet. A probabilistic round-off error propagation model. Application to the eigenvalue problem. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 139–160. Oxford University Press, 1990.

Cody:1993:AFS

[CC93] W. J. Cody and Jerome T. Coonen. ALGORITHM 722: Functions to support the IEEE standard for binary floating-point arithmetic. ACM Trans. Math. Software, 19(4):443–451, 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chaitin-Chatelin:1996:LFP

[CCF96] Françoise Chaitin-Chatelin and Valérie Frayssé. Lectures on Finite Precision Computations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-358-7. xv + 235 pp.

Cody:1984:PRW

[CCG^{+}84] W. J. Cody, J. T. Coonen, D. M. Gay, K. Hanson, D. Hough, W. Kahan, R. Karpinski, J. Palmer, F. N. Ris, and D. Stevenson. A proposed radix- and wordlength-independent standard for floating-point arithmetic. IEEE Micro, 4(4):86–100, 1984. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cline:1982:GLC

[CCV82] Alan K. Cline, Andrew R. Conn, and Charles F. Van Loan. Generalizing the LINPACK condition estimator. In J. P. Hennart, editor, Numerical Analysis, Mexico 1981, volume 909 of Lecture Notes in Mathematics, pages 73–83. Springer-Ver-lag, Berlin, 1982.

Carpenter:1986:MTA

[CD86] B. E. Carpenter and R. W. Doran, editors. A. M. Turing’s ACE Report of 1946 and Other Papers, volume 10 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-03114-0. vii + 141 pp.

Conte:1980:ENA

[CdB80] Samuel D. Conte and Carl de Boor. Elementary Numerical Analysis: An Algorithmic Approach. McGraw-Hill, Tokyo, third edition, 1980. ISBN 0-07-066228-2. xii + 432 pp.

Chan:1988:EWC

[CF88] Tony F. Chan and David E. Foulser. Effectively wellconditioned linear systems. SIAM J. Sci. Statist. Comput., 9(6):963–969, 1988.

Chatelin:1992:ECT

[CF92] Françoise Chatelin and Valérie Frayssé. Elements of a condition theory for the computational analysis of algorithms. In R. Beauwens and P. de Groen, editors, Iterative Methods in Linear Algebra, pages 15–25. Elsevier Science Publishers B.V. (North-Holland), Amsterdam, The Netherlands, 1992.

Chartres:1967:CEB

[CG67] Bruce A. Chartres and James C. Geuder. Computable error bounds for direct solution of linear equations. J. Assoc. Comput. Mach., 14(1):63–71, 1967.

Chu:1985:NEE

[CG85] Eleanor Chu and Alan George.

A note on estimating the error in Gaussian elimination without pivoting. ACM SIGNUM Newsletter, 20(2):2–7, 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Char:1991:MVL

[CGG^{+}91] Bruce W. Char, Keith O.

Geddes, Gaston H. Gonnet, Benton L. Leong, Michael B. Monagan, and Stephen M. Watt. Maple V Library Reference Manual. Springer-Verlag, Berlin, 1991. ISBN 3-540-97592-6. xxv + 698 pp.

Chan:1983:ACS

[CGL83] Tony F. Chan, Gene H. Golub, and Randall J. LeVeque. Algorithms for computing the sample variance: Analysis and recommendations. The American Statistician, 37(3):242–247, 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Concus:1985:BPC

[CGM85] P. Concus, G. H. Golub, and G. Meurant. Block preconditioning for the conjugate gradient method. SIAM J. Sci. Statist. Comput., 6(1):220–252,

Conn:1992:LFP

[CGT92] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. LANCELOT: A Fortran Package for Large-Scale Nonlinear Optimization (Release A). Springer-Verlag, Berlin, 1992. ISBN 0-387-55470-X. xviii + 330 pp.

Cox:1990:RNC

[CH90] M. G. Cox and S. J. Hammarling, editors. Reliable Numerical Computation. Oxford University Press, 1990. ISBN 0-19-853564-3. xvi + 339 pp.

Chan:1992:SAR

[CH92a] T. F. Chan and P. C. Hansen. Some applications of the rank revealing QR factorization. SIAM J. Sci. Statist. Comput., 13(3):727–741, 1992.

Croz:1992:SMM

[CH92b] Jeremy J. Du Croz and Nicholas J. Higham. Stability of methods for matrix inversion. IMA J. Numer. Anal., 12:1–19, 1992.

Christiansen:1994:ECN

[CH94] Søren Christiansen and Per Christian Hansen. The effective condition number applied to error analysis of certain boundary collocation methods. J. Comput. Appl. Math., 54:15–36, 1994.

Cheng:1998:MCA

[CH98a] Sheung Hun Cheng and Nicholas J. Higham. A modified Cholesky algorithm based on a symmetric indefinite factorization. SIAM J. Matrix Anal. Appl., 19(4): 1097–1110, October 1998.

Cox:1998:SHQ

[CH98b] Anthony J. Cox and Nicholas J.

Higham. Stability of Householder QR factorization for weighted least squares problems. In D. F. Griffiths, D. J. Higham, and G. A. Watson, editors, Numerical Analysis 1997, Proceedings of the 17th Dundee Biennial Conference, volume 380 of Pitman Research Notes in Mathematics, pages 57–73. Addison Wesley Longman, Harlow, Essex, UK, 1998.

Cox:1999:ASN

[CH99a] Anthony J. Cox and Nicholas J.

Higham. Accuracy and stability of the null space method for solving the equality constrained least squares problem. BIT, 39 (1):34–50, 1999. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cox:1999:BEB

[CH99b] Anthony J. Cox and Nicholas J.

Higham. Backward error bounds for constrained least squares problems. BIT, 39(2):210– 227, 1999. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Cox:1999:RWB

[CH99c] Anthony J. Cox and Nicholas J.

Higham. Row-wise backward stable elimination methods for the equality constrained least squares problem. SIAM J. Matrix Anal. Appl., 21(1):313–326, 1999.

Cheng:2001:ILB

[CH01a] Sheung Hun Cheng and Nicholas J. Higham. Implementation for LAPACK of a block algorithm for matrix 1-norm estimation. Numerical Analysis Report No. 393, Manchester Centre for Computational Mathematics, Manchester, England, August 2001. 19 pp. LAPACK Working Note 152.

Cheng:2001:PIB

[CH01b] Sheung Hun Cheng and Nicholas J. Higham. Parallel implementation of a block algorithm for matrix 1-norm estimation. In Rizos Sakellariou, John Keane, John Gurd, and Len Freeman, editors, Euro-Par 2001, Parallel Processing, volume 2150 of Lecture Notes in Computer Science, pages 568–577. Springer-Verlag, Berlin, 2001.

Chatelin:1993:EM

[Cha93] Françoise Chatelin. Eigenvalues of Matrices. Pure and applied mathematics. Wiley, New York, 1993. ISBN 0-471-93538-7. xviii + 382 pp. LCCN QA188 .C44 1993. URL

http://www.loc.gov/catdir/
enhancements/fy0707/93003430-
d.html;   http://www.loc.
gov/catdir/toc/onix04/93003430.

html. With exercises by Mario Ahués and the author, Translated from the French and with additional material by Walter Ledermann.

Cheng:1998:SIM

[Che98] Sheung Hun Cheng. Symmetric Indefinite Matrices: Linear System Solvers and Modified Inertia Problems. PhD thesis, University of Manchester, Manchester, England, January 1998. 150 pp.

Cohen:1983:CCP

[CHH83] Marty S. Cohen, T. E. Hull, and V. Carl Hamacher. CADAC: A controlled-precision decimal arithmetic unit. IEEE Trans. Comput., C-32(4):370– 377, 1983.

Cheng:2001:ALM

[CHKL01] Sheung Hun Cheng, Nicholas J.

Higham, Charles S. Kenney, and Alan J. Laub. Approximating the logarithm of a matrix to specified accuracy. SIAM J. Matrix Anal. Appl., 22(4):1112– 1125, 2001.

Cornea-Hasegan:1999:IFP

[CHN99] Marius Cornea-Hasegan and Bob Norin. IA-64 floating-point operations and the IEEE standard for binary floating-point arithmetic. Intel Technology Journal, Q4, 1999. http://

developer.intel.com/technology/
itj/.

Cho:1968:TDC

[Cho68] Choong Yun Cho. On the triangular decomposition of Cauchy matrices. Math. Comp., 2(104): 819–825, 1968.

Choi:1983:TTH

[Cho83] Man-Duen Choi. Tricks or treats with the Hilbert matrix. Amer. Math. Monthly, 90:301– 312, 1983.

Chu:1987:SME

[Chu87] King-wah Eric Chu. The solution of the matrix equations AXB − CXD = E and (Y A − DZ, Y C−BZ) = (E, F). Linear Algebra Appl., 93:93–105, 1987.

Chu:1999:CMP

[Chu99] Moody T. Chu. On constructing matrices with prescribed singular values and diagonal elements. Linear Algebra Appl., 288:11–22, 1999.

Chu:2000:FRA

[Chu00] Moody T. Chu. A fast recursive algorithm for constructing matrices with prescribed eigenvalues and singular values. SIAM J. Numer. Anal., 37(3):1004– 1020, 2000.

Chandrasekaran:1994:BEE

[CI94] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. Backward errors for eigenvalue and singular value decompositions. Numer. Math., 68:215–223, 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandrasekaran:1995:SSC

[CI95] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. On the sensitivity of solution components in linear systems of equations. SIAM J. Matrix Anal. Appl., 16 (1):93–112, 1995.

Campbell-Kelly:1981:PPA

[CK81] Martin Campbell-Kelly. Programming the Pilot ACE: Early programming activity at the National Physical Laboratory. Annals of the History of Computing, 3(2):133–162, 1981.

Campbell-Kelly:1984:RAT

[CK84] Martin Campbell-Kelly. Review of “Alan Turing: The Enigma”, by Andrew Hodges. Annals of the History of Computing, 6(2): 176–178, 1984.

Coomes:1995:RCS

[CKP95] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Rigorous computational shadowing of orbits of ordinary differential equations. Numer. Math., 69(4):401–421, 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1979:CSD

[CL79] Tony F. Chan and John Gregg Lewis. Computing standard deivations: Accuracy. Comm. ACM, 22(9):526–531, 1979. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chan:1983:AFS

[CL83a] N. N. Chan and Kim-Hung Li. Algorithm 115: A FOR-TRAN subroutine for finding a real symmetric matrix with prescribed diagonal elements and eigenvalues. Comput. J., 26(2): 184–186, 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Chan:1983:DEE

[CL83b] N. N. Chan and Kim-Hung Li. Diagonal elements and eigenvalues of a real symmetric matrix. Journal of Mathematical Analysis and Applications, 91:562– 566, 1983. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Clenshaw:1955:NSC

[Cle55] C. W. Clenshaw. A note on the summation of Chebyshev series. M.T.A.C., 9(51):118–120, 1955.

Cline:1973:EMS

[Cli73] A. K. Cline. An elimination method for the solution of linear least squares problems. SIAM J. Numer. Anal., 10(2):283–289, 1973.

Clinger:1990:HRF

[Cli90] William D. Clinger. How to read floating point numbers accurately. ACM SIGPLAN Notices, 25(6):92–101, 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Codenotti:2001:RAF

[CLP01] B. Codenotti, M. Leoncini, and F. P. Preparata. The role of arithmetic in fast parallel matrix inversion. Algorithmica, 30 (4):685–707, 2001. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Campbell:1979:GIL

[CM79] S. L. Campbell and C. D. Meyer, Jr. Generalized Inverses of Linear Transformations. Pitman, London, 1979. ISBN 0-486-66693-X. xi + 272 pp. Reprinted by Dover, New York, 1991.

Cline:1979:ECN

[CMSW79] A. K. Cline, C. B. Moler, G. W.

Stewart, and J. H. Wilkinson. An estimate for the condition number of a matrix. SIAM J. Numer. Anal., 16(2):368–375, 1979.

Chan:1996:CGM

[CN96] Raymond H. Chan and Michael K.

Ng. Conjugate gradient methods for Toeplitz systems. SIAM Rev., 38(3):427–482, 1996.

Clenshaw:1984:BFP

[CO84] C. W. Clenshaw and F. W. J.

Olver. Beyond floating point. J. Assoc. Comput. Mach., 31(2): 319–328, 1984.

Cody:1973:SDN

[Cod73] William J. Cody, Jr. Static and dynamic numerical characteristics of floating-point arithmetic. IEEE Trans. Comput., C-22(6): 598–601, 1973.

Cody:1982:ITF

[Cod82] W. J. Cody. Implementation and testing of function software. In Paul C. Messina and Almerico Murli, editors, Problems and Methodologies in Mathematical Software Production, volume 142 of Lecture Notes in Computer Science, pages 24–47. Springer-Verlag, Berlin, 1982.

Cody:1988:AMS

[Cod88a] W. J. Cody. ALGORITHM 665 MACHAR: A subroutine to dynamically determine machine parameters. ACM Trans. Math. Software, 14(4):303–311, 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1988:FPS

[Cod88b] W. J. Cody. Floating-point standards—Theory and practice. In Ramon E. Moore, editor, Reliability in Computing: The Role of Interval Methods in Scientific Computing, pages 99– 107. Academic Press, Boston, MA, USA, 1988.

Cody:1993:ACP

[Cod93] W. J. Cody. Algorithm 714. CELEFUNT: A portable test package for complex elementary functions. ACM Trans. Math. Software, 19(1):121, 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cohen:1974:NPS

[Coh74] A. M. Cohen. A note on pivot size in Gaussian elimination. Linear Algebra Appl., 8: 361–368, 1974.

Cohen:1975:IPM

[Coh75] A. M. Cohen. The inverse of a Pascal matrix. Mathematical Gazette, 59(408):111–112, 1975. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Computer:1988:ANM

[Com88] Apple Computer. Apple Numerics Manual. Addison-Wes-ley, Reading, MA, USA, second edition, 1988.

Computer:1994:IMP

[Com94] Apple Computer. Inside Macintosh: PowerPC Numerics. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-40728-0.

Coonen:1981:UDN

[Coo81] Jerome T. Coonen. Underflow and the denormalized numbers. Computer, 14:75–87, 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

cool90

[Coo90] James W. Cooley. How the FFT gained acceptance. In Stephen G. Nash, editor, A History of Scientific Computing, pages 133–140. Addison-Wesley, Reading, MA, USA, 1990.

Cooley:1994:LFD

[Coo94] James W. Cooley. Lanczos and the FFT: A discovery before its time. In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J.

Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 3–9. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Cottle:1974:MSC

[Cot74] Richard W. Cottle. Manifestations of the Schur complement. Linear Algebra Appl., 8: 189–211, 1974.

Clenshaw:1989:LIA

[COT89] C. W. Clenshaw, F. W. J. Olver, and P. R. Turner. Levelindex arithmetic: An introductory survey. In Peter R. Turner, editor, Numerical Analysis and Parallel Processing, Lancaster 1987, volume 1397 of Lecture Notes in Mathematics, pages 95–168. Springer-Verlag, Berlin, 1989.

Cline:1976:SUL

[CP76] R. E. Cline and R. J. Plemmons. l_{2}-solutions to underdetermined linear systems. SIAM Rev., 18 (1):92–106, 1976.

Chang:1998:SLF

[CP98] Xiao-Wen Chang and Christopher C. Paige. On the sensitivity of the LU factorization. BIT, 38(3):486–501, 1998. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Chang:2001:CPA

[CP01] Xiao-Wen Chang and Christopher C. Paige. Componentwise perturbation analyses for the QR factorization. Numer. Math., 88:319–345, 2001. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chang:1996:NPA

[CPS96] Xiao-Wen Chang, Christopher C. Paige, and G. W. Stewart. New perturbation analyses for the Cholesky factorization. IMA J. Numer. Anal., 16 (4):457–484, 1996.

Chang:1997:PAQ

[CPS97] Xiao-Wen Chang, Christopher C. Paige, and G. W. Stewart. Perturbation analyses for the QR factorization. SIAM J. Matrix Anal. Appl., 18(3):775– 791, 1997.

Curtis:1972:ASM

[CR72] A. R. Curtis and J. K. Reid.

On the automatic scaling of matrices for Gaussian elimination. J. Inst. Maths. Applics., 10:118– 124, 1972.

Cope:1979:BSL

[CR79] J. E. Cope and B. W. Rust.

Bounds on solutions of linear systems with inaccurate data. SIAM J. Numer. Anal., 16(6): 950–963, 1979.

Cline:1983:SCE

[CR83] A. K. Cline and R. K. Rew.

A set of counter-examples to three condition number estimators. SIAM J. Sci. Statist. Comput., 4(4):602–611, 1983.

Calvetti:1992:CVS

[CR92] D. Calvetti and L. Reichel. A Chebychev-Vandermonde solver. Linear Algebra Appl., 172:219–229, 1992.

Calvetti:1993:FIV

[CR93] D. Calvetti and L. Reichel. Fast inversion of Vandermondelike matrices involving orthogonal polynomials. BIT, 33: 473–484, 1993. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Crary:1979:VPN

[Cra79] Fred D. Crary. A versatile precompiler for nonstandard arithmetics. ACM Trans. Math. Software, 5(2):204–217, 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Crout:1941:SME

[Cro41] Prescott D. Crout. A short method for evaluating determinants and solving systems of linear equations with real or complex coefficients. Transactions of the American Institute of Electrical Engineers, 60:1235–1241, 1941. CODEN TAEEA5. ISSN 0096-3860.

Cryer:1968:PSG

[Cry68] Colin W. Cryer. Pivot size in Gaussian elimination. Numer. Math., 12:335–345, 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cryer:1973:FTP

[Cry73] Colin W. Cryer. The LUfactorization of totally positive matrices. Linear Algebra Appl., 7:83–92, 1973.

Cryer:1976:SPT

[Cry76] Colin W. Cryer. Some properties of totally positive matrices. Linear Algebra Appl., 15:1–25, 1976.

Cao:1986:NIH

[CS86] Wei-Lu Cao and William J.

Stewart. A note on inverses of Hessenberg-like matrices. Linear Algebra Appl., 76:233–240, 1986.

Csanky:1976:FPM

[Csa76] L. Csanky. Fast parallel matrix inversion algorithms. SIAM J. Comput., 5(4):618–623, 1976.

Cooley:1965:AMC

[CT65] James W. Cooley and John W.

Tukey. An algorithm for the machine calculation of complex Fourier series. Math. Comp., 19 (90):297–301, 1965.

Cooley:1993:OPF

[CT93] James W. Cooley and John W.

Tukey. On the origin and publication of the FFT paper. Current Contents, (51–52):8–9, 1993.

Cybenko:1986:CME

[CV86] George Cybenko and Charles F.

Van Loan. Computing the minimum eigenvalue of a symmetric positive definite Toeplitz matrix. SIAM J. Sci. Statist. Comput., 7(1):123–131, 1986.

Cody:1980:SME

[CW80] William J. Cody, Jr. and William Waite. Software Manual for the Elementary Functions. Prentice-Hall, Englewood Cliffs, NJ, USA, 1980. ISBN 0-13-822064-6. x + 269 pp.

Coppersmith:1987:MMA

[CW87] Don Coppersmith and Shmuel Winograd. Matrix multiplication via arithmetic progressions. In Proceedings of the Nineteenth Annual ACM Symposium on Theory of Computing, pages 1–6. ACM Press, New York, 1987.

Dahlquist:1983:MMM

[Dah83] G. Dahlquist. On matrix majorants and minorants, with applications to differential equations. Linear Algebra Appl., 52/53:199–216, 1983.

Danloy:1976:CSH

[Dan76] B. Danloy. On the choice of signs for Householder’s matrices. J. Comput. Appl. Math., 2(1):67– 69, 1976.

Datta:1988:MEA

[Dat88] Karabi Datta. The matrix equation XA − BX = R and its applications. Linear Algebra Appl., 109:91–105, 1988.

Davis:1979:CM

[Dav79] Philip J. Davis. Circulant Matrices. Wiley, New York, 1979.

ISBN 0-471-05771-1. xv + 250 pp.

Dax:1982:PPS

[Dax82] Achiya Dax. Partial pivoting strategies for symmetric Gaussian elimination. Math. Programming, 22:288–303, 1982. CODEN MHPGA4. ISSN 0025-5610.

Dax:1990:CLS

[Dax90] Achiya Dax. The convergence of linear stationary iterative processes for solving singular unstructured systems of linear equations. SIAM Rev., 32(4): 611–635, 1990.

Dahlquist:1974:NM

[DBA74] Germund Dahlquist, Åke Björck, and Ned Anderson. Numerical Methods. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1974. ISBN 0-13-627315-7. xviii + 573 pp. LCCN QA297 .D131 1969. Translated by Ned Anderson.

Dongarra:1979:LUG

[DBMS79] J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart. LINPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X.

Boor:1977:BEA

[dBP77] Carl de Boor and Allan Pinkus.

Backward error analysis for totally positive linear systems.

Numer. Math., 27:485–490, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dongarra:19xx:SLB

[DCDH90a] J. J. Dongarra, J. J. Du Croz, I. S. Duff, and S. J. Hammarling. A set of Level 3 basic linear algebra subprograms. ACM Trans. Math. Software, 16:1–17, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:ASL

[DCDH90b] J. J. Dongarra, J. J. Du Croz, I. S. Duff, and S. J. Hammarling. Algorithm 679. A set of Level 3 basic linear algebra subprograms: Model implementation and test programs. ACM Trans. Math. Software, 16:18– 28, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:ESF

[DCHH88a] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. An extended set of Fortran basic linear algebra subprograms. ACM Trans. Math. Software, 14(1):1– 17, 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:AES

[DCHH88b] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. Algorithm 656. An extended set of Fortran basic linear algebra subprograms: Model implementation and test programs. ACM Trans. Math. Software, 14(1): 18–32, 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demmel:1992:DPH

[DDK92] James W. Demmel, J. J. Dongarra, and W. Kahan. On designing portable high performance numerical libraries. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 69–84. Longman Scientific and Technical, Essex, UK, 1992.

Demmel:2001:CCE

[DDM01] James W. Demmel, Benjamin Diament, and Gregorio Malajovich. On the complexity of computing error bounds. Found. Comput. Math., 1(1):101–125, 2001.

Demmel:1995:CSB

[DDR95] James W. Demmel, Inderjit Dhillon, and Huan Ren. On the correctness of some bisectionlike parallel eigenvalue algorithms in floating point arithmetic. Electron. Trans. Numer. Anal., 3:116–149, 1995.

Dongarra:1998:NLA

[DDSvdV98] Jack J. Dongarra, Iain S.

Duff, Danny C. Sorensen, and Henk A. van der Vorst. Numerical Linear Algebra for High-Performance Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-428-1. xviii + 342 pp.

Dekker:1971:FPT

[Dek71] T. J. Dekker. A floatingpoint technique for extending the available precision. Numer. Math., 18:224–242, 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dongarra:2000:RAS

[DEL00] Jack Dongarra, Victor Eijkhout, and Piotr Luszczek. Recursive approach in sparse matrix LU factorization. In Proceedings of the 1st SGI Users Conference, pages 409–418. ACC Cyfronet UMM, Cracow, Poland, October 2000.

Demmel:1983:CNE

[Dem83] James W. Demmel. The condition number of equivalence transformations that block diagonalize matrix pencils. SIAM J. Numer. Anal., 20(3):599–610, 1983.

Demmel:1984:URN

[Dem84] James W. Demmel. Underflow and the reliability of numerical software. SIAM J. Sci. Statist. Comput., 5(4):887–919, 1984.

Demmel:1987:CND

[Dem87a] James W. Demmel. On condition numbers and the distance to the nearest ill-posed problem. Numer. Math., 51:251– 289, 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:1987:EAA

[Dem87b] James W. Demmel. On error analysis in arithmetic with varying relative precision. In Mary Jane Irwin and Renato Stefanelli, editors, Proceedings of the Eighth Symposium on Computer Arithmetic, Como, Italy, pages 148–152. IEEE Computer Society, Washington, D.C., 1987.

Demeure:1989:FFV

[Dem89a] Cédric J. Demeure. Fast QR factorization of Vandermonde matrices. Linear Algebra Appl., 122/3/4:165–194, 1989.

Demmel:1989:FPE

[Dem89b] James W. Demmel. On floating point errors in Cholesky. Technical Report CS-89-87, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 1989. 6 pp. LAPACK Working Note 14.

Demeure:1990:FCV

[Dem90] Cédric J. Demeure. QR factorization of confluent Vandermonde matrices. IEEE Trans. Acoust., Speech, Signal Processing, 38(10):1799–1802, 1990.

Demmel:1991:OIA

[Dem91] James W. Demmel. On the odor of IEEE arithmetic. NA Digest, Volume 91, Issue 39, 1991. (Response to a message “IEEE Arithmetic Stinks” in Volume 91, Issue 33).

demm92

[Dem92a] James W. Demmel. The componentwise distance to the nearest singular matrix. SIAM J. Matrix Anal. Appl., 13(1):10– 19, 1992.

Demmel:1992:OPN

[Dem92b] James W. Demmel. Open problems in numerical linear algebra. IMA Preprint Series #961, Institute for Mathematics and Its Applications, University of Minnesota, Minneapolis, MN, USA, April 1992. 21 pp. LAPACK Working Note 47.

Demmel:1992:SFP

[Dem92c] James W. Demmel. A specification for floating point parallel prefix. Technical Report CS-92-167, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, July 1992. 8 pp. LAPACK Working Note 49.

Demmel:1993:TPN

[Dem93] James W. Demmel. Trading off parallelism and numerical stability. In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 49–68. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Demmel:1997:ANL

[Dem97] James W. Demmel. Applied Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-389-7. xi + 419 pp.

Demmel:1999:ASV

[Dem99] James W. Demmel. Accurate singular value decompositions of structured matrices. SIAM J. Matrix Anal. Appl., 21(2):562– 580, 1999.

Demmel:2001:NAF

[Dem01] James W. Demmel. A note on accurate floating point summation. Manuscript, 2001.

Duff:1986:DMS

[DER86] I. S. Duff, A. M. Erisman, and J. K. Reid. Direct Methods for Sparse Matrices. Clarendon Press, Oxford, UK, 1986. ISBN 0-19-853408-6 (hardcover), 0-19-853421-3 (paperback). xiii + 341 pp. LCCN QA188 .D841 1986. US$37.50.

Descloux:1963:NRE

[Des63] J. Descloux. Note on the roundoff errors in iterative processes. Math. Comp., 17:18–27, 1963.

Dembo:1982:INM

[DES82] Ron S. Dembo, Stanley C.

Eisenstat, and Trond Steihaug. Inexact Newton methods. SIAM J. Numer. Anal., 19(2):400–408, 1982.

Dongarra:1987:DMS

[DG87] Jack J. Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. Comm. ACM, 30(5):403– 407, 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dubrulle:1994:MQI

[DG94] Augustin A. Dubrulle and Gene H. Golub. A multishift QR iteration without computation of the shifts. Numerical Algorithms, 7:173–181, 1994.

Demmel:1999:CSV

[DGE^{+}99] James W. Demmel, Ming Gu, Stanley Eisenstat, Ivan Slapničar, Krešimir Veselić, and Zlatko Drmač. Computing the singular value decomposition with high relative accuracy. Linear Algebra Appl., 299:21–80, 1999.

Dongarra:1984:ILA

[DGK84] J. J. Dongarra, F. G. Gustavson, and A. Karp. Implementing linear algebra algorithms for dense matrices on a vector pipeline machine. SIAM Rev., 26(1):91– 112, 1984.

Daniel:1976:RSA

[DGKS76] J. W. Daniel, W. B. Gragg, L. Kaufman, and G. W. Stewart. Reorthogonalization and stable algorithms for updating the Gram–Schmidt QR factorization. Math. Comp., 30(136): 772–795, 1976.

Duff:1989:SMT

[DGL89] Iain S. Duff, Roger G. Grimes, and John G. Lewis. Sparse matrix test problems. ACM Trans. Math. Software, 15(1):1– 14, 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Duff:1992:UGH

[DGL92] Iain S. Duff, Roger G. Grimes, and John G. Lewis. Users’ guide for the Harwell–Boeing sparse matrix collection (release 1). Report RAL-92-086, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, December 1992. 84 pp.

Duff:1991:FSS

[DGR^{+}91] I. S. Duff, N. I. M. Gould, J. K. Reid, J. A. Scott, and K. Turner. The factorization of sparse symmetric indefinite matrices. IMA J. Numer. Anal., 11: 181–204, 1991.

Drkosova:1995:NSG

[DGRS95] J. Drkošová, A. Greenbaum, M. Rozložník, and Z. Strakos̆. Numerical stability of GMRES. BIT, 35:309–330, 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dekker:1989:RGJ

[DH89] T. J. Dekker and W. Hoffmann. Rehabilitation of the Gauss–Jordan algorithm. Numer. Math., 54:591–599, 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:1992:SBA

[DH92] James W. Demmel and Nicholas J. Higham. Stability of block algorithms with fast level-3 BLAS. ACM Trans. Math. Software, 18(3):274–291, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demmel:1993:IEB

[DH93] James W. Demmel and Nicholas J. Higham. Improved error bounds for underdetermined system solvers. SIAM J. Matrix Anal. Appl., 14(1):1–14, January 1993.

Davies:2000:NSG

[DH00] Philip I. Davies and Nicholas J. Higham. Numerically stable generation of correlation matrices and their factors. BIT, 40(4): 640–651, 2000. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Dhillon:1998:RCC

[Dhi98] Inderjit Dhillon. Reliable computation of the condition number of a tridiagonal matrix in O(n) time. SIAM J. Matrix Anal. Appl., 19(3):776–796, 1998.

Demmel:1995:SBF

[DHS95] James W. Demmel, Nicholas J. Higham, and Robert S. Schreiber.

Stability of block LU factorization. Numerical Linear Algebra with Applications, 2(2):173–190, 1995.

Douglas:1994:GPL

[DHSS94] Craig C. Douglas, Michael Heroux, Gordon Slishman, and Roger M. Smith. GEMMW: A portable level 3 BLAS Winograd variant of Strassen’s matrix– matrix multiply algorithm. J. Comput. Phys., 110:1–10, 1994.

Diamond:1978:SRI

[Dia78] Harold G. Diamond. Stability of rounded off inverses under iteration. Math. Comp., 32(141): 227–232, 1978.

Dixon:1983:EEE

[Dix83] John D. Dixon. Estimating extremal eigenvalues and condition numbers of matrices. SIAM J. Numer. Anal., 20(4):812–814, 1983.

Jong:1977:TFD

[dJ77] Lieuwe Sytse de Jong. Towards a formal definition of numerical stability. Numer. Math., 28:211– 219, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dax:1977:PTS

[DK77] Achiya Dax and S. Kaniel. Pivoting techniques for symmetric Gaussian elimination. Numer. Math., 28:221–241, 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:1985:IAS

[DK85] James W. Demmel and F. Krückeberg. An interval algorithm for solving systems of linear equations to prespecified accuracy. Computing: Archiv für informatik und numerik, 34:117–129, 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Demmel:1987:CSE

[DK87] James W. Demmel and Bo Kågström. Computing stable eigendecompositions of matrix pencils. Linear Algebra Appl., 88/89:139– 186, 1987.

Demmel:1988:ASI

[DK88] James W. Demmel and Bo Kågström. Accurate solutions of ill-posed problems in control theory. SIAM J. Matrix Anal. Appl., 9 (1):126–145, 1988.

Demmel:1990:ASV

[DK90] James W. Demmel and W. Kahan. Accurate singular values of bidiagonal matrices. SIAM J. Sci. Statist. Comput., 11(5): 873–912, 1990.

Demmel:1994:FNA

[DL94] James W. Demmel and Xiaoye Li. Faster numerical algorithms via exception handling. IEEE Trans. Comput., 43(8):983–992, 1994.

Demmel:1989:TMG

[DM89] James W. Demmel and A. McKenney. A test matrix generation suite. Preprint MCS-P69-0389, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, March 1989. 16 pp. LAPACK Working Note 9.

Dongarra:2001:PVC

[Don01]  Jack J. Dongarra.  Perfor-

mance of various computers using standard linear equations software. Technical Report CS-89-85, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 2001. 58 pp.

Douglas:1959:REN

[Dou59] Jim Douglas, Jr. Round-off error in the numerical solution of the heat equation. J. Assoc. Comput. Mach., 6:48–58, 1959.

Drmac:1994:PCF

[DOV94] Zlatko Drmač, Matjaž Omladič, and Krešimir Veselić. On the perturbation of the Cholesky factorization. SIAM J. Matrix Anal. Appl., 15(4):1319–1332, 1994.

Doyle:1957:ISC

[Doy57] Thomas C. Doyle. Inversion of symmetric coefficient matrix of positive-definite quadratic form. M.T.A.C., 11:55–58, 1957.

Day:1988:GGE

[DP88] Jane M. Day and Brian Peterson. Growth in Gaussian elimination. Amer. Math. Monthly, 95(6):489–513, 1988.

Duff:1982:MSF

[DR82] Iain S. Duff and John K. Reid. MA27—A set of Fortran subroutines for solving sparse symmetric sets of linear equations. Technical Report AERE R10533, AERE Harwell Laboratory, July 1982. ISBN 0-7058-0985-4. Published by Her Majesty’s Stationary Office, London.

Davis:1984:MNI

[DR84] Philip J. Davis and Philip Rabinowitz. Methods of Numerical Integration. Academic Press, London, second edition, 1984. ISBN 0-12-206360-0. xiv + 612 pp.

dure95

[DR95] Iain S. Duff and John K. Reid. MA47, a Fortran code for direct solution of indefinite sparse symmetric linear systems. Report RAL-95-001, Atlas Centre, Rutherford Appleton Laboratory, Didcot, Oxon, UK, January 1995. 63 pp.

Duff:1979:DSS

[DRMN79] Iain S. Duff, John K. Reid, Neils Munskgaard, and Hans B. Nielsen. Direct solution of sets of linear equations whose matrix is sparse, symmetric and indefinite. J. Inst. Maths. Applics., 23:235–250, 1979.

Dennis:1983:NMU

[DS83] John E. Dennis, Jr. and Robert B. Schnabel. Numerical Methods for Unconstrained Optimization and Nonlinear Equations. Prentice-Hall, Englewood Cliffs, NJ, USA, 1983. ISBN 0-13-627216-9 (original), 0-89871-364-1 (reprint). xiii + 378 pp. Reprinted by SIAM (1996).

Douglas:1994:VMM

[DS94] Craig C. Douglas and Gordon Slishman. Variants of matrix-matrix multiplication for Fortran-90. ACM SIGNUM Newsletter, 29:4–6, 1994. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dennis:1991:DSM

[DT91] John E. Dennis, Jr. and Virginia Torczon. Direct search methods on parallel machines. SIAM J. Optim., 1(4):448–474, 1991.

Dubrulle:1983:CNM

[Dub83]  Augustin A. Dubrulle.  A

class of numerical methods for the computation of Pythagorean sums. IBM Journal of Research and Development, 27(6): 582–589, 1983. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Dumitrescu:1998:IEA

[Dum98] Bogdan Dumitrescu. Improving and estimating the accuracy of Strassen’s algorithm. Numer. Math., 79:485–499, 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dunham:1990:JTG

[Dun90]  William Dunham.  Journey

Through Genius: The Great Theorems of Mathematics. Penguin, New York, 1990. ISBN 0-14-014739-X. xiii + 300 pp.

Dennis:1984:IQN

[DW84] John E. Dennis, Jr. and Homer F. Walker. Inaccuracy in quasi-Newton methods: Local improvement theorems. Mathematical Programming Study, 22:70–85, 1984. CODEN MP-STDF. ISSN 0303-3929.

Dennis:1993:SLC

[DW93] John E. Dennis, Jr. and Henry Wolkowicz. Sizing and leastchange secant methods. SIAM J. Numer. Anal., 30(5):1291– 1314, 1993.

Dwyer:1951:LC

[Dwy51] Paul S. Dwyer. Linear Computations. Wiley, New York, 1951. xi + 344 pp.

Edelman:1988:ECN

[Ede88] Alan Edelman. Eigenvalues and condition numbers of random matrices. SIAM J. Matrix Anal. Appl., 9(4):543–560, 1988.

Edelman:1991:DMS

[Ede91a] Alan Edelman. The distribution and moments of the smallest eigenvalue of a random matrix of Wishart type. Linear Algebra Appl., 159:55–80, 1991.

Edelman:1991:FAL

[Ede91b] Alan Edelman. The first annual large dense linear system survey. ACM SIGNUM Newsletter, 26:6–12, October 1991. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Edelman:1992:CPC

[Ede92a] Alan Edelman. The complete pivoting conjecture for Gaussian elimination is false. Mathematica Journal, 2:58–61, 1992. ISSN 1047-5974 (print), 1097-1610 (electronic).

Edelman:1992:DSC

[Ede92b] Alan Edelman. On the distribution of a scaled condition number. Math. Comp., 58(197):185– 190, 1992.

Edelman:1993:ERR

[Ede93a]  Alan Edelman.  Eigenvalue

roulette and random test matrices. In Marc S. Moonen, Gene H. Golub, and Bart L. De Moor, editors, Linear Algebra for Large Scale and Real-Time Applications, volume 232 of NATO ASI Series E, pages 365–368. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993.

Edelman:1993:LDN

[Ede93b] Alan Edelman. Large dense numerical linear algebra in 1993: The parallel computing influence. Internat. J. Supercomputer Appl., 7(2):113–128, 1993.

CODEN IJSAE9. ISSN 0890-2720.

Edelman:1994:SDN

[Ede94a] Alan Edelman. Scalable dense numerical linear algebra in 1994: The multicomputer influence. In John G. Lewis, editor, Proceedings of the Fifth SIAM Conference on Applied Linear Algebra, pages 344–348. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Edelman:1994:W

[Ede94b] Alan Edelman. When is x ∗ (1/x) ≠ 1? Manuscript, 1994.

Edelman:1997:MPD

[Ede97] Alan Edelman. The mathematics of the Pentium division bug. SIAM Rev., 39(1):54–67, 1997.

edel91a

[Edi91] Editor’s note. SIAM J. Matrix Anal. Appl., 12(3), 1991.

Ewing:1991:PHC

[EG91] John H. Ewing and F. W. Gehring, editors. Paul Halmos: Celebrating 50 Years of Mathematics. Springer-Verlag, New York, 1991. ISBN 3-540-97509-8. viii + 320 pp.

Egecioglu:1990:PMF

[EGK90] Ömer Eǧecioğlu, E. Gallopoulos, and Ç. K. Koç. A parallel method for fast and practical high-order Newton interpolation. BIT, 30:268–288, 1990. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic).

Erisman:1987:EOU

[EGL^{+}87] A. M. Erisman, R. G. Grimes, J. G. Lewis, W. G. Poole, and H. D. Simon. Evaluation of orderings for unsymmetric sparse matrices. SIAM J. Sci. Statist. Comput., 8(4):600–624, 1987.

Eirola:1993:ABE

[Eir93] Timo Eirola. Aspects of backward error analysis of numerical ODEs. J. Comput. Appl. Math., 45(1-2):65–73, 1993.

Edelman:1994:HME

[EKS94] Alan Edelman, Eric Kostlan, and Michael Shub. How many eigenvalues of a random matrix are real? J. Amer. Math. Soc., 7(1):247–267, 1994.

Elden:1980:PTL

[Eld80] Lars Eldén. Perturbation theory for the least squares problem with linear equality constraints. SIAM J. Numer. Anal., 17(3): 338–350, 1980.

Edelman:1995:CPC

[EM95a] Alan Edelman and Walter Mascarenhas. On the complete pivoting conjecture for a Hadamard matrix of order 12. Linear and Multilinear Algebra, 38(3):181– 188, 1995.

Edelman:1995:PRC

[EM95b] Alan Edelman and H. Murakami. Polynomial roots from companion matrix eigenvalues. Math. Comp., 64(210):763–776, 1995.

Enright:1989:NEC

[Enr89] W. H. Enright. A new errorcontrol for initial value solvers. Appl. Math. Comput., 31:288– 301, 1989.

Epton:1980:MSA

[Ept80] Michael A. Epton. Methods for the solution of AXD − BXC = E and its application in the numerical solution of implicit ordinary differential equations. BIT, 20:341–345, 1980. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Erisman:1974:MST

[ER74] A. M. Erisman and J. K. Reid. Monitoring the stability of the triangular factorization of a sparse matrix. Numer. Math., 22:183–186, 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eldersveld:1992:BUL

[ES92] Samuel K. Eldersveld and Michael A. Saunders. A block-LU update for large-scale linear programming. SIAM J. Matrix Anal. Appl., 13(1):191–201, 1992.

Edelman:1993:SO

[ES93] Alan Edelman and G. W. Stewart. Scaling for orthogonality. IEEE Trans. Signal Processing, 41(4):1676–1677, 1993.

Espelid:1995:FPS

[Esp95] Terje O. Espelid. On floatingpoint summation. SIAM Rev., 37(4):603–607, 1995.

Embree:19xx:PG

[ETxx] Mark Embree and Lloyd N. Trefethen. Pseudospectra gateway,

19xx. http://www.comlab.ox.
ac.uk/pseudospectra/.

Evans:1976:IJH

[Eva76] Christopher Evans. Interview with J. H. Wilkinson. Number 10 in Pioneers of Computing, 60-Minute Recordings of Interviews. Science Museum, London, 1976.

Ewing:1994:CMT

[Ewi94] John Ewing, editor. A Century of Mathematics Through the Eyes of the Monthly. Mathematical Association of America, Washington, D.C., 1994. ISBN 0-88385-459-7. xi + 323 pp.

Eckart:1936:AOM

[EY36] Carl Eckart and Gale Young.

The approximation of one matrix by another of lower rank. Psychometrika, 1(3):211–218, 1936.

Faddeeva:1959:CML

[Fad59] V. N. Faddeeva. Computational Methods of Linear Algebra. Dover, New York, 1959. ISBN 0-486-60424-1. x + 252 pp.

Fallat:2001:BFT

[Fal01] Shaun M. Fallat. Bidiagonal factorizations of totally nonnegative matrices. Amer. Math. Monthly, 108(8):697–712, 2001.

Farebrother:1987:MLM

[Far87] R. W. Farebrother. A memoir of the life of M. H. Doolittle. IMA Bulletin, 23(6/7):102, 1987.

Farebrother:1988:LLS

[Far88a] R. W. Farebrother. Linear Least Squares Computations. Marcel Dekker, New York, 1988. ISBN 0-8247-7661-5. xiii + 293 pp.

Farnum:1988:CSF

[Far88b]  Charles Farnum.  Compiler

support for floating-point computation. Software—Practice and Experience, 18(7):701–709, 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fateman:1982:HLL

[Fat82] Richard J. Fateman. Highlevel language implications of the proposed IEEE floatingpoint standard. ACM Transactions on Programming Languages and Systems, 4(2):239– 257, April 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ferguson:1991:AMA

[FB91] Warren E. Ferguson, Jr. and Tom Brightman. Accurate and monotone approximations of some transcendental functions. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 237– 244. IEEE Computer Society Press, Los Alamitos, CA, USA, 1991.

Ferguson:1995:ECS

[Fer95] Warren E. Ferguson, Jr. Exact computation of a sum or difference with applications to argument reduction. In Simon Knowles and William H. McAllister, editors, Proc. 12th IEEE Symposium on Computer Arithmetic, Bath, England, pages 216–221. IEEE Computer Society Press, Los Alamitos, CA, USA, 1995.

Feldstein:1982:LSF

[FG82] Alan Feldstein and Richard Goodman. Loss of significance in floating point subtraction and addition. IEEE Trans. Comput., C-31(4):328–335, 1982.

Forsgren:1995:CMN

[FGM95] Anders Forsgren, Philip E. Gill, and Walter Murray. Computing modified Newton directions using a partial Cholesky factorization. SIAM J. Sci. Comput., 16 (1):139–150, 1995.

Ferng:1991:ALM

[FGP91] William R. Ferng, Gene H.

Golub, and Robert J. Plemmons. Adaptive Lanczos methods for recursive condition estimation. Numerical Algorithms, 1(1):1–19, 1991.

Forsgren:1996:SSI

[FGS96] Anders Forsgren, Philip E. Gill, and Joseph R. Shinnerl. Stability of symmetric ill-conditioned systems arising in interior methods for constrained optimization. SIAM J. Matrix Anal. Appl., 17(1):187–211, 1996.

Fan:1955:SMI

[FH55] Ky Fan and A. J. Hoffman. Some metric inequalities in the space of matrices. Proc. Amer. Math. Soc., 6:111–116, 1955.

Fox:1978:PMS

[FHS78] P. A. Fox, A. D. Hall, and N. L. Schryer. The PORT mathematical subroutine library. ACM Trans. Math. Software, 4(2): 104–126, 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fox:1948:NSA

[FHW48] L. Fox, H. D. Huskey, and J. H. Wilkinson. Notes on the solution of algebraic linear simultaneous equations. Quart. J. Mech. Appl. Math., 1:149–173, 1948.

Fike:1967:MEP

[Fik67] C. T. Fike. Methods of evaluating polynomial approximations in function evaluation routines. Comm. ACM, 10(3):175– 178, 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fike:1968:CEM

[Fik68] C. T. Fike. Computer Evaluation of Mathematical Functions. Prentice-Hall, Englewood Cliffs, NJ, USA, 1968. xii + 227 pp.

Fischer:1974:FSC

[Fis74] Patrick C. Fischer. Further schemes for combining matrix algorithms. In Jacques Loeckx, editor, Automata, Languages and Programming, volume 14 of Lecture Notes in Computer Science, pages 428–436. Springer-Verlag, Berlin, 1974.

Frigo:1998:FAS

[FJ98] Matteo Frigo and Steven G.

Johnson. FFTW: An adaptive software architecture for the FFT. In Proceedings of International Conference on Acoustics, Speech and Signal Processing, Volume 3, pages 1381 +, 1998.

Fletcher:1976:FSI

[Fle76] R. Fletcher. Factorizing symmetric indefinite matrices. Linear Algebra Appl., 14:257–272, 1976.

Fletcher:1985:EC

[Fle85] R. Fletcher. Expected conditioning. IMA J. Numer. Anal., 5:247–273, 1985.

Fletcher:1987:PMO

[Fle87] R. Fletcher. Practical Methods of Optimization. Wiley, Chichester, UK, second edition, 1987. ISBN 0-471-91547-5. xiv + 436 pp.

Forsythe:1967:CSL

[FM67] George E. Forsythe and Cleve B. Moler. Computer Solution of Linear Algebraic Systems. Prentice-Hall, Englewood Cliffs, NJ, USA, 1967. xi + 148 pp.

Francois:1991:SPM

[FM91] Philippe François and Jean-Michel Muller. The SCALP perturbation method. In Proceedings of IMACS ’91, 13th World Congress on Computation and Applied Mathematics, Dublin, Volume 1, pages 59–60, 1991.

Forsythe:1977:CMM

[FMM77] George E. (George Elmer) Forsythe, Michael A. Malcolm, and Cleve B. Moler. Computer Methods for Mathematical Computations. Prentice-Hall series in automatic computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1977. ISBN 0-13-165332-6. xi + 259 pp. LCCN QA297 .F5681. US$16.95. Cited in Åke Björck’s bibliography on least squares, which is available by anonymous ftp from math.liu.

se in pub/references.

Funderlic:1982:DGD

[FNP82] R. E. Funderlic, M. Neumann, and R. J. Plemmons. LU decompositions of generalized diagonally dominant matrices. Numer. Math., 40:57–69, 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forsythe:1951:GGR

[For51] George E. Forsythe. Gauss to Gerling on relaxation. Mathematical Tables and Other Aids to Computation, 5:255–258, 1951. CODEN MTTCAS. ISSN 0891-6837.

Forsythe:1953:SLA

[For53] George E. Forsythe. Solving linear algebraic equations can be interesting. Bull. Amer. Math. Soc., 59(4):299–329, 1953.

Forsythe:1959:RNR

[For59] George E. Forsythe. Reprint of a note on rounding-off errors. SIAM Rev., 1(1):66–67, 1959.

Forsythe:1960:ACP

[For60] George E. Forsythe. Algorithm 16: Crout with pivoting. Comm. ACM, 3(9):507– 508, September 1960. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Forsythe:1967:TCM

[For67] George E. Forsythe. Today’s computational methods of linear algebra. SIAM Rev., 9:489–515, 1967.

Forsythe:1969:SQE

[For69a] George E. Forsythe. Solving a quadratic equation on a computer. In National Research Council’s Committee on Support of Research in the Mathematical Sciences, editor, The Mathematical Sciences: A Collection of Essays, pages 138– 152. MIT Press, Cambridge, MA, USA, 1969.

Forsythe:1969:WSQ

[For69b] George E. Forsythe. What is a satisfactory quadratic equation solver? In Bruno Dejon and Peter Henrici, editors, Constructive Aspects of the Fundamental Theorem of Algebra, pages 53– 61. Wiley-Interscience, London, 1969.

Forsythe:1970:PCW

[For70] George E. Forsythe. Pitfalls in computation, or why a math book isn’t enough. Amer. Math. Monthly, 77:931–956, 1970.

Ford:1978:PET

[For78]  Brian Ford.  Parameteriza-

tion of the environment for transportable numerical software. ACM Trans. Math. Software, 4(2):100–103, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Foster:1991:MNE

[Fos91] Leslie Foster. Modifications of the normal equations method that are numerically stable. In G. H. Golub and P. M. Van Dooren, editors, Numerical Linear Algebra, Digital Signal Processing and Parallel Algorithms, volume F70 of NATO ASI Series, pages 501–512. Springer-Verlag, Berlin, 1991.

Foster:1994:GEP

[Fos94] Leslie V. Foster. Gaussian elimination with partial pivoting can fail in practice. SIAM J. Matrix Anal. Appl., 15(4):1354– 1362, 1994.

Foster:1997:GFE

[Fos97] Leslie V. Foster. The growth factor and efficiency of Gaussian elimination with rook pivoting. J. Comput. Appl. Math., 86: 177–194, 1997. Corrigendum in J. Comput. Appl. Math., 98:177, 1998.

Fox:1964:INL

[Fox64] L. Fox. An Introduction to Numerical Linear Algebra. Oxford University Press, 1964. xi + 328 pp.

Fox:1971:HGM

[Fox71] L. Fox. How to get meaningless answers in scientific computation (and what to do about it). IMA Bulletin, 7(10):296– 302, 1971.

Fox:1978:AAJ

[Fox78] L. Fox. All about Jim Wilkinson, with a commemorative snippet on backward error analysis. In The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, Symposium Proceedings Series No. 19, pages 1–20. The Institute of Mathematics and Its Applications, Southend-On-Sea, Essex, UK, 1978.

Fox:1987:JHW

[Fox87] L. Fox. James Hardy Wilkinson, 1919–1986. Biographical Memoirs of Fellows of the Royal Society, 33:671–708, 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Fletcher:1974:MF

[FP74] R. Fletcher and M. J. D. Powell. On the modification of LDL^{T} factorizations. Math. Comp., 28 (128):1067–1087, 1974.

Frank:1958:CEC

[Fra58] Werner L. Frank. Computing eigenvalues of complex matrices by determinant evaluation and by methods of Danilewski and Wielandt. J. Soc. Indust. Appl. Math., 6(4):378–392, 1958.

Fraysse:1992:RCS

[Fra92] V. Frayssé. Reliability of Computer Solutions. PhD thesis, L’Institut National Polytechnique de Toulouse, Toulouse, France, July 1992. CERFACS Report TH/PA/92/11.

Friedland:1991:RMS

[Fri91] Shmuel Friedland. Revisiting matrix squaring. Linear Algebra Appl., 154–156:59–63, 1991.

Frommer:2001:PCU

[Fro01] Andreas Frommer. Proving conjectures by use of interval arithmetic. In A. Facius, U. Kulisch, and R. Lohner, editors, Perspectives on Enclosure Methods, pages 1–13. Springer-Ver-lag, New York, 2001.

Forsythe:1955:BCM

[FS55] George E. Forsythe and E. G.

Straus. On best conditioned matrices. Proc. Amer. Math. Soc., 6:340–345, 1955.

Friedland:1980:GPN

[FS80] Shmuel Friedland and Hans Schneider. The growth of powers of a nonnegative matrix. SIAM J. Alg. Discrete Methods, 1(2): 185–200, 1980.

Feldstein:1986:OUS

[FT86] Alan Feldstein and Peter Turner. Overflow, underflow, and severe loss of significance in floating-point addition and subtraction. IMA J. Numer. Anal., 6:241–251, 1986.

Feingold:1962:BDD

[FV62] David G. Feingold and Richard S.

Varga. Block diagonally dominant matrices and generalizations of the Gerschgorin circle theorem. Pacific Journal of Mathematics, 12:1241–1250, 1962. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

Gustafson:1996:MPS

[GA96] John L. Gustafson and Srinivas Aluru. Massively parallel searching for better algorithms or, how to do a cross product with five multiplications. Scientific Programming, pages 203– 213, 1996. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Godunov:1993:GAN

[GAKK93] S. K. Godunov, A. G. Antonov, O. P. Kiriljuk, and V. I. Kostin. Guaranteed Accuracy in Numerical Linear Algebra. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-2352-1. xi + 535 pp.

Gantmacher:1959:TMa

[Gan59a] F. R. Gantmacher. The Theory of Matrices, volume one. Chelsea, New York, 1959. ISBN 0-8284-0131-4. x + 374 pp.

Gantmacher:1959:TMb

[Gan59b] F. R. Gantmacher. The Theory of Matrices, volume two. Chelsea, New York, 1959. ISBN 0-8284-0133-0. ix + 276 pp.

Gardner:1961:MMP

[Gar61] Martin Gardner. More Mathematical Puzzles and Diversions. Penguin, New York, 1961. ISBN 0-14-020748-1. 187 pp.

Garner:1976:SSR

[Gar76] Harvey L. Garner. A survey of some recent contributions to computer arithmetic. IEEE Trans. Comput., C-25(12):1277– 1282, 1976.

Gastinel:1983:LNA

[Gas83] Noel Gastinel. Linear Numerical Analysis. Kershaw Publishing, London, 1983. ISBN 0-901665-16-9. ix + 341 pp. First published in English by Academic Press, London, 1970.

Gautschi:1953:ABP

[Gau53] Walter Gautschi. The asymptotic behaviour of powers of matrices. Duke Mathematical Journal, 20:127–140, 1953. CO-DEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic).

Gautschi:1962:IVC

[Gau62] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. Numer. Math., 4:117–123, 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1975:NEI

[Gau75] Walter Gautschi. Norm estimates for inverses of Vandermonde matrices. Numer. Math., 23:337–347, 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:IVC

[Gau78] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices III. Numer. Math., 29:445–450, 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1983:CVL

[Gau83] Walter Gautschi. The condition of Vandermonde-like matrices involving orthogonal polynomials. Linear Algebra Appl., 52/53:293–300, 1983.

Gautschi:1990:HSV

[Gau90] Walter Gautschi. How (un)stable are Vandermonde systems? In R. Wong, editor, Asymptotic and Computational Analysis, volume 124 of Lecture Notes in Pure and Applied Mathematics, pages 193–210. Marcel Dekker, Inc., New York, NY, USA, 1990.

Gauss:1995:TCO

[Gau95] Carl Friedrich Gauss. Theory of the Combination of Observations Least Subject to Errors. Part One, Part Two, Supplement. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-347-1. xi + 241 pp. Translated from the Latin by G. W. Stewart.

Gay:1990:CRB

[Gay90]  David M. Gay.  Correctly

rounded binary-decimal and decimal-binary conversions. Numerical Analysis Manuscript 90-10, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1990. 16 pp.

Gal:1991:AEM

[GB91] Shmuel Gal and Boris Bachelis. An accurate elementary mathematical library for the IEEE floating point standard. ACM Trans. Math. Software, 17(1): 26–45, 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M. Boyle, J. J. Dongarra, and C. B. Moler. Matrix Eigensystem Routines— EISPACK Guide Extension, volume 51 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1977. ISBN 3-540-08254-9. viii + 343 pp.

Gu:1994:ECS

[GDD94] Ming Gu, James W. Demmel, and Inderjit Dhillon. Efficient computation of the singular value decomposition with applications to least squares problems. Technical Report CS-94-257, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 1994. 19 pp. LAPACK Working Note 88.

Guggenheimer:1995:SEC

[GEJ95] Heinrich W. Guggenheimer, Alan S. Edelman, and Charles R. Johnson. A simple estimate of the condition number of a linear system. College Mathematics Journal, 26(1):2–5, 1995.

Geman:1986:SRL

[Gem86]  Stuart Geman.  The spec-

tral radius of large random matrices. Annals of Probability, 14(4):1318–1328, October 1986. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.
aop/1176992372.

Gentleman:1969:EAG

[Gen69] W. M. Gentleman. An error analysis of Goertzel’s (Watt’s) method for computing Fourier coefficients. Comput. J., 12:160– 165, 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gentleman:1973:LSC

[Gen73] W. Morven Gentleman. Least squares computations by Givens transformations without square roots. J. Inst. Maths. Applics., 12:329–336, 1973.

Gentleman:1975:EAD

[Gen75] W. Morven Gentleman. Error analysis of QR decompositions by Givens transformations. Linear Algebra Appl., 10:189–197, 1975.

Geurts:1982:CTC

[Geu82] A. J. Geurts. A contribution to the theory of condition. Numer. Math., 39:85–96, 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goodman:1975:REP

[GF75] R. Goodman and A. Feldstein. Round-off error in products. Computing: Archiv für informatik und numerik, 15:263–273, 1975. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Goodman:1985:REF

[GFB85] R. H. Goodman, A. Feldstein, and J. Bustoz. Relative error in floating-point multiplication. Computing: Archiv für informatik und numerik, 35:127– 139, 1985. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

George:2002:GFG

[GIK02] Alan George, Khakim D.

Ikramov, and Andrey B. Kucherov. On the growth factor in Gaussian elimination for generalized Higham matrices. Numerical Linear Algebra with Applications, 9:107–114, 2002.

Gill:1951:PSS

[Gil51] S. Gill. A process for the stepby-step integration of differential equations in an automatic digital computing machine. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 47:96– 108, 1951. CODEN PCPSA4. ISSN 0008-1981.

Gillespie:1991:NVT

[Gil91] T. A. Gillespie. Noncommutative variations on theorems of Marcel Riesz and others. In John H. Ewing and F. W. Gehring, editors, Paul Halmos: Celebrating 50 Years of Mathematics, pages 221–236. Springer-Verlag, Berlin, 1991.

Givens:1954:NCC

[Giv54] Wallace J. Givens. Numerical computation of the characteristic values of a real symmetric matrix. Technical Report ORNL-1574, Oak Ridge National Laboratory, Oak Ridge, TN, USA, 1954. 107 pp.

Gregory:1969:CMT

[GK69] Robert T. Gregory and David L. Karney. A Collection of Matrices for Testing Computational Algorithms. Wiley, New York, 1969. ISBN 0-88275-649-4. ix + 154 pp. Reprinted with corrections by Robert E. Krieger, Huntington, New York, 1978.

Gohberg:1993:MCS

[GK93] I. Gohberg and I. Koltracht. Mixed, componentwise, and structured condition numbers. SIAM J. Matrix Anal. Appl., 14 (3):688–704, 1993.

Gudmundsson:1995:SSS

[GKL95] Thorkell Gudmundsson, Charles S. Kenney, and Alan J. Laub. Small-sample statistical estimates for matrix norms. SIAM J. Matrix Anal. Appl., 16(3): 776–792, 1995.

Graham:1994:CM

[GKP94] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-55802-5. xiii + 657 pp. LCCN QA39.2 .G733 1994.

George:1981:CSL

[GL81a] Alan George and Joseph W-H Liu. Computer Solution of Large Sparse Positive Definite Systems. Prentice-Hall, Englewood Cliffs, NJ, USA, 1981.

ISBN 0-13-165274-5. xii + 324 pp.

Grimes:1981:CNE

[GL81b] Roger G. Grimes and John G.

Lewis. Condition number estimation for sparse matrices. SIAM J. Sci. Statist. Comput., 2(4):384–388, 1981.

Gardiner:1991:PAA

[GL91] Judith D. Gardiner and Alan J.

Laub. Parallel algorithms for algebraic Riccati equations. Internat. J. Control, 54(6):1317– 1333, 1991.

Ghavimi:1995:BES

[GL95a] Ali R. Ghavimi and Alan J.

Laub. Backward error, sensitivity, and refinement of computed solutions of algebraic Riccati equations. Numerical Linear Algebra with Applications, 2(1):29–49, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ghavimi:1995:RBD

[GL95b] Ali R. Ghavimi and Alan J.

Laub. Residual bounds for discrete-time Lyapunov equations. IEEE Trans. Automat. Control, 40(7):1244–1249, 1995.

Glanz:1995:MLF

[Gla95] James Glanz. Mathematical logic flushes out the bugs in chip designs. Science, 267:332– 333, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). 20 January.

Gardiner:1992:SSM

[GLAM92] Judith D. Gardiner, Alan J. Laub, James J. Amato, and Cleve B. Moler. Solution of the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Trans. Math. Software, 18(2): 223–231, 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gahinet:1990:SSD

[GLKH90] Pascal M. Gahinet, Alan J. Laub, Charles S. Kenney, and Gary A. Hewer. Sensitivity of the stable discrete-time Lyapunov equation. IEEE Trans. Automat. Control, AC-35(11): 1209–1217, 1990.

Gill:1973:NSF

[GM73] P. E. Gill and W. Murray. A numerically stable form of the simplex algorithm. Linear Algebra Appl., 7:99–138, 1973.

Gentleman:1974:MAR

[GM74] W. Morven Gentleman and Scott B. Marovich. More on algorithms that reveal properties of floating point arithmetic units. Comm. ACM, 17(5):276– 277, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gill:1992:PIS

[GMPS92] Philip E. Gill, Walter Murray, Dulce B. Ponceleón, and Michael A. Saunders. Preconditioners for indefinite systems arising in optimization. SIAM J.

Matrix Anal. Appl., 13(1):292– 311, 1992.

Gill:1986:UGN

[GMSW86] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. User’s guide for NPSOL (version 4.0): A Fortran package for nonlinear programming. Technical Report SOL 86-2, Department of Operations Research, Stanford University, Stanford, CA, USA, January 1986. 53 pp.

Gill:1991:ICM

[GMSW91] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Inertiacontrolling methods for general quadratic programming. SIAM Rev., 33(1):1–36, 1991.

Gill:1981:PO

[GMW81] Philip E. Gill, Walter Murray, and Margaret H. Wright. Practical Optimization. Academic Press, London, 1981. ISBN 0-12-283952-8. xvi + 401 pp.

Golub:1979:HSM

[GNV79] G. H. Golub, S. Nash, and C. F.

Van Loan. A Hessenberg–Schur method for the problem AX + XB = C. IEEE Trans. Automat. Control, AC-24(6):909– 913, 1979.

Gohberg:1994:FIC

[GO94] I. Gohberg and V. Olshevsky.

Fast inversion of Chebyshev– Vandermonde matrices. Numer.

Math., 67(1):71–92, 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1962:BRE

[Gol62] Gene H. Golub. Bounds for the round-off errors in the Richardson second order method. BIT, 2:212–223, 1962. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Golub:1965:NMS

[Gol65] G. H. Golub. Numerical methods for solving linear least squares problems. Numer. Math., 7:206–216, 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldberg:1967:BED

[Gol67] I. Bennett Goldberg. 27 bits are not enough for 8-digit accuracy. Comm. ACM, 10(2): 105–106, February 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldstine:1972:CPN

[Gol72] Herman H. Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0. xii + 378 pp. 1993 printing with new preface.

Goldstine:1977:HNA

[Gol77] Herman H. Goldstine. A History of Numerical Analysis from the 16th through the 19th Century. Springer-Verlag, New York, 1977. ISBN 0-387-90277-5. xiv + 348 pp.

Goldberg:1991:WEC

[Gol91] David Goldberg. What every computer scientist should know about floating-point arithmetic. ACM Computing Surveys, 23(1):5–48, 1991. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Goodnight:1979:TSO

[Goo79] James H. Goodnight. A tutorial on the SWEEP operator. The American Statistician, 33 (3):149–158, 1979. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gould:1991:GGE

[Gou91] N. I. M. Gould. On growth in Gaussian elimination with complete pivoting. SIAM J. Matrix Anal. Appl., 12(2):354–361, 1991.

Govaerts:2000:NMB

[Gov00] Willy J. F. Govaerts. Numerical Methods for Bifurcations of Dynamical Equilibria. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-442-7. xxii + 362 pp.

Govaerts:1990:BEO

[GP90] W. Govaerts and J. D. Pryce.

Block elimination with one iterative refinement solves bordered linear systems accurately. BIT, 30:490–507, 1990. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gasca:1992:TPN

[GP92] M. Gasca and J. M. Peña. Total positivity and Neville elimination. Linear Algebra Appl., 165:25–44, 1992.

Gallivan:1990:PAD

[GPS90] K. A. Gallivan, R. J. Plemmons, and A. H. Sameh. Parallel algorithms for dense linear algebra computations. SIAM Rev., 32 (1):54–135, 1990.

Granville:1992:ZBB

[Gra92] Andrew Granville. Zaphod Beeblebrox’s brain and the fiftyninth row of Pascal’s triangle. Amer. Math. Monthly, 99:318– 331, 1992.

Greenbaum:1989:BSP

[Gre89] Anne Greenbaum. Behavior of slightly perturbed Lanczos and conjugate-gradient recurrences. Linear Algebra Appl., 113:7–63, 1989.

Greenbaum:1994:LCG

[Gre94] Anne Greenbaum. The Lanczos and conjugate gradient algorithms in finite precision arithmetic. In J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors, Proceedings of the Cornelius Lanczos International Centenary Conference, pages 49–60. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994.

Greenbaum:1997:EAA

[Gre97a] Anne Greenbaum. Estimating the attainable accuracy of recursively computed residual methods. SIAM J. Matrix Anal. Appl., 18(3):535–551, 1997.

Greenbaum:1997:IMS

[Gre97b] Anne Greenbaum. Iterative Methods for Solving Linear Systems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-396-X. xiii + 220 pp.

Griewank:1985:SNE

[Gri85] Andreas Griewank. On solving nonlinear equations with simple singularities or nearly singular solutions. SIAM Rev., 27(4): 537–563, 1985.

Gentleman:1966:FFT

[GS66] W. M. Gentleman and G. Sande.

Fast Fourier transforms—for fun and profit. In Fall Joint Computer Conference, volume 29 of AFIPS Conference Proceedings, pages 563–578. Spartan Books, Washington, D.C., 1966.

Gragg:1976:SVS

[GS76] W. B. Gragg and G. W. Stewart. A stable variant of the secant method for solving nonlinear equations. SIAM J. Numer. Anal., 13(6):889–903, 1976.

Goldberg:1983:MNM

[GS83] Moshe Goldberg and E. G. Straus. Multiplicativity of l_{p} norms for matrices. Linear Algebra Appl., 52/53:351–360, 1983.

Gluchowska:1990:SLL

[GS90] J. Gluchowska and A. Smoktunowicz. Solving the linear least squares problem with very high relative accuracy. Computing: Archiv für informatik und numerik, 45:345–354, 1990. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Greenbaum:1992:PBF

[GS92] Anne Greenbaum and Zdenĕk Strakos̆. Predicting the behavior of finite precision Lanczos and conjugate gradient computations. SIAM J. Matrix Anal. Appl., 13(1):121–137, 1992.

Gill:1996:SCF

[GSS96] Philip E. Gill, Michael A. Saunders, and Joseph R. Shinnerl. On the stability of Cholesky factorization for symmetric quasidefinite systems. SIAM J. Matrix Anal. Appl., 17 (1):35–46, 1996.

Gu:1998:BPB

[Gu98] Ming Gu. Backward perturbation bounds for linear least squares problems. SIAM J. Matrix Anal. Appl., 20(2):363–372, 1998.

Gulliksson:1994:IRC

[Gul94] Mårten Gulliksson. Iterative refinement for constrained and weighted linear least squares. BIT, 34:239–253, 1994. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Gulliksson:1995:BEA

[Gul95] Mårten Gulliksson. Backward error analysis for the constrained and weighted linear least squares problem when using the weighted QR factorization. SIAM J. Matrix Anal. Appl., 16(2):675–687, 1995.

Gustavson:1997:RLA

[Gus97]  F. G. Gustavson.  Recur-

sion leads to automatic variable blocking for dense linear-algebra algorithms. IBM Journal of Research and Development, 41(6): 737–755, November 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Golub:1974:CBS

[GV74] G. H. Golub and J. M. Varah.

On a characterization of the best ℓ_{2}-scaling of a matrix. SIAM J. Numer. Anal., 11(3):472–479, 1974.

Golub:1979:UPD

[GV79] Gene H. Golub and Charles F.

Van Loan. Unsymmetric positive definite linear systems. Linear Algebra Appl., 28:85–97,

Golub:1996:MC

[GV96] Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins Studies in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, third edition, 1996. ISBN 0-8018-5413-X (hardcover), 0-8018-5414-8 (paperback). xxvii + 694 pp. LCCN QA188 .G65 1996. URL

http://www.loc.gov/catdir/
bios/jhu052/96014291.html;
 http://www.loc.gov/catdir/
description/jhu051/96014291.
html.

Goldstine:1951:NIM

[GvN51] Herman H. Goldstine and John von Neumann. Numerical inverting of matrices of high order II. Proc. Amer. Math. Soc., 2:188–202, 1951. Reprinted in [Tau63, pp. 558–572].

Golub:1966:NIR

[GW66] G. H. Golub and J. H. Wilkinson. Note on the iterative refinement of least squares solution. Numer. Math., 9:139–148, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1976:ICE

[GW76] G. H. Golub and J. H. Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. SIAM Rev., 18(4):578–619, 1976.

Gulliksson:1992:MQD

[GW92] Mårten Gulliksson and Per-Åke Wedin. Modifying the QRdecomposition to constrained and weighted linear least squares. SIAM J. Matrix Anal. Appl., 13(4):1298–1313, 1992.

Gardiner:1992:AFS

[GWL^{+}92] Judith D. Gardiner, Matthew R.

Wette, Alan J. Laub, James J. Amato, and Cleve B. Moler. Algorithm 705: A FORTRAN-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Trans. Math. Software, 18(2): 232–238, 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hull:1985:NT

[HAC^{+}85] T. E. Hull, A. Abraham, M. S.

Cohen, A. F. X. Curley, C. B. Hall, D. A. Penny, and J. T. M. Sawchuk. Numerical TUR-ING. ACM SIGNUM Newsletter, 20(3):26–34, 1985. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hager:1984:CE

[Hag84] William W. Hager. Condition estimates. SIAM J. Sci. Statist. Comput., 5(2):311–316, 1984.

Hall:1967:CT

[Hal67] Marshall Hall, Jr. Combinatorial Theory. Blaisdell, Waltham, MA, USA, 1967. x + 310 pp.

Hamming:1973:NMS

[Ham73] R. W. Hamming. Numerical Methods for Scientists and Engineers. McGraw-Hill, New York, second edition, 1973. ISBN 0-07-025887-2. ix + 721 pp.

Hammarling:1974:NMG

[Ham74] Sven Hammarling. A note on modifications to the Givens plane rotation. J. Inst. Maths. Applics., 13:215–218, 1974.

Hammarling:1982:NSS

[Ham82] S. J. Hammarling. Numerical solution of the stable, non-negative definite Lyapunov equation. IMA J. Numer. Anal., 2:303–323, 1982.

Hamada:1987:NRN

[Ham87a] Hozumi Hamada. A new real number representation and its operation. In Mary Jane Irwin and Renato Stefanelli, editors, Proceedings of the Eighth Symposium on Computer Arithmetic, Como, Italy, pages 153– 157. IEEE Computer Society, Washington, D.C., 1987.

Hammarling:1987:NSG

[Ham87b] Sven J. Hammarling. The numerical solution of the general Gauss–Markov linear model. In T. S. Durrani, J. B. Abbiss, and J. E. Hudson, editors, Mathematics in Signal Processing, pages 451–456. Oxford University Press, 1987.

Hanson:1997:AML

[Han97] Richard J. Hanson. Aasen’s method for linear systems with self-adjoint matrices. Visual Numerics, Inc., http://www.

vni.com/books/whitepapers/
Aasen.html, July 1997.

harv46

[Har46] A Manual of Operation for the Automatic Sequence Controlled Calculator. Harvard University Press, Cambridge, MA, USA, 1946. Reprinted, with new foreword and introduction, Volume 8 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

harv48

[Har48] Proceedings of a Symposium on Large-Scale Digital Calculating Machinery, volume 16 of The Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1948. Reprinted, with a new introduction by William Aspray, Volume 7 in the Charles Babbage Institute Reprint Series for the History of Computing, MIT Press, Cambridge, MA, USA, 1985.

Hardy:1967:CPM

[Har67] G. H. Hardy. A Course of Pure Mathematics. Cambridge University Press, tenth edition, 1967. ISBN 0-521-09227-2. xii + 509 pp.

Harter:1972:OWF

[Har72] Richard Harter. The optimality of Winograd’s formula. Comm. ACM, 15(5):352, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hartfiel:1980:CSS

[Har80] D. J. Hartfiel. Concerning the solution set of Ax = b where P ≤ A ≤ Q and p ≤ b ≤ q. Numer. Math., 35:355–359, 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hauser:1996:HFP

[Hau96] John R. Hauser. Handling floating-point exceptions in numeric programs. ACM Trans. Program. Lang. Syst., 18(2): 139–174, 1996.

Holt:1988:TPL

[HC88] R. C. Holt and J. R. Cordy. The Turing programming language. Comm. ACM, 31(12): 1410–1423, 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hearon:1977:NS

[Hea77] John Z. Hearon. Nonsingular solutions of T A−BT = C. Linear Algebra Appl., 16:57–63, 1977.

Heath:1984:NML

[Hea84] Michael T. Heath. Numerical methods for large sparse linear least squares problems. SIAM J.

Sci. Statist. Comput., 5(3):497– 513, 1984.

Hein:1992:G

[Hei92] Piet Hein. Grooks. Number 85 in Borgens Pocketbooks. Borgens Forlag, Copenhagen, Denmark, second edition, 1992. ISBN 87-418-1079-1. 53 pp.

Henrici:1962:BII

[Hen62a] Peter Henrici. Bounds for iterates, inverses, spectral variation and fields of values of non-normal matrices. Numer. Math., 4:24–40, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henrici:1962:DVM

[Hen62b] Peter Henrici. Discrete Variable Methods in Ordinary Differential Equations. John Wiley, New York, 1962. xi + 407 pp.

Henrici:1963:EPD

[Hen63] Peter Henrici. Error Propagation for Difference Methods. John Wiley, New York, 1963. vi + 73 pp.

Henrici:1964:ENA

[Hen64] Peter Henrici. Elements of Numerical Analysis. Wiley, New York, 1964. xv + 328 pp.

Henrici:1966:TPM

[Hen66] Peter Henrici. Test of probabilistic models for the propagation of roundoff errors. Comm. ACM, 9(6):409–410, 1966. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henrici:1980:MPR

[Hen80] Peter Henrici. A model for the propagation of rounding error in floating arithmetic. In Karl L. E. Nickel, editor, Interval Mathematics 1980, pages 49–73. Academic Press, New York, 1980.

hp82

[Hew82] HP-15C Advanced Functions Handbook. Hewlett–Packard, Portable Computer Division, Corvallis, OR, USA, 1982. 221 pp. Part number 00015-90011 Rev. C.

hp93

[Hew93] HP 48G Series User’s Guide. Hewlett–Packard, Corvallis Division, Corvallis, OR, USA, 1993. Part number 00048-90126, Edition 3.

Hull:1994:ICE

[HFT94] T. E. Hull, Thomas F. Fairgrieve, and Ping Tak Peter Tang. Implementing complex elementary functions using exception handling. ACM Trans. Math. Software, 20(2):215–244, 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Heath:1990:EEI

[HGD90] M. T. Heath, G. A. Geist, and J. B. Drake. Early experience with the Intel iPSC/860 at Oak Ridge National Laboratory. Report ORNL/TM-11655, Oak Ridge National Laboratory, Oak Ridge, TN, USA, September 1990. 26 pp.

Higham:1989:LGF

[HH89] Nicholas J. Higham and Desmond J.

Higham. Large growth factors in Gaussian elimination with pivoting. SIAM J. Matrix Anal. Appl., 10(2):155–164, April 1989.

Higham:1992:BEC

[HH92a] Desmond J. Higham and Nicholas J. Higham. Backward error and condition of structured linear systems. SIAM J. Matrix Anal. Appl., 13(1):162– 175, January 1992.

Higham:1992:CPT

[HH92b] Desmond J. Higham and Nicholas J. Higham. Componentwise perturbation theory for linear systems with multiple right-hand sides. Linear Algebra Appl., 174:111–129, 1992.

Higham:2000:MG

[HH00] Desmond J. Higham and Nicholas J. Higham. MAT-LAB Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-516-4. xxii + 283 pp.

Hammer:1993:NTV

[HHKR93] Rolf Hammer, Matthias Hocks, Ulrich Kulisch, and Dietmar Ratz. Numerical Toolbox for Verified Computing I. Basic Numerical Problems: Theory, Algorithms, and Pascal-XSC Programs. Springer-Verlag, Berlin, 1993. ISBN 3-540-57118-3. xiii + 337 pp.

Higham:1986:CPD

[Hig86a] Nicholas J. Higham. Computing the polar decomposition— with applications. SIAM J. Sci. Statist. Comput., 7(4):1160– 1174, October 1986.

Higham:1986:EAC

[Hig86b] Nicholas J. Higham. Efficient algorithms for computing the condition number of a tridiagonal matrix. SIAM J. Sci. Statist. Comput., 7(1):150–165, 1986.

Higham:1987:CRS

[Hig87a] Nicholas J. Higham. Computing real square roots of a real matrix. Linear Algebra Appl., 88/89:405–430, 1987.

Higham:1987:EAB

[Hig87b] Nicholas J. Higham. Error analysis of the Björck–Pereyra algorithms for solving Vandermonde systems. Numer. Math., 50(5):613–632, 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1987:SCN

[Hig87c] Nicholas J. Higham. A survey of condition number estimation for triangular matrices. SIAM Rev., 29(4):575–596, December 1987.

Higham:1988:CNS

[Hig88a] Nicholas J. Higham. Computing a nearest symmetric positive semidefinite matrix. Linear Algebra Appl., 103:103–118, 1988.

Higham:1988:FSV

[Hig88b] Nicholas J. Higham. Fast solution of Vandermonde-like systems involving orthogonal polynomials. IMA J. Numer. Anal., 8:473–486, 1988.

Higham:1988:FCE

[Hig88c] Nicholas J. Higham. FOR-TRAN codes for estimating the one-norm of a real or complex matrix, with applications to condition estimation (Algorithm 674). ACM Trans. Math. Software, 14(4):381–396, December 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1989:AST

[Hig89a] Nicholas J. Higham. The accuracy of solutions to triangular systems. SIAM J. Numer. Anal., 26(5):1252–1265, October 1989.

Higham:1989:MNP

[Hig89b] Nicholas J. Higham. Matrix nearness problems and applications. In M. J. C. Gover and S. Barnett, editors, Applications of Matrix Theory, pages 1–27. Oxford University Press, 1989.

Higham:1990:ACD

[Hig90a] Nicholas J. Higham. Analysis of the Cholesky decomposition of a semi-definite matrix. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 161–185. Oxford University Press, 1990.

Higham:1990:BEG

[Hig90b] Nicholas J. Higham. Bounding the error in Gaussian elimination for tridiagonal systems. SIAM J. Matrix Anal. Appl., 11 (4):521–530, October 1990.

Higham:1990:CEB

[Hig90c] Nicholas J. Higham. Computing error bounds for regression problems. In Philip J. Brown and Wayne A. Fuller, editors, Statistical Analysis of Measurement Error Models and Applications, Contemporary Mathematics 112, pages 195–208. American Mathematical Society, Providence, RI, USA, 1990.

Higham:1990:EMN

[Hig90d] Nicholas J. Higham. Experience with a matrix norm estimator. SIAM J. Sci. Statist. Comput., 11(4):804–809, July 1990.

Higham:1990:EFM

[Hig90e] Nicholas J. Higham. Exploiting fast matrix multiplication within the level 3 BLAS. ACM Trans. Math. Software, 16(4): 352–368, December 1990. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Higham:1990:HAG

[Hig90f] Nicholas J. Higham. How accurate is Gaussian elimination? In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1989, Proceedings of the 13th Dundee Conference, volume 228 of Pitman Research Notes in Mathematics, pages 137–154. Longman Scientific and Technical, Essex, UK, 1990.

Higham:1990:IRE

[Hig90g] Nicholas J. Higham. Iterative refinement enhances the stability of QR factorization methods for solving linear equations. Numerical Analysis Report No. 182, University of Manchester, Manchester, England, April 1990.

Higham:1990:SAA

[Hig90h] Nicholas J. Higham. Stability analysis of algorithms for solving confluent Vandermondelike systems. SIAM J. Matrix Anal. Appl., 11(1):23–41, January 1990.

Higham:1991:IRE

[Hig91a] Nicholas J. Higham. Iterative refinement enhances the stability of QR factorization methods for solving linear equations. BIT, 31:447–468, 1991. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:1991:TMP

[Hig91b] Nicholas J. Higham. Three measures of precision in floating point arithmetic. NA Digest, Volume 91, Issue 16, 1991. Electronic mail magazine: na.help@

na-net.ornl.gov.

Higham:1992:EMN

[Hig92a] Nicholas J. Higham. Estimating the matrix p-norm. Numer. Math., 62:539–555, 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1992:SMM

[Hig92b] Nicholas J. Higham. Stability of a method for multiplying complex matrices with three real matrix multiplications. SIAM J. Matrix Anal. Appl., 13(3):681– 687, July 1992.

Higham:1993:AFP

[Hig93a] Nicholas J. Higham. The accuracy of floating point summation. SIAM J. Sci. Comput., 14 (4):783–799, July 1993.

Higham:1993:ODS

[Hig93b] Nicholas J. Higham. Optimization by direct search in matrix computations. SIAM J. Matrix Anal. Appl., 14(2):317–333, April 1993.

Higham:1993:PTB

[Hig93c] Nicholas J. Higham. Perturbation theory and backward error for AX − XB = C. BIT, 33: 124–136, 1993. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:1994:MSD

[Hig94a] Nicholas J. Higham. The matrix sign decomposition and its relation to the polar decomposition.

Linear Algebra Appl., 212/213: 3–20, 1994.

Higham:1994:SCP

[Hig94b] Nicholas J. Higham. A survey of componentwise perturbation theory in numerical linear algebra. In Walter Gautschi, editor, Mathematics of Computation 1943–1993: A Half Century of Computational Mathematics, volume 48 of Proceedings of Symposia in Applied Mathematics, pages 49–77. American Mathematical Society, Providence, RI, USA, 1994.

Higham:1995:CNT

[Hig95a] Desmond J. Higham. Condition numbers and their condition numbers. Linear Algebra Appl., 214:193–213, 1995.

Higham:1995:SPT

[Hig95b] Nicholas J. Higham. Stability of parallel triangular system solvers. SIAM J. Sci. Comput., 16(2):400–413, March 1995.

Higham:1995:TMT

[Hig95c] Nicholas J. Higham. The Test Matrix Toolbox for MATLAB (version 3.0). Numerical Analysis Report No. 276, Manchester Centre for Computational Mathematics, Manchester, England, September 1995. 70 pp.

Higham:1997:IRL

[Hig97a] Nicholas J. Higham. Iterative refinement for linear systems and LAPACK. IMA J. Numer. Anal., 17(4):495–509, 1997.

Higham:1997:SDP

[Hig97b] Nicholas J. Higham. Stability of the diagonal pivoting method with partial pivoting. SIAM J. Matrix Anal. Appl., 18(1):52– 65, January 1997.

Higham:1997:SIM

[Hig97c] Nicholas J. Higham. Stable iterations for the matrix square root. Numerical Algorithms, 15 (2):227–242, 1997.

Higham:1998:FCS

[Hig98a] Nicholas J. Higham. Factorizing complex symmetric matrices with positive definite real and imaginary parts. Math. Comp., 67(224):1591–1599, October 1998.

Higham:1998:HWM

[Hig98b] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1998. ISBN 0-89871-420-6. xvi + 302 pp.

Higham:1999:NAS

[Hig99a] Nicholas J. Higham. Notes on accuracy and stability of algorithms in numerical linear algebra. In Mark Ainsworth, Jeremy Levesley, and Marco Marletta, editors, The Graduate Student’s Guide to Numerical Analysis ’98, pages 48–82. Springer-Ver-lag, Berlin, 1999.

Higham:1999:SBF

[Hig99b] Nicholas J. Higham. Stability of block LDL^{T} factorization of a symmetric tridiagonal matrix. Linear Algebra Appl., 287:181– 189, 1999.

Higham:2000:QFC

[Hig00] Nicholas J. Higham. QR factorization with complete pivoting and accurate computation of the SVD. Linear Algebra Appl., 309: 153–174, 2000.

Higham:2001:EPA

[Hig01] Nicholas J. Higham. Evaluating Padé approximants of the matrix logarithm. SIAM J. Matrix Anal. Appl., 22(4):1126– 1135, 2001.

Higham:2002:ASN

[Hig02] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2002. ISBN 0-89871-521-0. xxx + 680 pp.

Hilbert:1894:BTL

[Hil94] David Hilbert. Ein Beitrag zur Theorie des Legendre’schen Polynoms. Acta Mathematica, 18:155–159, 1894. CODEN AC-MAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Hildebrand:1974:INA

[Hil74] F. B. Hildebrand. Introduction to Numerical Analysis. Mc-Graw-Hill, New York, second edition, 1974. ISBN 0-486-65363-3. xiii + 669 pp. Reprinted by Dover, New York, 1987.

Horn:1985:MA

[HJ85] Roger A. Horn and Charles R. Johnson. Matrix Analysis. Cambridge University Press, 1985. ISBN 0-521-30586-1. xiii + 561 pp.

Hockney:1988:PCA

[HJ88] Roger W. Hockney and Chris R. Jesshope. Parallel Computers: Architecture, Programming, and Algorithms. Adam Hilger Ltd., Bristol, UK, second edition, 1988. ISBN 0-85274-811-6 (hardcover), 0-85274-812-4 (paperback). xv + 625 pp. LCCN QA76.5 .H57 1988.

Horn:1991:TMA

[HJ91] Roger A. Horn and Charles R. Johnson. Topics in Matrix Analysis. Cambridge University Press, 1991. ISBN 0-521-30587-X. viii + 607 pp.

Hewer:1988:SSL

[HK88] Gary Hewer and Charles Kenney. The sensitivity of the stable Lyapunov equation. SIAM J. Control Optim., 26(2):321–344, 1988.

Higham:1993:CEA

[HK93a] Nicholas J. Higham and Philip A. Knight. Componentwise error analysis for stationary iterative methods. In Carl D. Meyer and Robert J. Plemmons, editors, Linear Algebra, Markov Chains, and Queueing Models, volume 48 of IMA Volumes in Mathematics and Its Applications, pages 29–46. Springer-Verlag, New York, 1993.

Higham:1993:FPB

[HK93b] Nicholas J. Higham and Philip A.

Knight. Finite precision behavior of stationary iteration for solving singular systems. Linear Algebra Appl., 192:165–186, 1993.

Higham:1995:MPF

[HK95] Nicholas J. Higham and Philip A.

Knight. Matrix powers in finite precision arithmetic. SIAM J. Matrix Anal. Appl., 16(2):343– 358, April 1995.

Higham:2000:NAQ

[HK00] Nicholas J. Higham and Hyun-Min Kim. Numerical analysis of a quadratic matrix equation. IMA J. Numer. Anal., 20(4): 499–519, 2000.

Hida:2001:AQD

[HLB01] Yozo Hida, Xiaoye S. Li, and David H. Bailey. Algorithms for quad-double precision floating point arithmetic. In Proc. 15th IEEE Symposium on Computer Arithmetic, pages 155– 162. IEEE Computer Society Press, Los Alamitos, CA, USA, 2001.

Hardy:1952:I

[HLP52] G. H. Hardy, J. E. Littlewood, and G. Pólya. Inequalities. Cambridge University Press, second edition, 1952. xii + 324 pp.

Hodges:1983:ATE

[Hod83] Andrew Hodges. Alan Turing: The Enigma. Burnett Books, London, 1983. ISBN 0-09-911641-3. xix + 586 pp. 1992 edition with preface, Vintage, London.

Hodel:1992:RAL

[Hod92] A. Scottedward Hodel. Recent applications of the Lyapunov equation in control theory. In R. Beauwens and P. de Groen, editors, Iterative Methods in Linear Algebra, pages 217–227. Elsevier Science Publishers B.V. (North-Holland), Amsterdam, The Netherlands, 1992.

Hoffmann:1987:SLS

[Hof87] W. Hoffmann. Solving linear systems on a vector computer. J. Comput. Appl. Math., 18: 353–367, 1987.

Hoffmann:1989:PAR

[Hof89a] Christoph M. Hoffmann. The problems of accuracy and robustness in geometric computation. Computer, March:31– 41, 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoffmann:1989:IAG

[Hof89b] W. Hoffmann. Iterative algorithms for Gram–Schmidt orthogonalization. Computing: Archiv für informatik und numerik, 41:335–348, 1989. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Horning:1979:NPR

[Hor79] Jim Horning. Note on program reliability. ACM Sigsoft, 4(4): 6, 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). Cited in [Tro84a].

Hotelling:1943:SNM

[Hot43] Harold Hotelling. Some new methods in matrix calculation. Annals of mathematical statistics, 14(1):1–34, 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

hous58

[Hou58] Alston S. Householder. Unitary triangularization of a nonsymmetric matrix. J. Assoc. Comput. Mach., 5:339–342, 1958.

Householder:1964:TMN

[Hou64] Alston S. Householder. The Theory of Matrices in Numerical Analysis. Blaisdell, New York, 1964. ISBN 0-486-61781-5. xi + 257 pp. Reprinted by Dover, New York, 1975.

Hough:1981:API

[Hou81] David Hough. Applications of the proposed IEEE 754 standard for floating-point arithmetic. Computer, 14:70–74, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1989:RS

[Hou89] David Hough. Random story. NA Digest, Volume 89, Issue 1, 1989.

Higham:1994:PAC

[HP94a] Nicholas J. Higham and Pythagoras Papadimitriou. A parallel algorithm for computing the polar decomposition. Parallel Comput., 20(8):1161–1173, August 1994.

Higham:1994:SPI

[HP94b] Nicholas J. Higham and Alex Pothen. Stability of the partitioned inverse method for parallel solution of sparse triangular systems. SIAM J. Sci. Comput., 15(1):139–148, January 1994.

Hennessy:1996:CAQ

[HP96] John L. Hennessy and David A. Patterson. Computer Architecture: A Quantitative Approach. Morgan Kaufmann, San Francisco, CA, USA, second edition, 1996. ISBN 1-55860-329-8. xxiii + 760 + appendices pp.

Henderson:1983:HKP

[HPS83] Harold V. Henderson, Friedrich Pukelsheim, and Shayle R. Searle. On the history of the Kronecker product. Linear and Multilinear Algebra, 14:113–120, 1983.

Hu:1992:KSM

[HR92] D. Y. Hu and L. Reichel. Krylov subspace methods for the Sylvester equation. Linear Algebra Appl., 172:283–313, 1992.

Hull:1966:TPM

[HS66] T. E. Hull and J. R. Swenson. Tests of probabilistic models for propagation of roundoff errors. Comm. ACM, 9(2):108– 113, 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henderson:1981:DIS

[HS81a] H. V. Henderson and S. R.

Searle. On deriving the inverse of a sum of matrices. SIAM Rev., 23(1):53–60, 1981.

Henderson:1981:VPM

[HS81b] Harold V. Henderson and S. R.

Searle. The vec-permutation matrix, the vec operator and Kronecker products: A review. Linear and Multilinear Algebra, 9:271–288, 1981.

Higham:1987:NLA

[HS87] Nicholas J. Higham and G. W.

Stewart. Numerical linear algebra in statistical computing. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 41–57. Oxford University Press, 1987.

Hochbruck:1995:PKS

[HS95] Marlis Hochbruck and Gerhard Starke. Preconditioned Krylov subspace methods for Lyapunov matrix equations. SIAM J. Matrix Anal. Appl., 16(1):156–171, 1995.

Hedayat:1999:OAT

[HSS99] A. S. Hedayat, N. J. A. Sloane, and John Stufken. Orthogonal Arrays: Theory and Applications. Springer-Verlag, New York, 1999. ISBN 0-387-98766-5. xxiii + 416 pp.

Higham:1993:SO

[HT93] Desmond J. Higham and Lloyd N. Trefethen. Stiffness of ODEs. BIT, 33:285–303, 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Higham:2000:BAM

[HT00] Nicholas J. Higham and Françoise Tisseur. A block algorithm for matrix 1-norm estimation, with an application to 1-norm pseudospectra. SIAM J. Matrix Anal. Appl., 21(4):1185–1201, 2000.

Higham:2001:BEM

[HT01] Nicholas J. Higham and Françoise Tisseur. Bounds for eigenvalues of matrix polynomials. Numerical Analysis Report 371, Manchester Centre for Computational Mathematics, Manchester, England, January 2001. 16 pp. To appear in Linear Algebra Appl.

Hull:1979:CNS

[Hul79] T. E. Hull. Correctness of numerical software. In Lloyd D. Fosdick, editor, Performance Evaluation of Numerical Software, pages 3–15. North-Hol-land, Amsterdam, The Netherlands, 1979.

Hull:1982:PCE

[Hul82] T. E. Hull. Precision control, exception handling and A choice of numerical algorithms. In G. A. Watson, editor, Numerical Analysis Proceedings, Dundee 1981, volume 912 of Lecture Notes in Mathematics, pages 169–178. Springer-Verlag, Berlin, 1982.

Hunt:1997:ROA

[Hun97] Julian Hunt. Rounding and other approximations for measurements, records and targets. Mathematics Today, 33:73–77, 1997. ISSN 1361-2042.

Hough:1997:COD

[HV97] Patricia D. Hough and Stephen A.

Vavasis. Complete orthogonal decomposition for weighted least squares. SIAM J. Matrix Anal. Appl., 18(2):369–392, 1997.

Hammarling:1976:PBL

[HW76] S. J. Hammarling and J. H.

Wilkinson. The practical behaviour of linear iterative methods with particular reference to S.O.R. Report NAC 69, National Physical Laboratory, Teddington, UK, September 1976. 19 pp.

Hairer:1991:SOD

[HW91] E. Hairer and G. Wanner. Solving Ordinary Differential Equations II. Springer-Verlag, Berlin, 1991. ISBN 3-540-53775-9. xv + 601 pp.

Hyman:1957:EEG

[Hym57] M. A. Hyman. Eigenvalues and eigenvectors of general matrices. Presented at the 12th National Meeting of the Association for Computing Machinery, Houston, Texas, 1957. Cited in [Wil63b].

IBM:1988:ESS

[IBM88] IBM. Engineering and Scientific Subroutine Library, Guide and Reference, Release 3. Fourth Edition (Program Number 5668-863), 1988.

LIA94

[IEC94] ISO/IEC 10967-1 (1994-12): Information Technology—Language Independent Arithmetic—Part 1: Integer and Floating Point Arithmetic. International Electrotechnical Commission, Geneva, Switzerland, 1994. 92 pp.

Group:1981:PSB

[IEE81] IEEE Computer Society Microprocessor Standards Committee, Floating-Point Working Group. A proposed standard for binary floating-point arithmetic, Draft 8.0 of IEEE Task P754 (with introductory comments by David Stevenson). Computer, 14:51–62, 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

ieee85

[IEE85] IEEE Standard for Binary Floating-Point Arithmetic, ANSI/IEEE Standard 754-1985. Institute of Electrical and Electronics Engineers, New York, 1985. Reprinted in SIGPLAN Notices, 22(2):9–25, 1987.

ieee87

[IEE87] A Radix-Independent Standard for Floating-Point Arithmetic, IEEE Standard 854-1987. IEEE Computer Society, New York, 1987.

Isaacson:1966:ANM

[IK66] Eugene Isaacson and Herbert Bishop Keller. Analysis of Numerical Methods. Wiley, New York, 1966. ISBN 0-486-68029-0. xv + 541 pp. Reprinted by Dover, New York, 1994.

Ikramov:2000:BGF

[IK00] Khakim D. Ikramov and Andrey B. Kucherov. Bounding the growth factor in Gaussian elimination for Buckley’s class of complex symmetric matrices. Numerical Linear Algebra with Applications, 7(5):269–274, 2000.

Ikebe:1979:IHM

[Ike79] Yasuhiko Ikebe. On inverses of Hessenberg matrices. Linear Algebra Appl., 24:93–97, 1979.

Ikramov:1998:RNC

[Ikr98] Khakim D. Ikramov. A remark on “A note on constructing a symmetric matrix with specified diagonal entries and eigenvalues”. BIT, 38 (4):807, 1998. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Committee:1993:RGL

[ILA93] ILAS Education Committee. Report on graduate linear algebra courses. Manuscript from the International Linear Algebra Society, November 1993. http:/

                                        /gauss.technion.ac.il/iic/
                                        GRAD-ED.SYLLABI.

Ipsen:1994:USM

[IM94] Ilse C. F. Ipsen and Carl D. Meyer. Uniform stability of Markov chains. SIAM J. Matrix Anal. Appl., 15(4):1061– 1074, 1994.

Incertis:1985:FMC

[Inc85] F. Incertis. A faster method of computing matrix Pythagorean sums. IEEE Trans. Automat. Control, AC-30(3):273– 275, 1985.

Inc:1989:UMS

[Inc89] Cray Research Inc. UNICOS Math and Scientific Library Reference Manual. Number SR-2081, Version 5.0, Eagan, MN, USA, 1989.

Ince:1992:CWM

[Inc92]  D. C. Ince, editor.  Col-

lected Works of A. M. Turing: Mechanical Intelligence. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88058-5. xix + 227 pp.

Iri:1991:HAD

[Iri91] Masao Iri. History of automatic differentiation and rounding error estimation. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, pages 3–16. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991.

Jang:1997:CLS

[Jan97] Ho-Jong Jang. A constrained least-squares approach to the rapid reanalysis of structures. Linear Algebra Appl., 265:185– 202, 1997.

Jennings:1982:BSV

[Jen82] A. Jennings. Bounds for the singular values of a matrix. IMA J. Numer. Anal., 2:459–474, 1982.

Jennings:1974:DEA

[JO74] L. S. Jennings and M. R. Osborne. A direct error analysis for least squares. Numer. Math., 22:325–332, 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jordan:1968:EEG

[Jor68] T. L. Jordan. Experiments on error growth associated with some linear least-squares procedures. Math. Comp., 22:579– 588, 1968.

Joseph:1991:CPN

[Jos91] George Gheverghese Joseph.

The Crest of the Peacock: Non-European Roots of Mathematics. Penguin, London, 1991. ISBN 0-14-012529-9. xv + 371 pp.

Jalby:1991:SAI

[JP91] William Jalby and Bernard Philippe. Stability analysis and improvement of the block Gram–Schmidt algorithm. SIAM J. Sci. Statist. Comput., 12(5):1058–1073, 1991.

Jones:1993:BKF

[JP93] Mark T. Jones and Merrell L. Patrick. Bunch–Kaufman factorization for real symmetric indefinite banded matrices. SIAM J. Matrix Anal. Appl., 14(2): 553–559, 1993.

Jones:1994:FSI

[JP94] Mark T. Jones and Merrell L. Patrick. Factoring symmetric indefinite matrices on high-performance architectures. SIAM J. Matrix Anal. Appl., 15 (1):273–283, 1994.

Jankowski:1983:NFP

[JSW83] M. Jankowski, A. Smoktunowicz, and H. Woźniakowski. A note on floating-point summation of very many terms. Journal of Information Processing and Cybernetics: EIK, 19(9): 435–440, 1983. CODEN JI-CYE5. ISSN 0863-0593.

Jones:1974:NSE

[JT74] William B. Jones and W. J. Thron. Numerical stability in evaluating continued fractions. Math. Comp., 28(127):795–810, 1974.

Jankowski:1977:IRI

[JW77] M. Jankowski and H. Woźniakowski.

Iterative refinement implies numerical stability. BIT, 17: 303–311, 1977. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jankowski:1985:ASC

[JW85] M. Jankowski and H. Woźniakowski.

The accurate solution of certain continuous problems using only single precision arithmetic. BIT, 25:635–651, 1985. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Jansen:1986:HAA

[JW86] Paul Jansen and Peter Weidner. High-accuracy arithmetic software—some tests of the ACRITH problem-solving routines. ACM Trans. Math. Software, 12(1):62–70, 1986. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaagstrom:1994:PAG

[Kåg94] Bo Kågström. A perturbation analysis of the generalized Sylvester equation (AR − LB, DR − LE) = (C, F). SIAM J. Matrix Anal. Appl., 15(4): 1045–1060, 1994.

Kahan:1965:FRR

[Kah65] W. Kahan. Further remarks on reducing truncation errors. Comm. ACM, 8(1):40, 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1966:NLA

[Kah66] W. M. Kahan. Numerical linear algebra. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 9(??): 757–801, ???? 1966. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kahan:1972:SEA

[Kah72] W. Kahan. A survey of error analysis. In Proc. IFIP Congress, Ljubljana, Information Processing 71, pages 1214– 1239. North-Holland, Amsterdam, The Netherlands, 1972.

Kahan:1973:IAL

[Kah73] W. Kahan. Implementation of algorithms (lecture notes by W. S. Haugeland and D. Hough). Technical Report 20, Department of Computer Science, University of California, Berkeley, CA, USA, 1973.

Kahan:1980:IAO

[Kah80] W. Kahan. Interval arithmetic options in the proposed IEEE floating point arithmetic standard. In Karl L. E. Nickel, editor, Interval Mathematics 1980, pages 99–128. Academic Press, New York, 1980.

Kahan:1981:WDW

[Kah81] W. Kahan. Why do we need a floating-point arithmetic standard? Technical report, University of California, Berkeley, CA, USA, February 1981. 41 pp.

Kahan:1983:MWS

[Kah83] W. Kahan. Mathematics written in sand—the HP-15C, Intel 8087, etc. In Statistical Computing Section, Proceedings of the American Statistical Association, Toronto, pages 12–26, 1983.

Kahan:1986:SRC

[Kah86] W. Kahan. To solve a real cubic equation. Technical Report PAM-352, Center for Pure and Applied Mathematics, University of California, Berkeley, CA, USA, November 1986. 20 pp.

Kahan:1987:BCC

[Kah87a] W. Kahan. Branch cuts for complex elementary functions or much ado about nothing’s sign bit. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 165–211. Oxford University Press, 1987.

Kahan:1987:DPI

[Kah87b] W. Kahan. Doubled-precision IEEE standard 754 floatingpoint arithmetic. Manuscript, February 1987.

Kahan:1990:HCA

[Kah90] W. Kahan. How Cray’s arithmetic hurts scientific computation (and what might be done about it). Manuscript prepared for the Cray User Group meeting in Toronto, June 1990.

Kahan:1991:ARL

[Kah91] W. Kahan. Analysis and refutation of the LCAS. ACM SIGNUM Newsletter, 26(3):2– 15, 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1994:CBV

[Kah94] W. Kahan. Computer benchmarks versus accuracy. Draft manuscript, June 1994.

Kahan:1996:IPE

[Kah96a] W. Kahan. The improbability of probabilistic error analyses for numerical computations. Manuscript, March 1996.

Kahan:1996:LNS

[Kah96b] W. Kahan. Lecture notes on the status of IEEE Standard 754 for binary floating-point arithmetic. Manuscript, May 1996.

Kahan:1997:MAA

[Kah97] W. Kahan. Miscalculating area and angles of a needle-like triangle. Manuscript, July 1997.

Kahan:2001:WVT

[Kah01] W. Kahan. What has the volume of a tetrahedron to do with computer programming languages? Manuscript, April 2001.

Kaporin:1999:PAF

[Kap99] Igor Kaporin. A practical algorithm for faster matrix multiplication. Numerical Linear Algebra with Applications, 6:687– 700, 1999.

Karlin:1968:TP

[Kar68] Samuel Karlin. Total Positivity, volume 1. Stanford University Press, Stanford, CA, USA, 1968.

Karasalo:1974:CTD

[Kar74] Ilkka Karasalo. A criterion for truncation of the QRdecomposition algorithm for the singular linear least squares problem. BIT, 14:156–166, 1974. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Karpinski:1985:PFP

[Kar85] Richard Karpinski. Paranoia:

A floating-point benchmark. BYTE Magazine, 10(2):223– 235, February 1985. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Kato:1976:PTL

[Kat76]  Tosio Kato.  Perturbation

Theory for Linear Operators. Springer-Verlag, Berlin, second edition, 1976. ISBN 3-540-97588-5. xxi + 619 pp.

Kaufman:1983:MMQ

[Kau83] Linda Kaufman. Matrix methods for queuing problems. SIAM J. Sci. Statist. Comput., 4(3): 525–552, 1983.

Kuki:1973:SSA

[KC73] H. Kuki and W. J. Cody. A statistical study of the accuracy of floating point number systems. Comm. ACM, 16(4):223– 230, 1973. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Kearfott:1996:AIF

[Kea96a] R. Baker Kearfott. Algorithm 763: INTERVAL ARITHMETIC: A Fortran 90 module for an interval data type. ACM Trans. Math. Software, 22(4):385–392, 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kearfott:1996:ICI

[Kea96b] R. Baker Kearfott. Interval computations: Introduction, uses, and resources. Euromath Bulletin, 2(1):95–112, 1996. http:

//interval.louisiana.edu/
preprints.html.

Keller:1965:SSS

[Kel65] Herbert B. Keller. On the solution of singular and semidefinite linear systems by iteration. SIAM J. Numer. Anal., 2(2): 281–290, 1965.

Kelley:1995:IML

[Kel95] C. T. Kelley. Iterative Methods for Linear and Nonlinear Equations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-352-8. xiii + 165 pp.

Kelley:1999:IMO

[Kel99] C. T. Kelley. Iterative Methods for Optimization. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1999. ISBN 0-89871-433-8. xv + 180 pp.

Kerr:1990:FCT

[Ker90]  Thomas H. Kerr.  Falla-

cies in computational testing of matrix positive definiteness/semidefiniteness. IEEE Transactions on Aerospace and Electronic Systems, 26(2):415– 421, 1990. CODEN IEARAX. ISSN 0018-9251 (print), 1557-9603 (electronic).

Kahan:1963:ANIa

[KF63a] W. Kahan and I. Farkas. Algorithm 168: Newton interpolation with backward divided differences. Comm. ACM, 6(4): 165, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahan:1963:ANIb

[KF63b] W. Kahan and I. Farkas. Algorithm 169: Newton interpolation with forward divided differences. Comm. ACM, 6(4): 165, 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krishnan:2001:PEM

[KFC^{+}01] Shankar Krishnan, Mark Foskey, Tim Culver, John Keyser, and Dinesh Manocha. PRE-CISE: Efficient multiprecision evaluation of algebraic roots and predicates for reliable geometric computation. In David Eppstein, Dan Halperin, and Dinesh Manocha, editors, Proc. ACM Symposium on Computational Geometry, pages 274–283. ACM Press, 2001.

Kennedy:1980:SC

[KG80] William J. Kennedy, Jr. and James E. Gentle. Statistical Computing. Marcel Dekker, New York, 1980. ISBN 0-8247-6898-1. xi + 591 pp.

Kenney:1990:SAD

[KH90] Charles Kenney and Gary Hewer. The sensitivity of the algebraic and differential Riccati equations. SIAM J. Control Optim., 28(1):50–69, 1990.

Kielbasinski:1973:SAC

[Kie73] Andrzej Kiełbasiński. Summation algorithm with corrections and some of its applications. Math. Stos., 1:22–41, 1973. (In Polish, cited in [JSW83] and [JW85]).

Kielbasinski:1981:IRL

[Kie81] Andrzej Kiełbasiński. Iterative refinement for linear systems in variable-precision arithmetic. BIT, 21:97–103, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kielbasinski:1987:NRE

[Kie87] Andrzej Kiełbasiński. A note on rounding-error analysis of Cholesky factorization. Linear Algebra Appl., 88/89:487–494, 1987.

Kittaneh:1995:SVC

[Kit95] Fuad Kittaneh. Singular values of companion matrices and bounds on zeros of polynomials.

SIAM J. Matrix Anal. Appl., 16 (1):333–340, 1995.

Klatte:1993:CXC

[KKL^{+}93] R. Klatte, U. W. Kulisch, C. Lawo, M. Rauch, and A. Wiethoff. C-XSC: A C++ Class Library for Extended Scientific Computing. Springer-Verlag, Berlin, 1993. ISBN 0-387-56328-8.

Klatte:1992:PXL

[KKN^{+}92] R. Klatte, U. W. Kulisch, M. Neaga, D. Ratz, and Ch. Ullrich. PASCAL-XSC— Language Reference with Examples. Springer-Verlag, Berlin, 1992.

Kahan:1985:AIA

[KL85] W. Kahan and E. LeBlanc.

Anomalies in the IBM ACRITH package. In Kai Hwang, editor, Proceedings of the 7th Symposium on Computer Arithmetic, pages 322–331. IEEE Computer Society Press, Silver Spring, MD, USA, 1985.

Kenney:1988:CSR

[KL88] Charles Kenney and Alan J.

Laub. Controllability and stability radii for companion form systems. Mathematics of control, signals, and systems: MCSS, 1: 239–256, 1988. CODEN MC-SYE8. ISSN 0932-4194 (print), 1435-568x (electronic).

Kenney:1994:SSS

[KL94] Charles S. Kenney and Alan J.

Laub. Small-sample statistical condition estimates for general matrix functions. SIAM J. Sci. Comput., 15(1):36–61, 1994.

Kenney:1992:MSA

[KLP92] Charles S. Kenney, Alan J. Laub, and Philip M. Papadopoulos. Matrix-sign algorithms for Riccati equations. IMA J. Math. Control Inform., 9:331–344, 1992.

Kenney:1998:SCEb

[KLR98a] Charles Kenney, Alan J. Laub, and M. S. Reese. Statistical condition estimation for linear least squares. SIAM J. Matrix Anal. Appl., 19(4):906–923, 1998.

Kenney:1998:SCEa

[KLR98b] Charles Kenney, Alan J. Laub, and M. S. Reese. Statistical condition estimation for linear systems. SIAM J. Sci. Comput., 19 (2):566–583, 1998.

Kulisch:1981:CAT

[KM81] Ulrich W. Kulisch and Willard L. Miranker. Computer Arithmetic in Theory and in Practice. Academic Press, New York, 1981. ISBN 0-12-428650-X. xiii + 249 pp.

Kulisch:1983:NAS

[KM83] Ulrich W. Kulisch and Willard L. Miranker, editors. A New Approach to Scientific Computation. Academic Press, New York, 1983. ISBN 0-12-428660-7.

Kulisch:1986:ADC

[KM86] Ulrich W. Kulisch and Willard L.

Miranker. The arithmetic of the digital computer: A new approach. SIAM Rev., 28(1):1–40, 1986.

Karp:1997:HPD

[KM97] Alan H. Karp and Peter Markstein. High-precision division and square root. ACM Trans. Math. Software, 23(4):561–589, 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kahaner:1989:NMS

[KMN89] David K. Kahaner, Cleve B.

Moler, and Stephen G. Nash. Numerical Methods and Software. Prentice-Hall, Englewood Cliffs, NJ, USA, 1989. ISBN 0-13-627258-4. xii + 495 pp.

Knight:1993:EAS

[Kni93] Philip A. Knight. Error Analysis of Stationary Iteration and Associated Problems. PhD thesis, University of Manchester, Manchester, England, September 1993. 135 pp.

Knight:1995:FRM

[Kni95] Philip A. Knight. Fast rectangular matrix multiplication and QR decomposition. Linear Algebra Appl., 221:69–81, 1995.

Knuth:1962:EPC

[Knu62] Donald E. Knuth. Evaluation of polynomials by computer. Comm. ACM, 5(12):

595–599, December 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [?].

Knuth:1992:TNN

[Knu92] Donald E. Knuth. Two notes on notation. Amer. Math. Monthly, 99(5):403–422, 1992.

Knuth:1997:FA

[Knu97] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-89683-4. xix + 650 pp. LCCN QA76.6 .K64 1997. US$49.50.

Knuth:1998:SA

[Knu98] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Knuth:1999:DT

[Knu99] Donald E. Knuth. Digital Typography. CSLI Lecture Notes Number 78. Center for the Study of Language and Information, Stanford University, Stanford, CA, USA, 1999. ISBN 0-575-86010-4. xv + 685 pp.

Karatsuba:1963:MMN

[KO63] A. Karatsuba and Yu. Ofman. Multiplication of multidigit numbers on automata.

Soviet Physics—Doklady, 7(7): 595–596, January 1963. CO-DEN SPHDA9. ISSN 0038-5689.

Korner:1988:FA

[Kör88] T. W. Körner. Fourier Analysis. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-38991-7. xii + 591 pp.

Kostlan:1992:SGM

[Kos92] Eric Kostlan. On the spectra of Gaussian matrices. Linear Algebra Appl., 162–164:385–388, 1992.

Kovarik:1976:CAS

[Kov76] Z. V. Kovarik. Compatibility of approximate solutions of inaccurate linear equations. Linear Algebra Appl., 15:217–225, 1976.

Kahan:1979:PFP

[KP79] W. Kahan and J. Palmer. On a proposed floating-point standard. ACM SIGNUM Newsletter, 14:13–21, October 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kaagstrom:1992:DSM

[KP92] Bo Kågström and Peter Poromaa. Distributed and shared memory block algorithms for the triangular Sylvester equation with sep^{−1} estimators. SIAM J. Matrix Anal. Appl., 13(1):90– 101, 1992.

Kaagstrom:1996:CES

[KP96a] Bo Kågström and Peter Poromaa. Computing eigenspaces with specified eigenvalues of a regular matrix pair (A, B) and condition estimation: Theory, algorithms and software. Numerical Algorithms, 12:369–407, 1996.

Kaagstrom:1996:LSA

[KP96b] Bo Kågström and Peter Poromaa. LAPACK-style algorithms and software for solving the generalized Sylvester equation and estimating the separation between regular matrix pairs. ACM Trans. Math. Software, 22(1):78–103, 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kreczmar:1976:MRS

[Kre76] Antoni Kreczmar. On memory requirements of Strassen’s algorithms. volume 45 of Lecture Notes in Computer Science, pages 404–407. Springer-Verlag, Berlin, 1976.

Kielbasinski:1988:NLA

[KS88] Andrzej Kiełbasiński and Hubert Schwetlick. Numerische Lineare Algebra: Eine Computerorientierte Einführung. VEB Deutscher, Berlin, 1988. ISBN 3-87144-999-7. 472 pp.

Kielbasinski:1992:NAL

[KS92] Andrzej Kiełbasiński and Hubert Schwetlick. Numeryczna Algebra Liniowa: Wprowadzenie do Obliczeń Zautomatyzowanych. Wydawnictwa Naukowo-Techniczne, Warszawa, 1992. ISBN 83-204-1260-9. 502 pp.

Kosowski:2000:CUT

[KS00] Przemysław Kosowski and Alicja Smoktunowicz. On constructing unit triangular matrices with prescribed singular values. Computing: Archiv für informatik und numerik, 64:279– 285, 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Kubota:1991:PFP

[Kub91] Koichi Kubota. PADRE2, a Fortran precompiler yielding error estimates and second derivatives. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, pages 251–262. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991.

Kuperman:1971:ALA

[Kup71] I. B. Kuperman. Approximate Linear Algebraic Equations. Van Nostrand Reinhold, London, 1971. xi + 225 pp.

Kaagstrom:1989:GSM

[KW89] Bo Kågström and Lars Westin.

Generalized Schur methods with condition estimators for solving the generalized Sylvester equation. IEEE Trans. Automat. Control, AC-34(7):745– 751, 1989.

Kuczynski:1992:ELE

[KW92] J. Kuczyński and H. Woźniakowski.

Estimating the largest eigenvalue by the power and Lanczos algorithms with a random start. SIAM J. Matrix Anal. Appl., 13 (4):1094–1122, 1992.

Karlson:1997:EOB

[KW97] Rune Karlson and Bertil Waldén. Estimation of optimal backward perturbation bounds for the linear least squares problem. BIT, 37(4):862–869, 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Kao:2000:LTA

[KW00] Ming-Yang Kao and Jie Wang. Linear-time approximation algorithms for computing numerical summation with provably small errors. SIAM J. Comput., 29(5): 1568–1576, 2000.

Laderman:1976:NAM

[Lad76] J. D. Laderman. A noncommutative algorithm for multiplying 3 × 3 matrices using 23 multiplications. Bull. Amer. Math. Soc., 82(1):126–128, 1976.

Lanczos:1956:AA

[Lan56]  Cornelius Lanczos.  Applied

Analysis. Prentice-Hall, Englewood Cliffs, NJ, USA, 1956. ISBN 0-486-65656-X. xx + 539 pp. Reprinted by Dover, New York, 1988.

Lancaster:1966:EAN

[Lan66] Peter Lancaster. Error analysis for the Newton–Raphson method. Numer. Math., 9:55– 68, 1966. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Lancaster:1970:ESL

[Lan70] Peter Lancaster. Explicit solutions of linear matrix equations. SIAM Rev., 12(4):544– 566, 1970.

Langlois:2001:ALC

[Lan01] Philippe Langlois. Automatic linear correction of rounding errors. BIT, 41(3):515–539, 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Laszlo:1994:ALB

[Lás94] Lajos László. An attainable lower bound for the best normal approximation. SIAM J. Matrix Anal. Appl., 15(3):1035– 1043, 1994.

Lauchli:1961:JEA

[Läu61] Peter Läuchli. Jordan-Elimination und Ausgleichung nach kleinsten Quadraten. Numer. Math., 3:226–240, 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Laub:1979:SMS

[Lau79] Alan J. Laub. A Schur method for solving algebraic Riccati equations. IEEE Trans. Automat. Control, AC-24(6):913– 921, 1979.

Lavington:1980:EBC

[Lav80] Simon Lavington. Early British Computers: The Story of Vintage Computers and the People Who Built Them. Manchester University Press, 1980. ISBN 0-7190-0803-4. 139 pp.

Li:1998:MSG

[LD98] Xiaoye S. Li and James W. Demmel. Making sparse Gaussian elimination scalable by static pivoting. In Proceedings of SC98: High Performance Networking and Computing Conference, Orlando, Florida, 1998. CD ROM.

Li:2000:DIT

[LDB^{+}00] X. S. Li, J. W. Demmel, D. H. Bailey, G. Henry, Y. Hida, J. Iskandar, W. Kahan, A. Kapur, M. C. Martin, T. Tung, and D. J. Yoo. Design, implementation and testing of Extended and Mixed Precision BLAS. Technical Report CS-00-451, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 2000. 61 pp. LAPACK Working Note 149.

Lehmer:1943:TMP

[Leh43] D. H. Lehmer. Tables to many places of decimals. Mathematical Tables and Other Aids to Computation, 1(1):30–31, 1943. CODEN MTTCAS. ISSN 0891-6837.

Lehoucq:1996:CEU

[Leh96] R. B. Lehoucq. The computation of elementary unitary matrices. ACM Trans. Math. Software, 22(4):393–400, 1996. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lemeire:1975:BCN

[Lem75] Frans Lemeire. Bounds for condition numbers of triangular and trapezoid matrices. BIT, 15: 58–64, 1975. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Lawson:1995:SLS

[LH95] Charles L. Lawson and Richard J.

Hanson. Solving Least Squares Problems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-356-0. xii + 337 pp. Revised republication of work first published in 1974 by Prentice-Hall.

Lawson:1979:BLA

[LHKK79] C. L. Lawson, R. J. Hanson, D. R. Kincaid, and F. T. Krogh. Basic linear algebra subprograms for Fortran usage. ACM Trans. Math. Software, 5(3): 308–323, 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lickteig:1987:CCD

[Lic87] Thomas Lickteig. The computational complexity of division in quadratic extension fields. SIAM J. Comput., 16(2):278– 311, 1987.

Linz:1970:AFP

[Lin70] Peter Linz. Accurate floatingpoint summation. Comm. ACM, 13(6):361–362, 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linnainmaa:1974:ASK

[Lin74]  Seppo Linnainmaa.  Analy-

sis of some known methods of improving the accuracy of floating-point sums. BIT, 14: 167–202, 1974. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1975:TAS

[Lin75] Seppo Linnainmaa. Towards accurate statistical estimation of rounding errors in floatingpoint computations. BIT, 15: 165–173, 1975. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1976:TEA

[Lin76]  Seppo Linnainmaa.  Taylor

expansion of the accumulated rounding error. BIT, 16:146– 160, 1976. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Linnainmaa:1981:SDP

[Lin81] Seppo Linnainmaa. Software for doubled-precision floatingpoint computations. ACM Trans. Math. Software, 7(3): 272–283, 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Linzer:1992:STB

[Lin92] Elliot Linzer. On the stability of transform-based circular deconvolution. SIAM J. Numer. Anal., 29(5):1482–1492, 1992.

Liu:1987:PPS

[Liu87] Joseph W. H. Liu. A partial pivoting strategy for sparse symmetric matrix decomposition. ACM Trans. Math. Software, 13(2):173–182, 1987. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lyness:1976:CNA

[LK76] J. N. Lyness and J. J. Kaganove.

Comments on the nature of automatic quadrature routines. ACM Trans. Math. Software, 2 (1):65–81, 1976. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lyness:1966:VMS

[LM66] J. N. Lyness and C. B. Moler.

Van der Monde systems and numerical differentiation. Numer. Math., 8:458–464, 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lefevre:2001:WCC

[LM01] Vincent Lefèvre and Jean-Michel Muller. Worst cases for correct rounding of the elementary functions in double precision. In Proc. 15th IEEE Symposium on Computer Arithmetic. IEEE Computer Society Press, Los Alamitos, CA, USA, 2001.

Lefevre:1998:TCR

[LMT98] Vincent Lefèvre, Jean-Michel Muller, and Arnaud Tisserand.

Toward correctly rounded transcendentals. IEEE Trans. Comput., 47(11):1235–1243, 1998.

Leuprecht:1982:PAR

[LO82] H. Leuprecht and W. Oberaigner. Parallel algorithms for the rounding exact summation of floating point numbers. Computing: Archiv für informatik und numerik, 28:89–104, 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Loizou:1969:NJC

[Loi69] Georghios Loizou. Nonnormality and Jordan condition numbers of matrices. J. Assoc. Comput. Mach., 16(4):580–584, 1969.

Longley:1967:ALS

[Lon67] J. W. Longley. An appraisal of least squares programs for the electronic computer from the point of view of the user. J. Amer. Statist. Assoc., 62:819– 841, 1967.

Laderman:1992:PAA

[LPS92] Julian Laderman, Victor Pan, and Xuan-He Sha. On practical algorithms for accelerated matrix multiplication. Linear Algebra Appl., 162–164:557–588, 1992.

Larson:1983:ASR

[LPW83] John L. Larson, Mary E. Pasternak, and John A. Wisniewski. Algorithm 594: Software for relative error analysis. ACM Trans. Math. Software, 9(1): 125–130, 1983. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lagarias:1998:CPN

[LRWW98] Jeffrey C. Lagarias, James A.

Reeds, Margaret H. Wright, and Paul E. Wright. Convergence properties of the Nelder–Mead simplex method in low dimensions. SIAM J. Optim., 9(1): 112–147, 1998.

Larson:1978:ECE

[LS78] John L. Larson and Ahmed H.

Sameh. Efficient calculation of the effects of roundoff errors. ACM Trans. Math. Software, 4(3):228–236, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Errata 5(3):372, 1979.

Larson:1980:ARE

[LS80] John L. Larson and Ahmed H.

Sameh. Algorithms for roundoff error analysis—A relative error approach. Computing: Archiv für informatik und numerik, 24:275–297, 1980. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Lewis:2000:DSM

[LTT00] Robert Michael Lewis, Virginia J. Torczon, and Michael W. Trossett. Direct search methods: Then and now. J. Comput. Appl. Math., 124:191–207, 2000.

Lu:1994:FSC

[Lu94] Hao Lu. Fast solution of confluent Vandermonde linear systems. SIAM J. Matrix Anal. Appl., 15(4):1277–1289, 1994.

Lu:1995:FAC

[Lu95] Hao Lu. Fast algorithms for confluent Vandermonde linear systems and generalized Trummer’s problem. SIAM J. Matrix Anal. Appl., 16(2):655–674, 1995.

Lu:1996:SVL

[Lu96] Hao Lu. Solution of Vandermondelike systems and confluent Vandermonde-like systems. SIAM J. Matrix Anal. Appl., 17(1): 127–138, 1996.

Lay-Yong:1989:MSL

[LYK89] Lam Lay-Yong and Shen Kangshen. Methods of solving linear equations in traditional China. Historia Mathematica, 16(2):107–122, May 1989. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0315086089900335.

Lynn:1964:REM

[Lyn64] M. Stuart Lynn. On the roundoff error in the method of successive over-relaxation. Math. Comp., 18(85):36–49, 1964.

Li:1992:HDA

[LZ92] T. Y. Li and Zhonggang Zeng. Homotopy-determinant algorithm for solving nonsymmetric eigenvalue problems. Math. Comp., 59(200):483–502, 1992.

Macleod:1989:SSG

[Mac89]  Allan J. Macleod.  Some

statistics on Gaussian elimination with partial pivoting. ACM SIGNUM Newsletter, 24 (2/3):10–14, 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Maindonald:1984:SC

[Mai84] J. H. Maindonald. Statistical Computation. Wiley, New York, 1984. ISBN 0-471-86452-8. xviii + 370 pp.

Makhoul:1991:TDP

[Mak91] John Makhoul. Toeplitz determinants and positive semidefiniteness. IEEE Trans. Signal Processing, 39(3):743–746, 1991.

Malcolm:1971:AFP

[Mal71] Michael A. Malcolm. On accurate floating-point summation. Comm. ACM, 14(11): 731–736, November 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malcolm:1972:ARP

[Mal72]  Michael A. Malcolm.  Al-

gorithms to reveal properties of floating-point arithmetic. Comm. ACM, 15(11):949–951, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malyshev:2000:NSG

[Mal00] Alexander N. Malyshev. A note on the stability of Gauss–Jordan elimination for diagonally dominant matrices. Computing: Archiv für informatik und numerik, 65(3):281–284, 2000. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Malyshev:2001:OBP

[Mal01] Alexander N. Malyshev. Optimal backward perturbation bounds for the LSS problem. BIT, 41(2):430–432, 2001. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Manteuffel:1981:IAA

[Man81] Thomas A. Manteuffel. An interval analysis approach to rank determination in linear least squares problems. SIAM J. Sci. Statist. Comput., 2(3):335–348, 1981.

maple

[Map] Maple. Waterloo Maple Inc., Waterloo, Ontario, Canada.

http://www.maplesoft.com.

Markoff:1994:CFC

[Mar94] John Markoff. Circuit flaw causes Pentium chip to miscalculate, Intel admits. New York Times, 1994. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. 24 November.

mathematica

[Mata] Mathematica. Wolfram Research, Inc., Champaign, IL,

USA. http://www.wolfram.
com.

matlab

[Matb] MATLAB. The MathWorks, Inc., Natick, MA, USA. http:

//www.mathworks.com.

matlab:SMT

[Matc] Symbolic Math Toolbox Version 2: User’s Guide. The Math-Works, Inc., Natick, MA, USA. Online version.

gaze24

[Mat24] Gleanings far and near. Mathematical Gazette, 22(170):95, 1924. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Matula:1968:C

[Mat68] David W. Matula. In-and-out conversions. Comm. ACM, 11 (1):47–50, January 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Gol67, ?].

Matula:1970:FFP

[Mat70] David W. Matula. A formalization of floating-point numeric base conversion. IEEE Trans. Comput., C-19(8):681– 692, 1970.

Mattheij:1984:SBD

[Mat84a] R. M. M. Mattheij. Stability of block LU-decompositions of matrices arising from BVP. SIAM J. Alg. Discrete Methods, 5(3): 314–331, 1984.

Mattheij:1984:SLD

[Mat84b] R. M. M. Mattheij. The stability of LU-decompositions of block tridiagonal matrices. Bulletin of the Australian Mathematical Society, 29:177–205, 1984. CO-DEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic).

Mathias:1992:MPD

[Mat92] Roy Mathias. Matrices with positive definite Hermitian part: Inequalities and linear systems. SIAM J. Matrix Anal. Appl., 13 (2):640–654, 1992.

Mathias:1995:AEC

[Mat95a] Roy Mathias. Accurate eigensystem computations by Jacobi methods. SIAM J. Matrix Anal. Appl., 16(3):977–1003, 1995.

Mathias:1995:IPP

[Mat95b] Roy Mathias. The instability of parallel prefix matrix multiplication. SIAM J. Sci. Comput., 16(4):956–973, 1995.

Mathias:1996:AAO

[Mat96] Roy Mathias. Analysis of algorithms for orthogonalizing products of unitary matrices. Numerical Linear Algebra with Applications, 3(2):125–145, 1996.

Merikoski:1997:BUB

[MAV^{+}97] Jorma Kaarlo Merikoski, Uoti Arpala, Ari Virtanen, Tin-Yau Tam, and Frank Uhlig. A best upper bound for the 2-norm condition number of a matrix. Linear Algebra Appl., 254:355–365, 1997.

McKeeman:1962:ACE

[McK62] William Marshall McKeeman.

Algorithm 135: Crout with equilibration and iteration. Comm. ACM, 5:553–555, 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKinnon:1998:CNM

[McK98] K. I. M. McKinnon. Convergence of the Nelder–Mead simplex method to a nonstationary point. SIAM J. Optim., 9(1): 148–158, 1998.

McCracken:1964:NMF

[MD64] Daniel D. McCracken and William S. Dorn. Numerical Methods and Fortran Programming: With Applications in Science and Engineering. Wiley, New York, 1964. xii + 457 pp.

Meinguet:1969:ES

[Mei69] Jean Meinguet. On the estimation of significance. In Elden Hansen, editor, Topics in Interval Analysis, pages 47–64. Oxford University Press, 1969.

Meinguet:1983:REA

[Mei83] Jean Meinguet. Refined error analyses of Cholesky factorization. SIAM J. Numer. Anal., 20 (6):1243–1250, 1983.

Mendelsohn:1956:SEP

[Men56] N. S. Mendelsohn. Some elementary properties of ill conditioned matrices and linear equations. Amer. Math. Monthly, 63 (5):285–295, 1956.

Metropolis:1977:MSA

[Met77] N. Metropolis. Methods of significance arithmetic. In David A. H. Jacobs, editor, The State of the Art in Numerical Analysis, pages 179–192. Academic Press, London, 1977.

Meurant:1992:RIS

[Meu92] Gérard Meurant. A review on the inverse of symmetric tridiagonal and block tridiagonal matrices. SIAM J. Matrix Anal. Appl., 13(3):707–728, 1992.

Meyer:1957:MIL

[MH57] H. I. Meyer and B. J. Hollingsworth. A method of inverting large matrices of special form. M.T.A.C., 11:94–97, 1957.

Matsui:1981:OUF

[MI81] Shouichi Matsui and Masao Iri. An overflow/underflow-free floating-point representation of numbers. Journal of Information Processing, 4(3):123–133, 1981. CODEN JIPRDE. ISSN 0387-6101.

Miller:1975:CCN

[Mil75a] Webb Miller. Computational complexity and numerical stability. SIAM J. Comput., 4(2): 97–107, 1975.

Miller:1975:SRA

[Mil75b] Webb Miller. Software for roundoff analysis. ACM Trans. Math. Software, 1(2):108–128, 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Miller:1976:GTR

[Mil76] Webb Miller. Graph transformations for roundoff analysis. SIAM J. Comput., 5(2):204– 216, 1976.

Miller:1984:ENS

[Mil84] Webb Miller. The Engineering of Numerical Software. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984. ISBN 0-13-279043-2. viii + 167 pp.

Milenkovic:1988:VIG

[Mil88a] Victor J. Milenkovic. Verifiable implementations of geometric algorithms using finite precision arithmetic. Artificial Intelligence, 37:377–401, 1988. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Miller:1988:ISM

[Mil88b] D. F. Miller. The iterative solution of the matrix equation XA + BX + C = 0. Linear Algebra Appl., 105:131–137, 1988.

Mirsky:1961:ILA

[Mir61] L. Mirsky. An Introduction to Linear Algebra. Oxford University Press, 1961. ISBN 0-486-66434-1. viii + 440 pp. Reprinted by Dover, New York, 1990.

Mitchell:1948:IMO

[Mit48] Herbert F. Mitchell, Jr. Inversion of a matrix of order 38. M.T.A.C., 3:161–166, 1948.

Matula:1985:FPR

[MK85] David W. Matula and Peter Kornerup. Finite precision rational arithmetic: Slash number systems. IEEE Trans. Comput., C-34(1):3–18, 1985.

Moore:1998:MCP

[MLK98] J. Strother Moore, Thomas W. Lynch, and Matt Kaufmann. A mechanically checked proof of the AMD5_{K}86 floating-point division program. IEEE Trans. Comput., 47(9):913–926, 1998.

Moler:1983:RSR

[MM83a] Cleve B. Moler and Donald Morrison. Replacing square roots by Pythagorean sums. IBM Journal of Research and Development, 27(6):577–581, 1983. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Moler:1983:SVA

[MM83b] Cleve B. Moler and Donald Morrison. Singular value analysis of cryptograms. Amer. Math. Monthly, 90:78–87, 1983.

Magnus:1979:CMS

[MN79] Jan R. Magnus and Heinz Neudecker. The commutation matrix: Some properties and applications. Ann. Statist., 7 (2):381–394, 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

McCullagh:1989:GLM

[MN89] P. McCullagh and J. A. Nelder.

Generalized Linear Models. Chapman and Hall, London, second edition, 1989. ISBN 0-412-31760-5. xix + 511 pp.

Magnus:1999:MDC

[MN99] Jan R. Magnus and Heinz Neudecker. Matrix Differential Calculus with Applications in Statistics and Econometrics. Wiley, Chichester, UK, revised edition, 1999. ISBN 0-471-98633-X. xviii + 395 pp.

Moller:1965:NQD

[Møl65a] Ole Møller. Note on quasi double-precision. BIT, 5:251– 255, 1965. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Moller:1965:QDP

[Møl65b] Ole Møller. Quasi doubleprecision in floating point addition. BIT, 5:37–50, 1965. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). See also [Møl65a] for remarks on this article.

Moler:1967:IRF

[Mol67] Cleve B. Moler. Iterative refinement in floating point. J. Assoc. Comput. Mach., 14(2):316–321, 1967.

Moler:1974:CRS

[Mol74]  Cleve B. Moler.  Cramer’s

rule on 2-by-2 systems. ACM SIGNUM Newsletter, 9(4):13– 14, 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1978:TRP

[Mol78] Cleve B. Moler. Three research problems in numerical linear algebra. In G. H. Golub and J. Oliger, editors, Numerical Analysis, volume 22 of Proceedings of Symposia in Applied Mathematics, pages 1–18. American Mathematical Society, Providence, RI, USA, 1978.

Moler:1990:CCW

[Mol90a] Cleve B. Moler. Cleve’s corner: The world’s simplest impossible problem. The MathWorks Newsletter, 4(2):6–7, 1990.

Moler:1990:TND

[Mol90b] Cleve B. Moler. Technical note: Double-rounding and implications for numeric computations. The MathWorks Newsletter, 4:6, 1990.

Moler:1995:TTN

[Mol95] Cleve B. Moler. A tale of two numbers. SIAM News, 28: 1,16, January 1995. ISSN 0036-1437. Also in MATLAB News and Notes, Winter 1995, 10–12.

Moler:1972:MCF

[Mol2a] Cleve B. Moler. Matrix computations with Fortran and paging. Comm. ACM, 15(4):268– 270, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moler:1972:ALE

[Mol2b] Cleve B. Moler. Algorithm 423:

Linear equation solver. Comm. ACM, 15(4):274, 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moore:1966:IA

[Moo66] Ramon E. Moore. Interval Analysis. Prentice-Hall, Englewood Cliffs, NJ, USA, 1966. xi + 145 pp.

Moore:1979:MAI

[Moo79] Ramon E. Moore. Methods and Applications of Interval Analysis. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-161-4. xi + 190 pp.

Malcolm:1974:FMS

[MP74] Michael A. Malcolm and John Palmer. A fast method for solving a class of tridiagonal linear systems. Comm. ACM, 17(1):14–17, 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:1977:CPM

[MP77] Carl D. Meyer, Jr. and R. J.

Plemmons. Convergent powers of a matrix with applications to iterative methods for singular linear systems. SIAM J. Numer. Anal., 14(4):699–705, 1977.

Martin:1966:IRS

[MPW66] R. S. Martin, G. Peters, and J. H. Wilkinson. Iterative refinement of the solution of a positive definite system of equations. Numer. Math., 8:203–216, 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 31– 44], Contribution I/2.

Metcalf:1999:FE

[MR99] Michael Metcalf and John K. Reid. Fortran 90/95 Explained. Oxford University Press, second edition, 1999. ISBN 0-19-850558-2. 358 pp.

Mead:2001:SPS

[MRW01] J. L. Mead, R. E. Renaut, and B. D. Welfert. Stability of a pivoting strategy for parallel Gaussian elimination. BIT, 41(3): 633–639, 2001. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

McCarthy:1974:OCM

[MS74] Charles McCarthy and Gilbert Strang. Optimal conditioning of matrices. SIAM J. Numer. Anal., 10(2):370–388, 1974.

Miller:19xx:SRA

[MS78a] Webb Miller and David Spooner. Software for roundoff analysis, II. ACM Trans. Math. Software, 4(4):369–387, 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Miller:19xx:ASR

[MS78b] Webb Miller and David Spooner. Algorithm 532: Software for roundoff analysis. ACM Trans.

Math. Software, 4(4):388–390, 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Malyshev:2002:COB

[MS02] Alexander N. Malyshev and Miloud Sadkane. Computation of optimal backward perturbation bounds for large sparse linear least squares problems. BIT, 41(4):739–747, 2002. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Muirhead:1982:AMS

[Mui82] Robb J. Muirhead. Aspects of Multivariate Statistical Theory. Wiley, New York, 1982. ISBN 0-471-09442-0. xix + 673 pp.

Muller:1983:REA

[Mül83] K. H. Müller. Rounding error analysis of Horner’s scheme. Computing: Archiv für informatik und numerik, 30:285–303, 1983. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Muller:1989:AO

[Mul89] Jean-Michel Muller. Arithmétique des Ordinateurs. Masson, Paris, 1989. ISBN 2-225-81689-1. 214 pp. In French. Cited in [FM91].

Muller:1997:EFA

[Mul97] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser, Boston, MA, USA, 1997. ISBN 0-8176-3990-X. xv + 204 pp.

Moler:1978:NDW

[MV78] Cleve B. Moler and Charles F. Van Loan. Nineteen dubious ways to compute the exponential of a matrix. SIAM Rev., 20 (4):801–836, 1978.

McCullough:1999:NRE

[MV99] B. D. McCullough and H. D. Vinod. The numerical reliability of econometric software. Journal of economic literature, 37:633– 665, 1999. CODEN JECLB3. ISSN 0022-0515 (print), 1547-1101 (electronic).

Miller:1980:SRA

[MW80] Webb Miller and Celia Wrathall. Software for Roundoff Analysis of Matrix Algorithms. Academic Press, New York, 1980. ISBN 0-12-497250-0. x + 151 pp.

fpv86

[NAG86] FPV: A floating-point validation package. Release I. User’s guide. Technical Report NP 1201, NAG Ltd., Oxford, May 1986.

Nelson:1998:PDM

[Nel98] David Nelson, editor. The Penguin Dictionary of Mathematics. Penguin, London, second edition, 1998. ISBN 0-14-051342-6. 461 pp.

Neumaier:1974:REV

[Neu74] A. Neumaier. Rundungsfehleranalyse einiger Verfahren zur Summation endlicher Summen. Z. Angew. Math. Mech., 54:39– 51, 1974.

Neumaier:1985:IPR

[Neu85] A. Neumaier. Inner product rounding error analysis in the presence of underflow. Computing: Archiv für informatik und numerik, 34:365–373, 1985. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Neumaier:1986:CMM

[Neu86] A. Neumaier. On the comparison of H-matrices with Mmatrices. Linear Algebra Appl., 83:135–141, 1986.

Newcomb:1881:NFU

[New81] Simon Newcomb. Note on the frequency of use of the different digits in natural numbers. American Journal of Mathematics, 4:39–40, 1881. CODEN AJ-MAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Newman:1962:MC

[New62] Morris Newman. Matrix computations. In John Todd, editor, Survey of Numerical Analysis, pages 222–254. McGraw-Hill, New York, 1962.

Newbery:1973:EAF

[New73] A. C. R. Newbery. Error analysis for Fourier series evaluation. Math. Comp., 27(123):639–644, 1973.

Newbery:1974:EAP

[New74] A. C. R. Newbery. Error analysis for polynomial evaluation. Math. Comp., 28(127):789–793, 1974.

Nickel:1970:KBS

[Nic70]  K. Nickel.  Das Kahan-

Babus̆kasche Summierungsverfahren in Triplex-ALGOL 60. Z. Angew. Math. Mech., 50:369– 373, 1970.

Nickel:1977:IA

[Nic77] Karl Nickel. Interval-analysis. In David A. H. Jacobs, editor, The State of the Art in Numerical Analysis, pages 193–225. Academic Press, London, 1977.

Nelder:1965:SMF

[NM65] J. A. Nelder and R. Mead. A simplex method for function minimization. Comput. J., 7: 308–313, 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Notay:1993:CRC

[Not93] Yvan Notay. On the convergence rate of the conjugate gradients in presence of rounding errors. Numer. Math., 65:301– 317, 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neumann:1984:BEA

[NP84] M. Neumann and R. J. Plemmons. Backward error analysis for linear systems associated with inverses of H-matrices. BIT, 24:102–112, 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Neal:1992:GAG

[NP92] Larry Neal and George Poole. A geometric analysis of Gaussian elimination. II. Linear Algebra Appl., 173:239–264, 1992.

Nashed:1976:ABG

[NR76] M. Zuhair Nashed and L. B.

Rall. Annotated bibliography on generalized inverses and applications. In M. Zuhair Nashed, editor, Generalized Inverses and Applications, pages 771–1041. Academic Press, New York, 1976.

Newman:1958:EMI

[NT58] Morris Newman and John Todd.

The evaluation of matrix inversion programs. J. Soc. Indust. Appl. Math., 6(4):466–476, 1958.

Nocedal:1999:NO

[NW99] Jorge Nocedal and Stephen J.

Wright. Numerical Optimization. Springer-Verlag, New York, 1999. ISBN 0-387-98793-2. xx + 636 pp.

OCinneide:1993:EPT

[O’C93] Colm Art O’Cinneide. Entrywise perturbation theory and error analysis for Markov chains. Numer. Math., 65:109– 120, 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oettli:1965:SSL

[Oet65] W. Oettli. On the solution set of a linear system with inaccurate coefficients. SIAM J. Numer. Anal., 2(1):115–118, 1965.

OLeary:1980:EMC

[O’L80] Dianne Prost O’Leary. Estimating matrix condition numbers. SIAM J. Sci. Statist. Comput., 1(2):205–209, 1980.

Oliver:1977:EAM

[Oli77] J. Oliver. An error analysis of the modified Clenshaw method for evaluating Chebyshev and Fourier series. J. Inst. Maths. Applics., 20:379–391, 1977.

Oliver:1979:REP

[Oli79] J. Oliver. Rounding error propagation in polynomial evaluation schemes. J. Comput. Appl. Math., 5(2):85–97, 1979.

Olver:1978:NAE

[Olv78] F. W. J. Olver. A new approach to error arithmetic. SIAM J. Numer. Anal., 15(2):368–393, 1978.

Olver:1983:EAC

[Olv83] F. W. J. Olver. Error analysis of complex arithmetic. In Computational Aspects of Complex Analysis, volume 102 of NATO Advanced Study Institute Series C, pages 279–292. D. Reidel, Dordrecht, Holland, 1983.

Olver:1986:EBP

[Olv86] F. W. J. Olver. Error bounds for polynomial evaluation and complex arithmetic. IMA J. Numer. Anal., 6:373–379, 1986.

Oettli:1964:CAS

[OP64] W. Oettli and W. Prager. Compatibility of approximate solution of linear equations with given error bounds for coefficients and right-hand sides. Numer. Math., 6:405–409, 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oettli:1965:ASL

[OPW65] W. Oettli, W. Prager, and J. H.

Wilkinson. Admissible solutions of linear systems with not sharply defined coefficients. SIAM J. Numer. Anal., 2(2): 291–299, 1965.

Ortega:1963:EAH

[Ort63] James M. Ortega. An error analysis of Householder’s method for the symmetric eigenvalue problem. Numer. Math., 5:211– 225, 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ortega:1972:NAS

[Ort72] James M. Ortega. Numerical Analysis: A Second Course. Academic Press, New York, 1972. ISBN 0-89871-250-5 (original), 0-89871-250-5 (SIAM). xiii + 201 pp. Republished by SIAM (1990).

Ostrowski:1973:SEE

[Ost73] A. M. Ostrowski. Solution of Equations in Euclidean and Banach Spaces. Academic Press, New York, 1973. xx + 412 pp. Third edition of Solution of Equations and Systems of Equations.

Overton:2001:NCI

[Ove01] Michael L. Overton. Numerical Computing with IEEE Floating Point Arithmetic, Including One Theorem, One Rule of Thumb, and One Hundred and One Exercises. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-482-6 (hardcover), 0-89871-571-7 (paperback), 0-89871-807-4 (ebook). xiv + 104 pp. LCCN QA76.9.M35 O94 2001. US$40.00. URL http://

www.cs.nyu.edu/cs/faculty/
overton/book/;   http://www.
siam.org/catalog/mcc07/ot76.
htm.

Olver:1982:PEB

[OW82] F. W. J. Olver and J. H. Wilkinson. A posteriori error bounds for Gaussian elimination. IMA J. Numer. Anal., 2:377–406, 1982.

Paige:1973:EAM

[Pai73] C. C. Paige. An error analysis of a method for solving matrix equations. Math. Comp., 27 (122):355–359, 1973.

Paige:1990:SAG

[Pai90] C. C. Paige. Some aspects of generalized QR factorizations. In M. G. Cox and S. J. Hammarling, editors, Reliable Numerical Computation, pages 73–91. Oxford University Press, 1990.

Pan:1978:SAO

[Pan78] Victor Y. Pan. Strassen algorithm is not optimal. Trilinear technique of aggregating, uniting and canceling for constructing fast algorithms for matrix multiplication. In Proc. 19th Annual Symposium on the Foundations of Computer Science, pages 166–176. Ann Arbor, MI, USA, 1978.

Pan:1984:HCW

[Pan84]  Victor Y. Pan.  How can

we speed up matrix multiplication? SIAM Rev., 26(3):393– 415, 1984.

Pan:2000:ECR

[Pan00] C.-T. Pan. On the existence and computation of rank-revealing LU factorizations. Linear Algebra Appl., 316:199–222, 2000.

Pan:2001:SMP

[Pan01]  Victor Y. Pan.  Structured

Matrices and Polynomials: Unified Superfast Algorithms. Birkhäuser, Boston, MA, USA, 2001. ISBN 0-8176-4240-4. xxiv + 278 pp.

Parlett:1964:LMA

[Par64] Beresford N. Parlett. Laguerre’s method applied to the matrix eigenvalue problem. Math. Comp., 18:464–485, 1964.

Parlett:1965:MEP

[Par65] Beresford N. Parlett. Matrix eigenvalue problems. Amer. Math. Monthly, 72(2):59–66, February 1965.

Parlett:1971:AAR

[Par71] Beresford N. Parlett. Analysis of algorithms for reflections in bisectors. SIAM Rev., 13(2):197– 208, 1971.

Parlett:1990:CJH

[Par90]  Beresford N. Parlett.  The

contribution of J. H. Wilkinson to numerical analysis. In Stephen G. Nash, editor, A History of Scientific Computing, pages 17–30. Addison-Wes-ley, Reading, MA, USA, 1990.

Parlett:1998:SEP

[Par98]  Beresford N. Parlett.  The

Symmetric Eigenvalue Problem, volume 20 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-402-8 (paperback), 1-61197-116-0 (e-book). xxiv + 398 pp. LCCN QA188 .P37 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0708/97040623-

d.html. Corrected reprint [?].

Parhami:2000:CAA

[Par00] Behrooz Parhami. Computer Arithmetic: Algorithms and Hardware Designs. Oxford University Press, New York, 2000. ISBN 0-19-512583-5 (hardcover), 3-540-04907-X (print), 3-540-36246-0 (e-book). xx + 490 pp. LCCN QA76.9.C62P37 1999. US$85.00.

Paxson:1991:PTI

[Pax91] Vern Paxson. A program for testing IEEE decimal–binary conversion. Manuscript. ftp:

//ftp.ee.lbl.gov/testbase-
report.ps.Z, May 1991.

Pan:1999:NIS

[PBRZ99] Victor Y. Pan, Sheryl Branham, Rhys E. Rosholt, and Ai-Long Zheng. Newton’s iteration for structured matrices. In Thomas Kailath and Ali H. Sayed, editors, Fast Reliable Algorithms for Matrices with Structure, pages 189–210. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1999.

Puschmann:1983:CPR

[PC83] Heinrich Puschmann and Joaquín Cortés. The coordinex problem and its relation to the conjecture of Wilkinson. Numer. Math., 42:291–297, 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parlett:2000:RRR

[PD00] Beresford N. Parlett and Inderjit S. Dhillon. Relatively robust representations of symmetric tridiagonals. Linear Algebra Appl., 309:121–151, 2000.

Pena:1996:PSL

[Peñ96] J. M. Peña. Pivoting strategies leading to small bounds of the errors for certain linear systems. IMA J. Numer. Anal., 16 (2):141–153, 1996.

Pena:1998:PSL

[Peñ98a] J. M. Peña. Pivoting strategies leading to diagonal dominance by rows. Numer. Math., 81:293– 304, 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Penzl:1998:NSG

[Pen98b] Thilo Penzl. Numerical solution of generalized Lyapunov equations. Adv. in Comput. Math., 8:33–48, 1998.

Percival:2002:RMM

[Per02] Colin Percival. Rapid multiplication modulo the sum and difference of highly composite numbers. Math. Comp., 2002. To appear.

Petersen:1981:ET

[Pet81] Karl Petersen. Ergodic Theory. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-23632-0. xi + 329 pp.

Patterson:1998:COD

[PH98] David A. Patterson and John L. Hennessy. Computer Organization and Design: The Hardware/Software Interface. Morgan Kaufmann, San Francisco, CA, USA, second edition, 1998. ISBN 1-55860-428-6. xxix + 759 + appendices pp.

Pichat:1972:CSA

[Pic72] M. Pichat. Correction d’une somme en arithmétique à virgule flottante. Numer. Math., 19:400–406, 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peitgen:1992:FCP

[PJS92] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Part One: Introduction to Fractals and Chaos. Springer-Verlag, New York, 1992. ISBN 0-387-97041-X. xiv + 450 pp.

Plauger:1988:PFP

[Pla88] P. J. Plauger. Properties of floating-point arithmetic. Computer Language Magazine, 5(3): 17–22, 1988. CODEN COM-LEF. ISSN 0749-2839.

Plemmons:1974:LLS

[Ple74] Robert J. Plemmons. Linear least squares by elimination and MGS. J. Assoc. Comput. Mach., 21(4):581–585, 1974.

Plemmons:1976:RSD

[Ple76] R. J. Plemmons. Regular splittings and the discrete Neumann problem. Numer. Math., 25:153–161, 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Puschmann:1985:ZUV

[PN85] Heinrich Puschmann and Marcelo Nordio. Zwei Unzalässige Verstärkungen der Vermutung von Wilkinson. Linear Algebra Appl., 72:167–176, 1985. In German.

Poole:2000:RPS

[PN00] George Poole and Larry Neal. The rook’s pivoting strategy. J. Comput. Appl. Math., 123:353– 369, 2000.

Polman:1987:IBF

[Pol87] Ben Polman. Incomplete blockwise factorizations of (block) Hmatrices. Linear Algebra Appl., 90:119–132, 1987.

Powell:1970:SNM

[Pow70] M. J. D. Powell. A survey of numerical methods for unconstrained optimization. SIAM Rev., 12(1):79–97, 1970.

Powell:1975:VUM

[Pow75] M. J. D. Powell. A view of unconstrained minimization algorithms that do not require derivatives. ACM Trans. Math. Software, 1(2):97–107, 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Power:1986:CDH

[Pow86] Stephen Power. The Cholesky decomposition in Hilbert space. IMA Bulletin, 22(11/12):186– 187, 1986.

Powell:1998:DSA

[Pow98]  M. J. D. Powell.  Direct

search algorithms for optimization calculations. Acta Numerica, 7:287–336, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Pierce:1992:FAC

[PP92] Daniel J. Pierce and Robert J.

Plemmons. Fast adaptive condition estimation. SIAM J. Matrix Anal. Appl., 13(1):274–291, 1992.

Powell:1969:AHT

[PR69] M. J. D. Powell and J. K.

Reid. On applying Householder transformations to linear least squares problems. In Proc. IFIP Congress 1968, pages 122– 126. North-Holland, Amsterdam, The Netherlands, 1969.

Poljak:1993:CRN

[PR93] Svatopluk Poljak and Jiří Rohn.

Checking robust nonsingularity is NP-hard. Mathematics of control, signals, and systems: MCSS, 6:1–9, 1993. CODEN MCSYE8. ISSN 0932-4194 (print), 1435-568x (electronic).

Priest:1991:AAP

[Pri91] Douglas M. Priest. Algorithms for arbitrary precision floating point arithmetic. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 132–143. IEEE Computer Society Press, Los Alamitos, CA, USA, 1991.

Priest:1992:PFP

[Pri92] Douglas M. Priest. On Properties of Floating Point Arithmetics: Numerical Stability and the Cost of Accurate Computations. PhD thesis, Mathematics Department, University of California, Berkeley, CA, USA, November 1992. 126 pp. ftp://

                                        ftp.icsi.berkeley.edu/pub/
                                        theory/priest-thesis.ps.Z.

Pryce:1981:REA

[Pry81] J. D. Pryce. Round-off error analysis with fewer tears. IMA Bulletin, 17:40–47, 1981.

Pryce:1984:NMR

[Pry84] J. D. Pryce. A new measure of relative error for vectors. SIAM J. Numer. Anal., 21(1):202–215, 1984.

Pryce:1985:MEA

[Pry85] J. D. Pryce. Multiplicative error analysis of matrix transformation algorithms. IMA J. Numer. Anal., 5:437–445, 1985.

Paterson:1973:NNM

[PS73] Michael S. Paterson and Larry J. Stockmeyer. On the number of nonscalar multiplications necessary to evaluate polynomials. SIAM J. Comput., 2(1):60–66, 1973.

Pan:1991:INI

[PS91] Victor Y. Pan and Robert Schreiber. An improved Newton iteration for the generalized inverse of a matrix, with applications. SIAM J. Sci. Statist. Comput., 12(5):1109– 1130, 1991.

Press:1992:NRF

[PTVF92] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical Recipes in FORTRAN:

The Art of Scientific Computing. Cambridge University Press, second edition, 1992. ISBN 0-521-43064-X. xxvi + 963 pp.

Puglisi:1992:MHM

[Pug92] Chiara Puglisi. Modification of the Householder method based on the compact WY representation. SIAM J. Sci. Statist. Comput., 13(3):723–726, 1992.

Porte:1974:ESE

[PV74] M. La Porte and J. Vignes.

Etude statistique des erreurs dans l’arithmétique des ordinateurs; application au contrôle des Résultats d’algorithmes numériques. Numer. Math., 23:63–72, 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peters:1970:LSP

[PW70] G. Peters and J. H. Wilkinson. The least squares problem and pseudo-inverses. Comput. J., 13(3):309–316, 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Peters:1971:PPA

[PW71] G. Peters and J. H. Wilkinson. Practical problems arising in the solution of polynomial equations. J. Inst. Maths. Applics., 8:16–35, 1971.

Peters:1975:SGJ

[PW75] G. Peters and J. H. Wilkinson. On the stability of Gauss– Jordan elimination with pivoting. Comm. ACM, 18(1):20– 24, 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paige:1994:HGC

[PW94] C. C. Paige and M. Wei. History and generality of the CS decomposition. Linear Algebra Appl., 208/209:303–326, 1994.

Pan:2001:CNC

[PY01] Victor Y. Pan and Y. Yu. Certification of numerical computation of the sign of the determinant of a matrix. Algorithmica, 30(4):708–724, 2001. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Quintana:2001:NPM

[QQSvdG01] Enrique S. Quintana, Gregorio Quintana, Xiaobai Sun, and Robert van de Geijn. A note on parallel matrix inversion. SIAM J. Sci. Comput., 22(5):1762– 1771, 2001.

Quinn:1990:REL

[QT90] Thomas Quinn and Scott Tremaine. Roundoff error in long-term planetary orbit integrations. Astronomical Journal, 99(3):1016–1023, 1990. CO-DEN ANJOAA. ISSN 0004-6256 (print), 1538-3881 (electronic).

Quinn:1991:TMY

[QTD91] Thomas R. Quinn, Scott Tremaine, and Martin Duncan. A three million year integration of the earth’s orbit. Astronomical Journal, 101(6):2287– 2305, 1991. CODEN ANJOAA. ISSN 0004-6256 (print), 1538-3881 (electronic).

Quinn:1983:EPR

[Qui83] Kevin Quinn. Ever had problems rounding off figures? This stock exchange has. Wall Street Journal, 1983. CODEN WSJOAF. ISSN 0099-9660. 8 November.

Quinlan:1994:REL

[Qui94] Gerald D. Quinlan. Round-off error in long-term orbital integrations using multistep methods. Celestial Mechanics and Dynamical Astronomy, 58:339– 351, 1994. CODEN CLMCAV. ISSN 0923-2958 (print), 1572-9478 (electronic).

Raimi:1969:PDF

[Rai69] Ralph A. Raimi. The peculiar distribution of first digits. Scientific American, 221(6): 109–120, 1969. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Raimi:1976:FDP

[Rai76] Ralph A. Raimi. The first digit problem. Amer. Math. Monthly, 83:521–538, 1976.

Rall:1981:ADT

[Ral81] Louis B. Rall. Automatic Differentiation: Techniques and Applications, volume 120 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, 1981. viii + 165 pp.

Rall:1991:TMC

[Ral91]  Louis B. Rall.  Tools for

mathematical computation. In Kenneth R. Meyer and Dieter S. Schmidt, editors, Computer Aided Proofs in Analysis, volume 28 of IMA Volumes in Mathematics and Its Applications, pages 217–228. Springer-Verlag, New York, 1991.

Ramos:1971:REA

[Ram71] George U. Ramos. Roundoff error analysis of the fast Fourier transform. Math. Comp., 25 (116):757–768, 1971.

Randell:1975:ODC

[Ran75] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Springer-Ver-lag, Berlin, third edition, 1975. ISBN 3-540-11319-3. xvi + 580 pp.

Rath:1982:FGR

[Rat82] Wolfgang Rath. Fast Givens rotations for orthogonal similarity transformations. Numer. Math., 40:47–56, 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reid:1971:NSG

[Rei71] J. K. Reid. A note on the stability of Gaussian elimination. J. Inst. Maths. Applics., 8:374– 375, 1971.

Reid:1987:SM

[Rei87] J. K. Reid. Sparse matrices. In A. Iserles and M. J. D. Powell, editors, The State of the Art in Numerical Analysis, pages 59–85. Oxford University Press, 1987.

Reichel:1990:NIL

[Rei90] Lothar Reichel. Newton interpolation at Leja points. BIT, 30:332–346, 1990. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Reichel:1991:FQD

[Rei91] Lothar Reichel. Fast QR decomposition of Vandermonde-like matrices and polynomial least squares approximation. SIAM J. Matrix Anal. Appl., 12(3): 552–564, 1991.

Reid:2000:ISL

[Rei00] J. K. Reid. Implicit scaling of linear least squares problems. BIT, 40(1):146–157, 2000. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Rigal:1967:CGS

[RG67] J. L. Rigal and J. Gaches. On the compatibility of a given solution with the data of a linear system. J. Assoc. Comput. Mach., 14(3):543–548, 1967.

Rheinboldt:1976:MIC

[Rhe76] Werner C. Rheinboldt. On measures of ill-conditioning for nonlinear equations. Math. Comp., 30(133):104–111, 1976.

Rice:1966:EGS

[Ric66a] John R. Rice. Experiments on Gram–Schmidt orthogonalization. Math. Comp., 20:325– 328, 1966.

Rice:1966:TC

[Ric66b] John R. Rice. A theory of condition. SIAM J. Numer. Anal., 3(2):287–310, 1966.

Reiser:1975:EDF

[RK75] John F. Reiser and Donald E. Knuth. Evading the drift in floating-point addition. Inform. Process. Lett., 3(3):84–87, 1975.

Reichel:1991:CVS

[RO91] Lothar Reichel and Gerhard Opfer. Chebyshev–Vandermonde systems. Math. Comp., 57(196): 703–721, 1991.

Roberts:1980:LMR

[Rob80] J. D. Roberts. Linear model reduction and solution of the algebraic Riccati equation by use of the sign function. Internat. J. Control, 32(4):677–687, 1980. First issued as report CUED/B-Control/TR13, Department of Engineering, University of Cambridge, 1971.

Rohn:1989:NCN

[Roh89a] Jiří Rohn. New condition numbers for matrices and linear systems. Computing: Archiv für informatik und numerik, 41:167– 169, 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Rohn:1989:SLI

[Roh89b] Jiří Rohn. Systems of linear interval equations. Linear Algebra Appl., 126:39–78, 1989.

Rohn:1990:NUD

[Roh90] Jiří Rohn. Nonsingularity under data rounding. Linear Algebra Appl., 139:171–174, 1990.

Rohn:1995:NHR

[Roh95] Jiří Rohn. NP-hardness results for some linear and quadratic problems. Technical Report No. 619, Institute of Computer Science, Academy of Sciences of the Czech Republic, Prague, January 1995. 11 pp.

Ross:1965:RTE

[Ros65] D. R. Ross. Reducing truncation errors using cascading accumulators. Comm. ACM, 8(1):32– 33, 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rowan:1990:FSA

[Row90] Thomas Harvey Rowan. Functional Stability Analysis of Numerical Algorithms. PhD thesis, University of Texas at Austin, Austin, TX, USA, May 1990. xii + 205 pp.

Robertazzi:1988:BOF

[RS88] T. G. Robertazzi and S. C.

Schwartz. Best “ordering” for floating-point addition. ACM Trans. Math. Software, 14(1): 101–110, 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Routh:1977:PSM

[RSD77] M. W. Routh, P. A. Swartz, and M. B. Denton. Performance of the super modified simplex. Analytical Chemistry, 49(9):1422– 1428, 1977. CODEN ANCHAM. ISSN 0003-2700 (print), 1520-6882 (electronic).

Reddy:1992:SML

[RT92a] Satish C. Reddy and Lloyd N. Trefethen. Stability of the method of lines. Numer. Math., 62:235–267, 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reichel:1992:EPE

[RT92b] Lothar Reichel and Lloyd N. Trefethen. Eigenvalues and pseudo-eigenvalues of Toeplitz matrices. Linear Algebra Appl., 162–164:153–185, 1992.

Ruhe:1983:NAG

[Ruh83] Axel Ruhe. Numerical aspects of Gram–Schmidt orthogonalization of vectors. Linear Algebra Appl., 52/53:591–601, 1983.

Rump:1998:SPS

[Rum98] Siegfried M. Rump. Structured perturbations and symmetric matrices. Linear Algebra Appl., 278:121–132, 1998.

Rump:1999:FPI

[Rum99a] Siegfried M. Rump. Fast and parallel interval arithmetic. BIT, 39(3):534–554, 1999. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Rump:1999:ICM

[Rum99b] Siegfried M. Rump. Illconditioned matrices are componentwise near to singularity. SIAM Rev., 41(1):102–112, 1999.

Rump:1999:IIL

[Rum99c] Siegfried M. Rump. INTLAB— INTerval LABoratory. In Tibor Csendes, editor, Developments in Reliable Computing, pages 77–104. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1999.

Rump:2001:SVM

[Rum01]  Siegfried M. Rump.  Self-

validating methods. Linear Algebra Appl., 324:3–13, 2001.

Rump:2002:OSN

[Rum02] Siegfried M. Rump. Optimal scaling for P -norms and componentwise distance to singularity. IMA J. Numer. Anal., 2002. To appear.

Russinoff:1998:MCP

[Rus98] David M. Russinoff. A mechanically checked proof of IEEE compliance of the floating point multiplication, division and square root algorithms of the AMD-K7 processor. LMS J. Computer Math., 1:148–200, 1998.

Rutherford:1947:SCD

[Rut47] D. E. Rutherford. Some continuant determinants arising in physics and chemistry. Proceedings of the Royal Society of Edinburgh, 62,A:229–236, 1947. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Rutherford:1952:SCD

[Rut52] D. E. Rutherford. Some continuant determinants arising in physics and chemistry—II. Proceedings of the Royal Society of Edinburgh, 63,A:232–241, 1952. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Rutishauser:1958:SEP

[Rut58] Heinz Rutishauser. Solution of eigenvalue problems with the LR-transformation. In Further Contributions to the Solution of Simultaneous Linear Equations and the Determination of Eigenvalues, number 49 in Applied Mathematics Series, pages 47– 81. National Bureau of Standards, United States Department of Commerce, Washington, D. C., 1958.

Rutishauser:1968:TM

[Rut68]  H. Rutishauser.  On test

matrices. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 (no.

165) of Éditions Centre Nat. Recherche Sci., Paris, pages 349–365, 1968.

Stegun:1956:PC

[SA56] Irene A. Stegun and Milton Abramowitz. Pitfalls in computation. J. Soc. Indust. Appl. Math., 4(4):207–219, 1956.

shal73

[SA73] Lawrence F. Shampine and Richard C. Allen, Jr. Numerical Computing: An Introduction. W. B. Saunders, Philadelphia, PA, USA, 1973. ISBN 0-7216-8150-6. vii + 258 pp.

Swartzlander:1975:SLN

[SA75] Earl E. Swartzlander, Jr. and Aristides G. Alexopolous. The sign/logarithm number system. IEEE Trans. Comput., C-24 (12):1238–1242, 1975.

Saunders:1972:LSL

[Sau72] M. A. Saunders. Large-scale linear programming using the Cholesky factorization. Report CS 252, Department of Computer Science, Stanford University, January 1972.

Sautter:1978:EAG

[Sau78] Werner Sautter. Error analysis of Gauss elimination for the best least squares problem. Numer. Math., 30:165–184, 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Samelson:1953:ORR

[SB53] Klaus Samelson and Friedrich L.

Bauer. Optimale Rechengenauigkeit bei Rechenanlagen mit gleitendem Komma. Z. Angew. Math. Phys., 4:312–316, 1953.

Sameh:1977:STS

[SB77] Ahmed H. Sameh and Richard P. Brent. Solving triangular systems on a parallel computer. SIAM J. Numer. Anal., 14(6): 1101–1113, 1977.

Stoer:1980:INA

[SB80] Josef Stoer and R. Bulirsch. Introduction to Numerical Analysis. Springer-Verlag, New York, 1980. ISBN 0-387-90420-4. ix + 609 pp.

Shroff:1992:ACE

[SB92] Gautam M. Shroff and Christian H. Bischof. Adaptive condition estimation for rank-one updates of QR factorizations. SIAM J. Matrix Anal. Appl., 13 (4):1264–1278, 1992.

Sharangpani:1994:SAF

[SB94] H. P. Sharangpani and M. L. Barton. Statistical analysis of floating point flaw in the Pentium processor (1994). Technical report, Intel Corporation, November 1994. 31 pp.

Sun:1995:BKR

[SB95] Xiaobai Sun and Christian H. Bischof. A basis-kernel representation of orthogonal matrices. SIAM J. Matrix Anal. Appl., 16 (4):1184–1196, 1995.

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines—EISPACK Guide, volume 6 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, second edition, 1976. ISBN 3-540-06710-8. vii + 551 pp.

Scarborough:1950:NMA

[Sca50] James B. Scarborough. Numerical Mathematical Analysis. Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1950. xviii + 511 pp.

schu33

[Sch33]  Günther Schulz.  Iterative

Berechnung der reziproken Matrix. Z. Angew. Math. Mech., 13:57–59, 1933.

Schryer:1981:TCF

[Sch81] N. L. Schryer. A test of a computer’s floating-point arithmetic unit. Computing Science Technical Report No. 89, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981.

Schelin:1983:CFA

[Sch83] Charles W. Schelin. Calculator function approximation. Amer. Math. Monthly, 90(5):317–325, 1983.

Schryer:1984:DCF

[Sch84] N. L. Schryer. Determination of correct floating-point model parameters. In Wayne R. Cowell, editor, Sources and Development of Mathematical Software, pages 360–366. Prentice-Hall, Englewood Cliffs, NJ, USA, 1984.

Schreiber:1988:BAP

[Sch88] Robert S. Schreiber. Block algorithms for parallel machines. In M. H. Schultz, editor, Numerical Algorithms for Modern Parallel Computer Architectures, number 13 in IMA Volumes In Mathematics and Its Applications, pages 197–207. Springer-Verlag, Berlin, 1988.

Schreiber:1989:HRS

[Sch89] Robert S. Schreiber. Hough’s random story explained. NA Digest, Volume 89, Issue 3, 1989.

Schatzman:1996:ADF

[Sch96] James C. Schatzman. Accuracy of the discrete Fourier transform and the fast Fourier transform. SIAM J. Sci. Comput., 17(5): 1150–1166, 1996.

Sedgewick:1988:A

[Sed88] Robert Sedgewick. Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 0-201-06673-4. xii + 657 pp. LCCN QA76.6 .S435 1988. US$34.95.

Severance:1998:IOM

[Sev98] Charles Severance. An interview with the old man of floating-point. Reminiscences elicited from William Kahan by Charles Severance, Febru-

                                        ary 1998.  http://www.cs.
                                        berkeley.edu/~wkahan/ieee754status/
                                        754story.html. A shortened

version appears in Computer, 31(3):114–115, March 1998.

Shapiro:1982:OSM

[Sha82] Alexander Shapiro. Optimally scaled matrices, necessary and sufficient conditions. Numer. Math., 39:239–245, 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shapiro:1985:OBD

[Sha85] Alexander Shapiro. Optimal block diagonal l_{2}-scaling of matrices. SIAM J. Numer. Anal., 22(1):81–94, 1985.

Shapiro:1991:UBN

[Sha91]  Alexander Shapiro.  Upper

bounds for nearly optimal diagonal scaling of matrices. Linear and Multilinear Algebra, 29:147, 1991.

Shampine:1994:NSO

[Sha94] Lawrence F. Shampine. Numerical Solution of Ordinary Differential Equations. Chapman and Hall, New York, 1994. ISBN 0-412-05151-6. x + 484 pp.

Shewchuk:1997:APF

[She97] Jonathan Richard Shewchuk.

Adaptive precision floatingpoint arithmetic and fast robust geometric predicates. Discrete and Computational Geometry, 18(3):305–363, 1997. CO-DEN DCGEER. ISSN 0179-5376 (print), 1432-0444 (electronic).

Skeel:1993:ENC

[SK93] Robert D. Skeel and Jerry B.

Keiper. Elementary Numerical Computing with Mathematica. McGraw-Hill, New York, 1993. ISBN 0-07-057820-6. xiv + 434 pp.

Skeel:1979:SNS

[Ske79] Robert D. Skeel. Scaling for numerical stability in Gaussian elimination. J. Assoc. Comput. Mach., 26(3):494–526, 1979.

Skeel:1980:IRI

[Ske80] Robert D. Skeel. Iterative refinement implies numerical stability for Gaussian elimination. Math. Comp., 35(151):817–832, 1980.

Skeel:1981:EER

[Ske81] Robert D. Skeel. Effect of equilibration on residual size for partial pivoting. SIAM J. Numer. Anal., 18(3):449–454, 1981.

Skeel:1989:SNB

[Ske89] Robert D. Skeel. Safety in numbers: The boundless errors of numerical computation. Working Document 89-3, Department of Computer Science, University of Illinois, Urbana, IL, USA, 1989. 9 pp.

Savage:1954:TIF

[SL54] I. Richard Savage and Eugene Lukacs. Tables of inverse of finite segments of the Hilbert matrix. In Olga Taussky, editor, Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, number 39 in Applied Mathematics Series, pages 105– 108. National Bureau of Standards, United States Department of Commerce, Washington, D. C., 1954.

Smale:1990:SRF

[Sma90] Steve Smale. Some remarks on the foundations of numerical analysis. SIAM Rev., 32(2):211– 220, 1990.

Smith:1962:ACD

[Smi62] Robert L. Smith. Algorithm 116: Complex division. Comm. ACM, 5(8):435, 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:1965:ASO

[Smi65] Francis J. Smith. An algorithm for summing orthogonal polynomial series and their derivatives with applications to curvefitting and interpolation. Math. Comp., 19:33–36, 1965.

Smith:1975:SAP

[Smi75] Jon M. Smith. Scientific Analysis on the Pocket Calculator. Wiley, New York, 1975. ISBN 0-471-79997-1. xii + 380 pp.

Smith:1991:AFP

[Smi91] David M. Smith. Algorithm 693: A FORTRAN package for floating-point multiple-precision arithmetic. ACM Trans. Math. Software, 17(2):273–283, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smith:1998:AMP

[Smi98] David M. Smith. Algorithm 786: Multiple-precision complex arithmetic and functions. ACM Trans. Math. Software, 24(4): 359–367, 1998. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smoktunowicz:1995:NSC

[Smo95] Alicja Smoktunowicz. A note on the strong componentwise stability of algorithms for solving symmetric linear systems. Demonstratio Mathematica, 28 (2):443–448, 1995. CODEN DE-MADO. ISSN 0420-1213.

Starke:1991:S

[SN91] Gerhard Starke and Wilhelm Niethammer. SOR for AX−XB = C. Linear Algebra Appl., 154– 156:355–375, 1991.

Snyder:1955:ISS

[Sny55] James N. Snyder. On the improvement of the solutions to a set of simultaneous linear equations using the ILLIAC. Mathematical Tables and Other Aids to Computation, 9:177– 184, 1955. CODEN MTTCAS. ISSN 0891-6837.

Sorensen:1985:APP

[Sor85] D. C. Sorensen. Analysis of pairwise pivoting in Gaussian elimination. IEEE Trans. Comput., C-34:274–278, 1985.

Schreiber:1988:BRT

[SP88] Robert S. Schreiber and Beresford N. Parlett. Block reflectors:

Theory and computation. SIAM J. Numer. Anal., 25(1):189–205, 1988.

Spiess:1976:UZD

[Spi76] J. Spieß. Untersuchungen des Zeitgewinns durch neue Algorithmen zur Matrix-Multiplikation. Computing: Archiv für informatik und numerik, 17:23–36, 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Schonfelder:1980:ECP

[SR80] J. L. Schonfelder and M. Razaz.

Error control with polynomial approximations. IMA J. Numer. Anal., 1:105–114, 1980.

Shampine:1997:MOS

[SR97] Lawrence F. Shampine and Mark W. Reichelt. The MAT-LAB ODE suite. SIAM J. Sci. Comput., 18(1):1–22, 1997.

Schneider:1962:CTS

[SS62] Hans Schneider and W. Gilbert Strang. Comparison theorems for supremum norms. Numer. Math., 4:15–20, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Soderstrom:1974:NPI

[SS74] Torsten Söderström and G. W.

Stewart. On the numerical properties of an iterative method for computing the Moore–Penrose generalized inverse. SIAM J. Numer. Anal., 11(1):61–74, 1974.

Smoktunowicz:1984:BCI

[SS84] Alicja Smoktunowicz and Jolanta Sokolnicka. Binary cascades iterative refinement in doubledmantissa arithmetics. BIT, 24: 123–127, 1984. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Stewart:1990:MPT

[SS90] Gilbert Wright (Gilbert W.) Stewart and Ji-guang Sun. Matrix Perturbation Theory. Computer science and scientific computing. Academic Press, New York, NY, USA, 1990. ISBN 0-12-670230-6 (hardcover). xv + 365 pp. LCCN QA871 .S775 1990. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/21227976.
html;   http://www.loc.gov/
catdir/description/els032/
90033378.htm;   http://www.
loc.gov/catdir/description/
els032/90033378.html;   http:
//www.loc.gov/catdir/toc/
els031/90033378.htm;   http:
//www.loc.gov/catdir/toc/
els031/90033378.html.

Sun:1997:OBPb

[SS97] Ji-guang Sun and Zheng Sun. Optimal backward perturbation bounds for underdetermined systems. SIAM J. Matrix Anal. Appl., 18(2):393–402, 1997.

Sanz-Serna:1993:SCS

[SSL93] J. M. Sanz-Serna and S. Larsson. Shadows, chaos, and saddles. Appl. Numer. Math., 13: 181–190, 1993.

Stewart:1971:EAA

[Ste71] G. W. Stewart. Error analysis of the algorithm for shifting the zeros of a polynomial by synthetic division. Math. Comp., 25(113): 135–139, 1971.

Stewart:1973:EPB

[Ste73a] G. W. Stewart. Error and perturbation bounds for subspaces associated with certain eigenvalue problems. SIAM Rev., 15 (4):727–764, 1973.

Stewart:1973:IMC

[Ste73b] G. W. Stewart. Introduction to Matrix Computations. Academic Press, New York, 1973. ISBN 0-12-670350-7. xiii + 441 pp.

Sterbenz:1974:FPC

[Ste74a] Pat H. Sterbenz. Floating-Point Computation. Prentice-Hall, Englewood Cliffs, NJ, USA, 1974. ISBN 0-13-322495-3. xiv + 316 pp.

Stewart:1974:MPE

[Ste74b] G. W. Stewart. Modifying pivot elements in Gaussian elimination. Math. Comp., 28(126): 537–542, 1974.

Stewart:1977:PPI

[Ste77a] G. W. Stewart. On the perturbation of pseudo-inverses, projections and linear least squares problems. SIAM Rev., 19(4): 634–662, 1977.

Stewart:1977:PBF

[Ste77b] G. W. Stewart. Perturbation bounds for the QR factorization of a matrix. SIAM J. Numer. Anal., 14(3):509–518, 1977.

Stewart:1977:RDL

[Ste77c] G. W. Stewart. Research, development, and LINPACK. In John R. Rice, editor, Mathematical Software III, pages 1– 14. Academic Press, New York, 1977.

Stewart:1980:EGR

[Ste80] G. W. Stewart. The efficient generation of random orthogonal matrices with an application to condition estimators. SIAM J. Numer. Anal., 17(3):403–409, 1980.

Stewart:1985:NCD

[Ste85] G. W. Stewart. A note on complex division. ACM Trans. Math. Software, 11(3):238–241, 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stewart:1990:SPT

[Ste90] G. W. Stewart. Stochastic perturbation theory. SIAM Rev., 32 (4):579–610, 1990.

Stewart:1992:NGS

[Ste92] G. W. Stewart. Note on a generalized Sylvester equation. IMA Preprint Series #985, Institute for Mathematics and Its Applications, University of Minnesota, Minneapolis, MN, USA, May 1992. 3 pp.

Stewart:1993:EHS

[Ste93a] G. W. Stewart. On the early history of the singular value decomposition. SIAM Rev., 35(4):551– 566, 1993.

Stewart:1993:PLC

[Ste93b] G. W. Stewart. On the perturbation of LU, Cholesky, and QR factorizations. SIAM J. Matrix Anal. Appl., 14(4):1141– 1145, 1993.

Stewart:1995:GSG

[Ste95] G. W. Stewart. Gauss, statistics, and Gaussian elimination. Journal of Computational and Graphical Statistics, 4(1):1–11, 1995.

Stewart:1997:MCS

[Ste97a] G. W. Stewart. On Markov chains with sluggish transients. Communication in Statistics, Stochastic Models, 13:85–95, 1997.

Stewart:1997:PCF

[Ste97b] G. W. Stewart. On the perturbation of LU and Cholesky factors. IMA J. Numer. Anal., 17 (1):1–6, 1997.

Stewart:1997:TMG

[Ste97c] G. W. Stewart. The triangular matrices of Gaussian elimination and related decompositions. IMA J. Numer. Anal., 17 (1):7–16, 1997.

Stewart:1997:CFS

[Ste97d]  Michael Stewart.  Cholesky

factorization of semidefinite Toeplitz matrices. Linear Algebra Appl., 254:497–525, 1997.

Stewart:1998:AM

[Ste98] G. W. Stewart. On the adjugate matrix. Linear Algebra Appl., 283:151–164, 1998.

Stewart:2001:MAV

[Ste01] G. W. Stewart. Matrix Algorithms. Volume II: Eigensystems. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-503-2. xix + 469 pp.

Stone:1962:BPR

[Sto62] Betty Jane Stone. Best possible ratios of certain matrix norms. Numer. Math., 4:114–116, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1972:ENM

[Sto72] Josef Stoer. Einführung in die Numerische Mathematik I. Springer-Verlag, Berlin, 1972. ISBN 0-387-05750-1. ix + 250 pp.

Stoutemyer:1977:AEA

[Sto77] David R. Stoutemyer. Automatic error analysis using computer algebraic manipulation. ACM Trans. Math. Software, 3 (1):26–43, 1977. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Stoutemyer:1991:CMC

[Sto91] David R. Stoutemyer. Crimes and misdemeanours in the computer algebra trade. Notices Amer. Math. Soc., 38(7):778– 785, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Strassen:1969:GEO

[Str69] V. Strassen. Gaussian elimination is not optimal. Numer. Math., 13:354–356, 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strang:1986:PTM

[Str86] Gilbert Strang. A proposal for Toeplitz matrix calculations. Studies in Applied Mathematics, 74:171–176, 1986. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic).

Strang:1993:ILA

[Str93] Gilbert Strang. Introduction to Linear Algebra. Wellesley-Cambridge Press, Wellesley, MA, USA, 1993. ISBN 0-9614088-5-5. viii + 472 pp.

Stummel:1980:REA

[Stu80] F. Stummel. Rounding error analysis of elementary numerical algorithms. Computing: Archiv für informatik und numerik, Suppl. 2:169–195, 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stummel:1981:PTE

[Stu81] F. Stummel. Perturbation theory for evaluation algorithms of arithmetic expressions. Math. Comp., 37(156):435–473, 1981.

Stummel:1982:OEE

[Stu82] F. Stummel. Optimal error estimates for Gaussian elimination in floating-point arithmetic. Z. Angew. Math. Mech., 62:T355– T357, 1982.

Stummel:1985:SOE

[Stu85a] F. Stummel. Strict optimal error estimates for Gaussian elimination. Z. Angew. Math. Mech., 65:T405–T407, 1985.

Stummel:1985:FEAa

[Stu85b] Friedrich Stummel. Forward error analysis of Gaussian elimination, Part I: Error and residual estimates. Numer. Math., 46:365–395, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stummel:1985:FEAb

[Stu85c] Friedrich Stummel. Forward error analysis of Gaussian elimination, Part II: Stability theorems. Numer. Math., 46:397– 415, 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1991:PBC

[Sun91] Ji-guang Sun. Perturbation bounds for the Cholesky and QR factorizations. BIT, 31: 341–352, 1991. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

sunM92

[Sun92a]  SPARCompiler  FORTRAN:

Numerical Computation Guide. Sun Microsystems, Inc., Mountain View, CA, USA, October 1992. Part No. 800-7097-11, Revision A.

sunM92a

[Sun92b] SPARCompiler FORTRAN 2.0.1:

User’s Guide. Sun Microsystems, Inc., Mountain View, CA, USA, October 1992. Part No. 800-6552-11, Revision A.

Sun:1992:CPB

[Sun92c] Ji-guang Sun. Componentwise perturbation bounds for some matrix decompositions. BIT, 32(4):702–714, 1992. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1992:REP

[Sun92d] Ji-guang Sun. Rounding-error and perturbation bounds for the Cholesky and LDL^{T} factorizations. Linear Algebra Appl., 173: 77–97, 1992.

Sun:1995:NBP

[Sun95a] Ji-guang Sun. A note on backward perturbations for the Hermitian eigenvalue problem. BIT, 35:385–393, 1995. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1995:PBQ

[Sun95b] Ji-guang Sun. On perturbation bounds for the QR factorization.

Linear Algebra Appl., 215:95– 111, 1995.

Sun:1996:OBP

[Sun96] Ji-guang Sun. Optimal backward perturbation bounds for the linear least-squares problem with multiple right-hand sides. IMA J. Numer. Anal., 16(1):1– 11, 1996.

Sun:1997:OBPa

[Sun97] Ji-guang Sun. On optimal backward perturbation bounds for the linear least squares problem. BIT, 37(1):179–188, 1997. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Sun:1998:BSB

[Sun98] Ji-guang Sun. Bounds for the structured backward errors of Vandermonde systems. SIAM J. Matrix Anal. Appl., 20(1):45– 59, 1998.

Schreiber:1989:SEW

[SV89] Robert S. Schreiber and Charles F. Van Loan. A storage efficient WY representation for products of Householder transformations. SIAM J. Sci. Statist. Comput., 10:53–57, 1989.

Sleijpen:2000:DER

[SvdVM00] Gerard L. G. Sleijpen, Henk A. van der Vorst, and Jan Modersitzki. Differences in the effects of rounding errors in Krylov solvers for symmetric indefinite linear systems. SIAM J. Matrix Anal. Appl., 22(3):726–751, 2000.

Stoer:1962:TDM

[SW62] Josef Stoer and C. Witzgall.

Transformations by diagonal matrices in a normed space. Numer. Math., 4:158–171, 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shepherd:1989:MCW

[SW89] David Shepherd and Greg Wilson. Making chips that work. New Scientist, pages 61–64, 1989. CODEN NWSCAL. ISSN 0262-4079, 0028-6664. 13 May.

Steele:1990:HPF

[SW90] Guy L. Steele, Jr. and Jon L.

White. How to print floatingpoint numbers accurately. ACM SIGPLAN Notices, 25(6):112– 126, 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1998:MWN

[SW98] Hans Schneider and Hans F.

Weinberger. On matrices for which norm bounds are attained. Linear Algebra Appl., 275–276:563–577, 1998.

Swann:1972:DSM

[Swa72] W. H. Swann. Direct search methods. In W. Murray, editor, Numerical Methods for Unconstrained Optimization, pages 13–28. Academic Press, London,

Swann:1974:COD

[Swa74] W. H. Swann. Constrained optimization by direct search. In P. E. Gill and W. Murray, editors, Numerical Methods for Constrained Optimization, pages 191–217. Academic Press, London, 1974.

Swartzlander:1980:CA

[Swa80] Earl E. Swartzlander, Jr., editor. Computer Arithmetic, volume 21 of Benchmark Papers in Electrical Engineering and Computer Science. Dowden, Hutchinson and Ross, Stroudsburg, PA, USA, 1980.

Sweeney:1965:AFP

[Swe65] D. W. Sweeney. An analysis of floating-point addition. IBM Systems Journal, 4:31–42, 1965. CODEN IBMSA7. ISSN 0018-8670. Reprinted in [Swa80, pp. 317–328].

Sylvester:1850:AAN

[Syl50] J. J. Sylvester. Additions to the articles, “On a New Class of Theorems,” and “On Pascal’s Theorem”. Philosophical Magazine, 37:363–370, 1850. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Syl04, pp. 1451– 151].

Sylvester:1851:RBM

[Syl51] J. J. Sylvester. On the relation between the minor determinants of linearly equivalent quadratic functions. Philosophical Magazine, (Fourth Series) 1:295–305, 1851. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Syl04, pp. 241–250].

Sylvester:1884:EM

[Syl84] J. J. Sylvester. Sur l’equation en matrices px = xq. Comptes Rendus de l’Académie des Sciences, pages 67–71 and 115–116, 1884.

sylv04

[Syl04] The Collected Mathematical Papers of James Joseph Sylvester, volume 1 (1837–1853). Cambridge University Press, 1904. xii + 650 pp.

Scherer:1980:SNR

[SZ80] R. Scherer and K. Zeller. Shorthand notation for rounding errors. Computing: Archiv für informatik und numerik, Suppl. 2:165–168, 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Turnbull:1932:ITC

[TA32] H. W. Turnbull and A. C.

Aitken. An Introduction to the Theory of Canonical Matrices. Blackie, London and Glasgow, 1932. xiii + 200 pp. Reprinted with appendix, 1952.

Tang:1989:TDI

[Tan89] Ping Tak Peter Tang. Tabledriven implementation of the exponential function in IEEE floating-point arithmetic. ACM Trans. Math. Software, 15(2): 144–157, 1989. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1990:AET

[Tan90a] Ping Tak Peter Tang. Accurate and efficient testing of the exponential and logarithm functions. ACM Trans. Math. Software, 16(3):185–200, 1990. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1990:TDI

[Tan90b] Ping Tak Peter Tang. Tabledriven implementation of the logarithm function in IEEE floating-point arithmetic. ACM Trans. Math. Software, 16(3): 378–400, 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tang:1991:TLA

[Tan91] Ping Tak Peter Tang. Tablelookup algorithms for elementary functions and their error analysis. In Peter Kornerup and David W. Matula, editors, Proc. 10th IEEE Symposium on Computer Arithmetic, pages 232– 236. IEEE Computer Society Press, Los Alamitos, CA, USA, 1991.

Tang:1992:TDI

[Tan92] Ping Tak Peter Tang. Tabledriven implementation of the expm1 function in IEEE floatingpoint arithmetic. ACM Trans. Math. Software, 18(2):211–222, 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tao:1984:CSM

[Tao84a] Pham Dinh Tao. Convergence of a subgradient method for computing the bound norm of matrices. Linear Algebra Appl., 62: 163–182, 1984. In French.

Tao:1984:SMC

[Tao84b] Pham Dinh Tao. Some methods for computing the maximum of quadratic form on the unit ball of the maximum norm. Numer. Math., 45:377–401, 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). In French.

Taussky:1949:RCC

[Tau49] Olga Taussky. A remark concerning the characteristic roots of the finite segments of the Hilbert matrix. Quarterly Journal of Mathematics, 20:80–83, 1949. CODEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Taub:1963:JNC

[Tau63] A. H. Taub, editor. John von Neumann Collected Works, volume V, Design of Computers, Theory of Automata and Numerical Analysis. Pergamon, Oxford, 1963. ix + 784 pp.

Taussky:1988:HBT

[Tau88] Olga Taussky. How I became a torchbearer for matrix theory. Amer. Math. Monthly, 95 (9):801–812, November 1988.

Trefethen:1997:NLA

[TB97] Lloyd N. Trefethen and David Bau III. Numerical Linear Algebra. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-361-7. xii + 361 pp.

Tang:1981:BDV

[TG81] W. P. Tang and G. H. Golub. The block decomposition of a Vandermonde matrix and its applications. BIT, 21:505– 517, 1981. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Thacher:1961:ACP

[Tha61] Henry C. Thacher, Jr. Algorithm 43: Crout with pivoting II. Comm. ACM, 4(4):176– 177, 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thisted:1988:ESC

[Thi88] Ronald A. Thisted. Elements of Statistical Computing: Numerical Computation. Chapman and Hall, New York, 1988.

Thompson:1942:GFC

[Tho42] D’Arcy Wentworth Thompson. On Growth and Form. The Complete Revised Edition. Cambridge University Press, 1942. ISBN 0-486-67135-6. viii + 1116 pp. Reprinted by Dover, New York, 1992.

Tienari:1970:SMR

[Tie70]  Martti Tienari.  A statis-

tical model of roundoff error for varying length floatingpoint arithmetic. BIT, 10: 355–365, 1970. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Tisseur:2001:NMF

[Tis01] Françoise Tisseur. Newton’s method in floating point arithmetic and iterative refinement of generalized eigenvalue problems. SIAM J. Matrix Anal. Appl., 22 (4):1038–1057, 2001.

Taussky:1962:EFM

[TM62] Olga Taussky and Marvin Marcus. Eigenvalues of finite matrices. In John Todd, editor, Survey of Numerical Analysis, pages 279–313. McGraw-Hill, New York, 1962.

Tisseur:2001:QEP

[TM01] Françoise Tisseur and Karl Meerbergen. The quadratic eigenvalue problem. SIAM Rev., 43(2):235–286, 2001.

Todd:1954:CFS

[Tod54] John Todd. The condition of the finite segments of the Hilbert matrix. In Olga Taussky, editor, Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, number 39 in Applied Mathematics Series, pages 109– 116. National Bureau of Standards, United States Department of Commerce, Washington, D. C., 1954.

Todd:1961:CPC

[Tod61] John Todd. Computational problems concerning the Hilbert matrix. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 65(1):19–22, 1961. CODEN JNBBAU. ISSN 0022-4340.

Todd:1968:CN

[Tod68] J. Todd. On condition numbers. In Programmation en Mathématiques Numériques, Besançon, 1966, volume 7 (no.

165) of Éditions Centre Nat. Recherche Sci., Paris, pages 141–159, 1968.

Todd:1977:BNM

[Tod77] John Todd. Basic Numerical Mathematics, Vol. 2: Numerical Algebra. Birkhäuser, Basel, and Academic Press, New York, 1977.

Toledo:1997:LRL

[Tol97] Sivan Toledo. Locality of reference in LU decomposition with partial pivoting. SIAM J. Matrix Anal. Appl., 18(4):1065– 1081, 1997.

Tornheim:1965:MTP

[Tor65] L. Tornheim. Maximum third pivot for Gaussian reduction. Technical report, Calif. Research Corp., Richmond, CA, 1965. Cited in [Cry68].

Torczon:1989:MDS

[Tor89] Virginia J. Torczon. Multi-Directional Search: A Direct Search Algorithm for Parallel Machines. PhD thesis, Rice University, Houston, TX, USA, May 1989. vii + 85 pp.

Torczon:1991:CMS

[Tor91] Virginia J. Torczon. On the convergence of the multidirectional search algorithm. SIAM J. Optim., 1(1):123–145, 1991.

Torczon:1997:CPS

[Tor97] Virginia J. Torczon. On the convergence of pattern search algorithms. SIAM J. Optim., 7(1): 1–25, 1997.

Traub:1966:APU

[Tra66] J. F. Traub. Associated polynomials and uniform methods for the solution of linear problems. SIAM Rev., 8(3):277–301, 1966.

Trefethen:1985:TMG

[Tre85] Lloyd N. Trefethen. Three mysteries of Gaussian elimination. ACM SIGNUM Newsletter, 20: 2–5, 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Trefethen:1990:ATN

[Tre90] Lloyd N. Trefethen. Approximation theory and numerical linear algebra. In J. C. Mason and M. G. Cox, editors, Algorithms for Approximation II, pages 336–360. Chapman and Hall, London, 1990.

Trefethen:1992:DNA

[Tre92a]  Lloyd N. Trefethen.  The

definition of numerical analysis. SIAM News, 25:6 and 22, November 1992. ISSN 0036-1437.

Trefethen:1992:PM

[Tre92b] Lloyd N. Trefethen. Pseudospectra of matrices. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1991, Proceedings of the 14th Dundee Conference, volume 260 of Pitman Research Notes in Mathematics, pages 234–266. Longman Scientific and Technical, Essex, UK, 1992.

Trefethen:1999:SP

[Tre99] Lloyd N. Trefethen. Spectra and pseudospectra. In Mark Ainsworth, Jeremy Levesley, and Marco Marletta, editors, The Graduate Student’s Guide to Numerical Analysis ’98, pages 217–250. Springer-Verlag, Berlin, 1999.

Tropp:1984:FA

[Tro84a] Henry S. Tropp. FORTRAN anecdotes. Annals of the History of Computing, 6(1):59–64, 1984.

Tropp:1984:OTE

[Tro84b] Henry S. Tropp. Origin of the term bit. Annals of the History of Computing, 6(2):152– 155, 1984.

Trefethen:1990:ACS

[TS90] Lloyd N. Trefethen and Robert S. Schreiber. Average-case stability of Gaussian elimination. SIAM J. Matrix Anal. Appl., 11 (3):335–360, 1990.

Tsao:1975:NIH

[Tsa75] Nai-kuan Tsao. A note on implementing the Householder transformation. SIAM J. Numer. Anal., 12(1):53–58, 1975.

Tsao:1983:SAE

[Tsa83] Nai-kuan Tsao. A simple approach to the error analysis of division-free numerical algorithms. IEEE Trans. Comput., C-32(4):343–351, 1983.

Trefethen:1987:IPS

[TT87] Lloyd N. Trefethen and Manfred R. Trummer. An instability phenomenon in spectral methods. SIAM J. Numer. Anal., 24 (5):1008–1023, 1987.

Toh:1994:PPP

[TT94] Kim-Chuan Toh and Lloyd N.

Trefethen. Pseudozeros of polynomials and pseudospectra of companion matrices. Numer. Math., 68(3):403–425, 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tucker:2002:ROS

[Tuc02]  Warwick Tucker.  A rigor-

ous ODE solver and Smale’s 14th problem. Found. Comput. Math., 2(1):53–117, 2002.

Turnbull:1929:TDM

[Tur29] H. W. Turnbull. The Theory of Determinants, Matrices, and Invariants. Blackie, London and Glasgow, 1929. xvi + 338 pp.

Turing:1936:CNA

[Tur36] A. M. Turing. On computable numbers, with an application to the Entscheidungsproblem. Proceedings of the London Mathematical Society. Third Series, 42:230–265, 1936. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic).

Turing:1945:PDM

[Tur45] A. M. Turing. Proposal for development in the Mathematics Division of an Automatic Computing Engine (ACE). Report E.882, Executive Committee, National Physical Laboratory, Teddington, UK, 1945. Reprinted in [CD86, pp. 20-105] and [Inc92, pp. 1–86].

Turing:1948:REM

[Tur48] A. M. Turing. Rounding-off errors in matrix processes. Quart. J. Mech. Appl. Math., 1:287– 308, 1948. Reprinted in [Bri92] with summary and notes (including corrections).

Turner:1982:DLS

[Tur82] Peter R. Turner. The distribution of leading significant digits. IMA J. Numer. Anal., 2: 407–412, 1982.

Turner:1984:FRD

[Tur84] Peter R. Turner. Further revelations on L.S.D. IMA J. Numer. Anal., 4:225–231, 1984.

Turner:1991:WRR

[Tur91] Peter R. Turner. Will the “real” real arithmetic please stand up?

Notices Amer. Math. Soc., 38 (4):298–304, 1991. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tyrtyshnikov:1991:CTM

[Tyr91] Evgenij E. Tyrtyshnikov. Cauchy-Toeplitz matrices and some applications. Linear Algebra Appl., 149:1–18, 1991.

Tasche:2001:WAC

[TZ01] Manfred Tasche and Hansmartin Zeuner. Worst and average case roundoff error analysis for FFT. BIT, 41(3): 563–581, 2001. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

patriot92

[Uni92] Patriot missile defense: Software problem led to system failure at Dhahran, Saudi Arabia. Report GAO/IMTEC-92-26, Information Management and Technology Division, United States General Accounting Office, Washington, D.C., February 1992. 16 pp.

Urabe:1968:RED

[Ura68] Minoru Urabe. Roundoff error distribution in fixed-point multiplication and a remark about the rounding rule. SIAM J. Numer. Anal., 5(2):202–210, 1968.

Uspensky:1948:TE

[Usp48] J. V. Uspensky. Theory of Equations. McGraw-Hill, New York, 1948. vii + 353 pp.

Vignes:1986:ESM

[VA86] J. Vignes and R. Alt. An efficient stochastic method for round-off error analysis. In Willard L. Miranker and Richard A. Toupin, editors, Accurate Scientific Computations, Proceedings, 1985, volume 235 of Lecture Notes in Computer Science, pages 183–205. Springer-Verlag, Berlin, 1986.

vanVeldhuizen:1977:NPP

[van77] M. van Veldhuizen. A note on partial pivoting and Gaussian elimination. Numer. Math., 29: 1–10, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Loan:1979:NEM

[Van79] Charles F. Van Loan. A note on the evaluation of matrix polynomials. IEEE Trans. Automat. Control, AC-24(2):320– 321, 1979.

Loan:1985:MWE

[Van85] Charles F. Van Loan. On the method of weighting for equality-constrained least-squares problems. SIAM J. Numer. Anal., 22(5):851–864, 1985.

Loan:1987:ECE

[Van87] Charles F. Van Loan. On estimating the condition of eigenvalues and eigenvectors. Linear Algebra Appl., 88/89:715– 732, 1987.

Loan:1992:CFF

[Van92] Charles F. Van Loan. Computational Frameworks for the Fast Fourier Transform. Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-285-8 (paperback), 1-61197-099-7 (ebook). xiii + 273 pp. LCCN QA403.5 .V35 1992.

Vanderbei:1995:SQD

[Van95] Robert J. Vanderbei. Symmetric quasi-definite matrices. SIAM J. Optim., 5(1):100–113, 1995.

Varah:1972:SBT

[Var72] J. M. Varah. On the solution of block-tridiagonal systems arising from certain finite-difference equations. Math. Comp., 26 (120):859–868, 1972.

Varah:1975:LBS

[Var75] J. M. Varah. A lower bound for the smallest singular value of a matrix. Linear Algebra Appl., 11:3–5, 1975.

Varga:1976:DDA

[Var76] Richard S. Varga. On diagonal dominance arguments for bounding ∥A^{−1}∥_{∞}. Linear Algebra Appl., 14:211–217, 1976.

Varah:1979:STM

[Var79] J. M. Varah. On the separation of two matrices. SIAM J. Numer. Anal., 16(2):216–222, 1979.

Varga:1990:SCM

[Var90] Richard S. Varga. Scientific Computation on Mathematical Problems and Conjectures. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-257-2. vi + 122 pp.

Vavasis:1989:GEP

[Vav89] S. A. Vavasis. Gaussian elimination with partial pivoting is P-complete. SIAM Journal on Discrete Mathematics, 2:413– 423, 1989. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Verdonk:2001:PRIa

[VCV01a] Brigitte Verdonk, Annie Cuyt, and Dennis Verschaeren. A precision and range independent tool for testing floatingpoint arithmetic I: Basic operations, square root and remainder. ACM Trans. Math. Software, 27(1):92–118, 2001. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Verdonk:2001:PRIb

[VCV01b] Brigitte Verdonk, Annie Cuyt, and Dennis Verschaeren. A precision and range independent tool for testing floatingpoint arithmetic II: Conversions. ACM Trans. Math. Software, 27(1):119–140, 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sluis:1969:CNE

[vdS69] A. van der Sluis. Condition numbers and equilibration of matrices. Numer. Math., 14:14– 23, 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1970:CEP

[vdS70] A. van der Sluis. Condition, equilibration and pivoting in linear algebraic systems. Numer. Math., 15:74–86, 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sluis:1975:SSL

[vdS75]  A. van der Sluis.  Stabil-

ity of the solutions of linear least squares problems. Numer. Math., 23:241–254, 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderVorst:1990:CBP

[vdV90] H. A. van der Vorst. The convergence behaviour of preconditioned CG and CG-S in the presence of rounding errors. In Owe Axelsson and Lily Yu. Kolotilina, editors, Preconditioned Conjugate Gradient Methods, volume 1457 of Lecture Notes in Mathematics, pages 126–136. Springer-Verlag, Berlin, 1990.

Verzuh:1949:SSL

[Ver49] Frank M. Verzuh. The solution of simultaneous linear equations with the aid of the 602 calculating punch. M.T.A.C., 3:453– 462, 1949.

Vitenko:1968:OAA

[Vit68] I. V. Viten’ko. Optimum algorithms for adding and multiplying on computers with a floating point. U.S.S.R. Computational Mathematics and Mathematical Physics, 8(5):183–195, 1968. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Vitasek:1969:NSS

[Vit69] Emil Vitasek. The numerical stability in solution of differential equations. In J. Ll. Morris, editor, Conference on the Numerical Solution of Differential Equations, volume 109 of Lecture Notes in Mathematics, pages 87–111. Springer-Verlag, Berlin, 1969.

Neumann:1947:NIM

[vNG47] John von Neumann and Herman H. Goldstine. Numerical inverting of matrices of high order. Bull. Amer. Math. Soc., 53: 1021–1099, 1947. Reprinted in [Tau63, pp. 479–557].

Rienen:2001:NMC

[vR01] Ursula van Rienen. Numerical Methods in Computational Electrodynamics: Linear Systems in Practical Applications. Number 12 in Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, 2001. ISBN 3-540-67629-5. xiii + 375 pp.

Viswanath:1998:CNR

[VT98] D. Viswanath and L. N. Trefethen. Condition numbers of random triangular matrices. SIAM J. Matrix Anal. Appl., 19 (2):564–581, 1998.

Wijngaarden:1966:NAI

[vW66] A. van Wijngaarden. Numerical analysis as an independent science. BIT, 6:66–81, 1966. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Wachpress:1988:ISL

[Wac88] Eugene L. Wachpress. Iterative solution of the Lyapunov matrix equation. Applied Mathematics Letters, 1(1):87–90, 1988. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Wallis:1990:IFP

[Wal90] Peter J. L. Wallis, editor.

Improving Floating-Point Programming. Wiley, London, 1990.

Wallis:1993:HM

[Wal93] W. D. Wallis. Hadamard matrices. In Richard A. Brualdi, Shmuel Friedland, and Victor Klee, editors, Combinatorial and Graph-Theoretical Problems in Linear Algebra, volume 50 of IMA Volumes in Mathematics and Its Applications, pages 235– 243. Springer-Verlag, New York, 1993. ISBN 0-387-94086-3.

Wareeditor:1960:SCT

[War60] Willis H. Ware, editor. Soviet computer technology— 1959. Comm. ACM, 3(3):131– 166, 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ward:1976:QAH

[War76] Robert C. Ward. The QR algorithm and Hyman’s method on vector computers. Math. Comp., 30(133):132–142, 1976.

Watson:1991:AOS

[Wat91] G. A. Watson. An algorithm for optimal ℓ_{2} scaling of matrices. IMA J. Numer. Anal., 11:481– 492, 1991.

Williams:1989:EBC

[WCK89] M. R. Williams and Martin Campbell-Kelly, editors. The Early British Computer Conferences, volume 14 of Charles Babbage Institute Reprint Series for the History of Computing. MIT Press, Cambridge, MA, USA, 1989.

Wedderburn:1934:LM

[Wed34] J. H. M. Wedderburn. Lectures on Matrices, volume 17 of American Mathematical Society Colloquium Publications. American Mathematical Society, Providence, RI, USA, 1934. vii + 205 pp.

Wedin:1973:PTP

[Wed73] Per-Åke Wedin. Perturbation theory for pseudo-inverses. BIT, 13:217–232, 1973. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Wei:1990:PLS

[Wei90] Musheng Wei. Perturbation of the least squares problem. Linear Algebra Appl., 141:177–182, 1990.

Wendroff:1966:TNA

[Wen66] Burton Wendroff. Theoretical Numerical Analysis. Academic Press, New York, 1966. xi + 239 pp.

Werner:1984:PIL

[Wer84] Wilhelm Werner. Polynomial interpolation: Lagrange versus Newton. Math. Comp., 43(167): 205–217, 1984.

Westlake:1968:HNM

[Wes68] Joan R. Westlake. A Handbook of Numerical Matrix Inversion and Solution of Linear Equations. Wiley, New York, 1968.

White:1997:INL

[Whi97] Michael White. Isaac Newton:

The Last Sorcerer. Fourth Estate, London, 1997. ISBN 1-85702-706-X. 402 pp.

Wichmann:1989:TFS

[Wic89] B. A. Wichmann. Towards a formal specification of floating point. Comput. J., 32:432– 436, 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wichmann:1992:NUF

[Wic92] B. A. Wichmann. A note on the use of floating point in critical systems. Comput. J., 35 (1):41–44, 1992. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wilkinson:1948:ACE

[Wil48a] J. H. Wilkinson. The Automatic Computing Engine at the National Physical Laboratory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195:285– 286, 1948. CODEN PRLAAZ. ISSN 0080-4630.

Wilkinson:1948:PRA

[Wil48b] J. H. Wilkinson. Progress report on the Automatic Computing Engine. Report MA/17/1024, Mathematics Division, Department of Scientific and Industrial Research, National Physical Laboratory, Teddington, UK, April 1948. 127 pp.

Wilkinson:1954:LAP

[Wil54a] J. H. Wilkinson. Linear algebra on the Pilot ACE. In Automatic Digital Computation. Her Majesty’s Stationery Office, London, 1954.

Wilkinson:1954:PA

[Wil54b] J. H. Wilkinson. The Pilot ACE. In Automatic Digital Computation, pages 5–14. Her Majesty’s Stationery Office, London, 1954.

Wilkinson:1955:UIM

[Wil55] J. H. Wilkinson. The use of iterative methods for finding the latent roots and vectors of matrices. Mathematical Tables and Other Aids to Computation, 9:184–191, 1955. CO-DEN MTTCAS. ISSN 0891-6837.

Wilkinson:1960:EAF

[Wil60] J. H. Wilkinson. Error analysis of floating-point computation. Numer. Math., 2:319– 340, 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1961:EAD

[Wil61] J. H. Wilkinson. Error analysis of direct methods of matrix inversion. J. Assoc. Comput. Mach., 8:281–330, 1961.

Wilkinson:1962:EAE

[Wil62] J. H. Wilkinson. Error analysis of eigenvalue techniques based on orthogonal transformations. J. Soc. Indust. Appl. Math., 10 (1):162–195, 1962.

Wilkinson:1963:PRF

[Wil63a] J. H. Wilkinson. Plane-rotations in floating-point arithmetic. In Experimental Arithmetic, High Speed Computing and Mathematics, volume 15 of Proceedings of Symposia in Applied Mathematics, pages 185–198. American Mathematical Society, Providence, RI, USA, 1963.

Wilkinson:1963:REA

[Wil63b] J. H. Wilkinson. Rounding Errors in Algebraic Processes. Notes on Applied Science No. 32, Her Majesty’s Stationery Office, London, 1963. ISBN 0-486-67999-3. vi + 161 pp. Also published by Prentice-Hall, Englewood Cliffs, NJ, USA. Reprinted by Dover, New York, 1994.

Wilkinson:1965:AEP

[Wil65a] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Oxford University Press, 1965. ISBN 0-19-853403-5 (hardback), 0-19-853418-3 (paperback). xviii + 662 pp.

Wilkinson:1965:EAT

[Wil65b] J. H. Wilkinson. Error analysis of transformations based on the use of matrices of the form I − 2ww^{H}. In Louis B. Rall, editor, Error in Digital Computation, volume 2, pages 77–101. Wiley, New York, 1965.

Wilkinson:1967:BZW

[Wil67]  J. H. Wilkinson.  Bledy

Zaokragleń w Procesach Algebraicznych. PWW, Warszawa, 1967. Polish translation of [Wil63b].

Wilkinson:1968:PEA

[Wil68] J. H. Wilkinson. A priori error analysis of algebraic processes. In I. G. Petrovsky, editor, Proc. International Congress of Mathematicians, Moscow 1966, pages 629–640. Mir Publishers, Moscow, 1968.

Wilkinson:1969:R

[Wil69] J. H. Wilkinson. Rundungsfehler. Springer-Verlag, Berlin, 1969. German translation of [Wil63b].

Wilkinson:1970:AEP

[Wil70] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Nauka, U.S.S.R. Academy of Sciences, 1970. 564 pp.

Wilkinson:1971:MEA

[Wil71a] J. H. Wilkinson. Modern error analysis. SIAM Rev., 13(4):548– 568, 1971.

Wilkinson:1971:SCN

[Wil71b] J. H. Wilkinson. Some comments from a numerical analyst. J. Assoc. Comput. Mach., 18(2): 137–147, 1971.

Wilkinson:1974:CEA

[Wil74a] J. H. Wilkinson. The classical error analyses for the solution of linear systems. IMA Bulletin, 10 (5/6):175–180, 1974.

Wilkinson:1974:NLA

[Wil74b] J. H. Wilkinson. Numerical linear algebra on digital computers. IMA Bulletin, 10(9/10): 354–356, 1974.

Wilkinson:1980:TWN

[Wil80] J. H. Wilkinson. Turing’s work at the National Physical Laboratory and the construction of Pilot ACE, DEUCE, and ACE. In N. Metropolis, J. Howlett, and Gian-Carlo Rota, editors, A History of Computing in the Twentieth Century: A Collection of Essays, pages 101– 114. Academic Press, New York, 1980.

Wilkinson:1984:PP

[Wil84] James H. Wilkinson. The perfidious polynomial. In G. H. Golub, editor, Studies in Numerical Analysis, volume 24 of Studies in Mathematics, pages 1–28. Mathematical Association of America, Washington, D.C., 1984.

Wilkinson:1985:SAE

[Wil85] J. H. Wilkinson. The state of the art in error analysis. In NAG Newsletter 2/85, pages 5–28. Numerical Algorithms Group, Oxford, UK, November 1985.

Wilkinson:1986:EAR

[Wil86] J. H. Wilkinson. Error analysis revisited. IMA Bulletin, 22 (11/12):192–200, 1986.

Wilkes:1997:AE

[Wil97] M. V. Wilkes. Arithmetic on the EDSAC. IEEE Annals of the History of Computing, 19(1):13– 15, 1997.

Winograd:1968:NAI

[Win68] S. Winograd. A new algorithm for inner product. IEEE Trans. Comput., C-18:693–694, 1968.

Winograd:1971:MM

[Win71] S. Winograd. On multiplication of 2×2 matrices. Linear Algebra Appl., 4:381–388, 1971.

Walden:1995:OBP

[WKS95] Bertil Waldén, Rune Karlson, and Ji-guang Sun. Optimal backward perturbation bounds for the linear least squares problem. Numerical Linear Algebra with Applications, 2(3):271–286, 1995.

Woehr:1997:CWK

[Woe97] Jack Woehr. A conversation with William Kahan. Dr. Dobb’s Journal of Software Tools, 271: 18–32, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Wolfe:1964:RTE

[Wol64]  Jack M. Wolfe.  Reducing

truncation errors by programming. Comm. ACM, 7(6):355– 356, 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wolfram:1999:MB

[Wol99] Stephen Wolfram. The Mathematica Book. Wolfram Media, Champaign, IL, USA and Cambridge University Press, Cambridge, UK, fourth edition, 1999. ISBN 1-57955-004-5 (Wolfram), 0-521-64314-7 (Cambridge). xxvi + 1470 pp.

wood58

[Woo58] Michael Woodger. The history and present use of digital computers at the National Physical Laboratory. Process Control and Automation, pages 437– 442, November 1958. CODEN PCAUA3. ISSN 0552-234X.

Reprinted in [CD86, pp. 125-140].

Wozniakowski:1977:NSS

[Woź77a]  H. Woźniakowski.  Numeri-

cal stability for solving nonlinear equations. Numer. Math., 27:373–390, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1977:NSC

[Woź77b]  H. Woźniakowski.  Numeri-

cal stability of the Chebyshev method for the solution of large linear systems. Numer. Math., 28:191–209, 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1978:REA

[Woź78] H. Woźniakowski. Roundofferror analysis of iterations for large linear systems. Numer. Math., 30:301–314, 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1980:REA

[Woź80] H. Woźniakowski. Roundofferror analysis of a new class of conjugate-gradient algorithms. Linear Algebra Appl., 29:507– 529, 1980.

Whaley:2001:AEO

[WPD01] R. Clint Whaley, Antoine Petitet, and Jack J. Dongarra. Automated empirical optimization of software and the ATLAS project. Parallel Comput., 27(1– 2):3–25, 2001.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.

Wright:1992:IMC

[Wri92] Margaret H. Wright. Interior methods for constrained optimization. Acta Numerica, 1:341–407, 1992. CODEN ANUMFU. ISBN 0-521-41026-6. ISSN 0962-4929 (print), 1474-0508 (electronic). URL ftp://

netlib.att.com/netlib/att/
cs/doc/91/4-10.ps.Z.

Wright:1993:CPW

[Wri93] Stephen J. Wright. A collection of problems for which Gaussian elimination with partial pivoting is unstable. SIAM J. Sci. Statist. Comput., 14(1):231–238, 1993.

Wright:1996:DSM

[Wri96] Margaret H. Wright. Direct search methods: Once scorned, now respectable. In D. F. Griffiths and G. A. Watson, editors, Numerical Analysis 1995, Proceedings of the 16th Dundee Conference, volume 344 of Pitman Research Notes in Mathematics, pages 191–208. Addison Wesley Longman, Harlow, Essex, UK, 1996.

Wright:1997:PDI

[Wri97] Stephen J. Wright. Primal-Dual Interior-Point Methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-382-X. xx + 289 pp.

Wallis:1972:CRS

[WSW72] W. D. Wallis, Anne Penfold Street, and Jennifer Seberry Wallis. Combinatorics: Room Squares, Sum-Free Sets, Hadamard Matrices, volume 292 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, 1972. ISBN 3-540-06035-9. 508 pp.

Wegert:1994:BNP

[WT94] Elias Wegert and Lloyd N. Trefethen. From the Buffon needle problem to the Kreiss matrix theorem. Amer. Math. Monthly, 101(2):132–139, 1994.

Weiss:1986:ACN

[WWWS86] N. Weiss, G. W. Wasilkowski, H. Woźniakowski, and M. Shub. Average condition number for solving linear equations. Linear Algebra Appl., 83:79–102, 1986.

Yalamov:1995:GSA

[Yal95] P. Y. Yalamov. Graphs and stability of algorithms. Preprint N27, University of Rousse, Bulgaria, February 1995. 17 pp.

Yalamov:2000:ASS

[Yal00] Plamen Y. Yalamov. On the almost strong stability of the circular deconvolution algorithm. SIAM J. Matrix Anal. Appl., 22 (2):358–363, 2000.

Yeung:1997:PAG

[YC97] Man-Chung Yeung and Tony. F.

Chan. Probabilistic analysis of Gaussian elimination without pivoting. SIAM J. Matrix Anal. Appl., 18(2):499–517, 1997.

Young:1972:SNM

[YG72] David M. Young and Robert T.

Gregory. A Survey of Numerical Mathematics, volume 1. Addison-Wesley, Reading, MA, USA, 1972. ISBN 0-486-65691-8. x + 492 + appendices pp. Reprinted by Dover, New York, 1988.

Yip:1986:NSS

[Yip86] E. L. Yip. A note on the stability of solving a rank-p modification of a linear system by the Sherman–Morrison– Woodbury formula. SIAM J. Sci. Statist. Comput., 7(3):507– 513, 1986.

Yohe:1979:INA

[Yoh79a] J. M. Yohe. Implementing nonstandard arithmetics. SIAM Rev., 21(1):34–56, 1979.

Yohe:1979:SIA

[Yoh79b] J. M. Yohe. Software for interval arithmetic: A reasonably portable package. ACM Trans. Math. Software, 5(1):50– 63, 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yohe:1980:PSI

[Yoh80] J. M. Yohe. Portable software for interval arithmetic. Computing: Archiv für informatik und numerik, Suppl. 2:211–229, 1980. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Young:1971:ISL

[You71] David M. Young. Iterative Solution of Large Linear Systems. Academic Press, New York, 1971. ISBN 0-12-773050-8. xxiv + 570 pp.

Young:1989:HOI

[You89] David M. Young. A historical overview of iterative methods. Computer Physics Communications, 53(1–3):1–17, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0010465589901458.

Ypma:1983:ERE

[Ypm83a] Tjalling J. Ypma. The effect of rounding errors on Newton-like methods. IMA J. Numer. Anal., 3:109–118, 1983.

Ypma:1983:FMZ

[Ypm83b] Tjalling J. Ypma. Finding a multiple zero by transformations and Newton-like methods. SIAM Rev., 25(3):365–378, 1983.

Ypma:1995:HDN

[Ypm95] Tjalling J. Ypma. Historical development of the Newton– Raphson method. SIAM Rev., 37(4):531–551, 1995.

Yuval:1978:SPS

[Yuv78] Gideon Yuval. A simple proof of Strassen’s result. Inform. Process. Lett., 7(6):285–286, 1978.

Zawilski:1991:NSC

[Zaw91] Adam T. Zawilski. Numerical stability of the cyclic Richardson iteration. Numer. Math., 60:251–290, 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zha:1993:CPA

[Zha93a] Hongyuan Zha. A componentwise perturbation analysis of the QR decomposition. SIAM J. Matrix Anal. Appl., 14(4):1124– 1131, 1993.

Zha:1993:P

[Zha93b] Hongyuan Zha. Problem 10312.

Amer. Math. Monthly, 100(5): 499, 1993.

Zietak:1985:PCI

[Ziȩ85] K. Ziȩtak. On a particular case of the inconsistent linear matrix equation AX + Y B = C. Linear Algebra Appl., 66:249–258, 1985.

Zielke:1986:RTM

[Zie86] G. Zielke. Report on test matrices for generalized inverses.

Computing: Archiv für informatik und numerik, 36:105–162, 1986. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Zielke:1988:SRM

[Zie88]  Gerhard Zielke.  Some re-

marks on matrix norms, condition numbers, and error estimates for linear equations. Linear Algebra Appl., 110:29–41, 1988.

Ziv:1982:RDE

[Ziv82] Abraham Ziv. Relative distance— An error measure in round-off error analysis. Math. Comp., 39 (160):563–569, 1982.

Ziv:1991:FEE

[Ziv91] Abraham Ziv. Fast evaluation of elementary mathematical functions with correctly rounded last bit. ACM Trans. Math. Software, 17(3):410–423, 1991. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ziv:1995:CAE

[Ziv95] Abraham Ziv. Converting approximate error bounds into exact ones. Math. Comp., 64(209): 265–277, 1995.

Zlatev:1986:CNE

[ZWS86] Zahari Zlatev, Jerzy Wasniewski, and Kjeld Schaumburg. Condition number estimators in a sparse matrix software. SIAM J. Sci. Statist. Comput., 7(4): 1175–1189, 1986.

Zha:1995:NCS

[ZZ95] Hongyuan Zha and Zhenyue Zhang. A note on constructing a symmetric matrix with specified diagonal entries and eigenvalues. BIT, 35:448–452, 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).